### v4

# KEN KANGDA WREN1

Abstract
Existing approaches to asset-pricing under model-uncertainty adapt classical utility-maximisation frameworks and seek theoretical comprehensiveness. We move toward practice by considering binary model-uncertainties and by switching attention from ’preference’ to ’constraints’. This decomposes economic asset-pricing into the viable pricing of model-risk and non-model risk separately such that the former has a unique and intuitive formula with convenient properties. Its parameter, dynamically conserved under model-risk inference, allows an integrated representation of ex-ante risk-pricing and bias, with ex-post price-effects that can be disentangled, through well-known anomalies, Momentum and Low-Risk, whose risk-reward curves acquire a new significance: peak-rewards measure ex-ante risk-pricing, and peak-locations, bias.
JEL Classification: C11, C12, C22, D81, G02, G12
Conflict-of-Interest Disclosure Statement
Author K.K.Wren
I have nothing to disclose.
Introduction
Motivation and Results
We study the pricing of model-risks that may be subject to uncertainty in the sense of Knight (1921) and Bewley (2002), where ’uncertainty’ refers to randomness governed by probability laws whose parameters are unknown. The aim is to move closer to practice and to examine relevant issues missed by existing literature. For our purpose, ’risk’ shall refer to randomness generally, with laws known or unknown, and ’model’, some chosen probability law.
Our approach exploits potential risk and data hierarchies, exposes model-risk as a dominant source of value gained/lost in the market, and furnishes ways to disentangle risk-pricing from bias, through price-anomalies such as Momentum (Jegadeesh and Titman (1993)) and Low-Risk (Ang et al. (2006)), without ad hoc criteria for what part of an observed drift is ’by design’ and what, ’by mistake’. The bias identified this way happens to be Status Quo Bias, a well-known behavioural trait (Samuelson and Zeckhauser (1988), Kahneman et al. (1991)), with a prominent role in Knightian Decision Theory (Bewley (2002)), making our framework a tool for the exploration of these concepts in practical and concrete terms.
Existing works adapting classical theories to incorporate model-risk come under ambiguity-aversion or robust-control, where agents perform recursive optimisation, against goals such as max-min utility, optimal smooth-ambiguity or variational preference, over the set of model choices (e.g. Ju and Miao (2012) or Machina and Siniscalchi (2014)). Related is parameter learning (e.g. Guidolin and Timmermann (2007), Collin-Dufresne et al. (2016)), carried out usually under a definitive model-risk law, without uncertainty.
Existing focus is the nature of preference/utility under model-risk and model-risk discount in a general equilibrium. Our focus is how model-risk pricing may be achieved or characterised ’minimally’, with fewest assumptions and parameters. A natural start is thus the Risk-Neutral Equivalent (RNE) formulation, given the First Theorem of Asset Pricing (FTAP) [^1] Early max-min utility struggled with FTAP-compliance (e.g. Epstein and Wang (1995)). Its modern form (e.g. Chen and Epstein (2002)) and other approaches have no such issues (e.g. Ju and Miao (2012)), especially when reformulated under the banner of smooth-ambiguity (Proposition 3.6 of Burzoni et al. (2021))..
However, viability, No Arbitrage in the FTAP sense, is too weak for most pricing tasks: there are far more viable processes than economically valid ones. Moreover, learning about model-risk makes RNE probabilities path-dependent, and so intractable, even in simple cases (proved formally here, but already observed in Guidolin and Timmermann (2007)).
One way to address such obstacles is to add incremental conditions to the FTAP baseline till RNE formulations that do provide economic asset-pricing can be characterised. It inverts the routine of ’theory making followed by viability checking’, and switches attention from ’agent preference’ to ’structural constraints’. It culminates in the decomposition of viable economic asset-pricing into separately viable components and the identification of a unique and intuitive RNE model-risk pricing formula, whose sole parameter, a constant of motion of the model-risk inferential dynamics, allows an integrated treatment of ex-ante risk-premia and bias such that their ex-post price-effects, only jointly observable, can be disentangled.
Scope and Premise
We take the existence of representative agents/beliefs for granted, and use the language of rational (Bayesian) inference to mean no more than adherence to the laws of probabilities; likewise, ’rational’ and ’expectation’ signify the same without implying ’objective correctness’, unlike in classical Rational Expectation (RE) frameworks. Three types of probabilities are thus involved: 1) objective ones governing events; 2) subjective ones representing inferential beliefs about the objective world; 3) RNE ones for pricing.
Let our model-risk $B$, of micro or macro economic origin, be binary, with outcome $b$ or $\overline{b}\not=b$, and our $B$-sure models, elementary in the sense of being model-risk free. Such $B$-sure processes may represent the familiar, which by classical RE principles must be well-modelled already. In our context therefore, only change, breaks from routines, may create model-risk (e.g. new CEO, R&D results, or regulatory/governmental/geopolitical events). Many real-world asset valuation challenges fall into this category. Generalisation to model-risks that are non-binary is outlined towards the end, by involving composite $B$-sure processes, which contain discounts for other model-risks.
Remark 1.
Some model-risks of this type may be hedged or diversified away. Our results apply only to the economically priced: any deemed worthy of being assigned at each moment a risk-premium through some recursive optimisation scheme taking said risk as an input.
Our model-risk shall be considered ’one-off’. How it ’flares up’ and the pricing of ’flare-risk’ remains implicit; that is, we study episodic ’model crises’, one at a time. This contrasts with the workhorse Markov regime-switching setting of the literature, where the pricing of the entire process is treated. Further, despite possible uncertainty and indefinite priors, our model-risk is resolvable via regular inference, albeit not before any finite horizon; this is also the case in Hansen-Sargent’s robust-control setting (e.g. Hansen et al. (2006)). Correspondingly, let horizons be sufficiently long for the meaningful development of model-risk inference, and let asset-pricing be time-homogeneous (’identical risk-levels identically priced’) during such developments (i.e. while the model-risk remains active and relevant).
The law of our model-risk, the unconditional distribution of $B$-outcomes, may be unknown. If the true unconditional probability, denoted $\mathbf{p}^{B}_{0}$ throughout, is known, model-detection (learning) and asset-pricing remain classical, with familiar conclusions. If unknown, the same may involve non-classical cognitive and economic elements. Nonetheless, asset-pricing may be formulated in terms of classical inference, with non-classical aspects reflected through preference and discount mechanisms; this is the usual, and our, strategy.
Lastly, data about model-risks (e.g. weekly/monthly indicators of economic regimes) tend to be more frequent and impactful than data about ’routine’ risks (e.g. biannual dividend). Such structures, absent in existing studies but relevant to practice, are treated explicitly here.
The setup as outlined above makes model-risk a prominent feature of asset-pricing, a major source of excess value, and so a focus of market effort and activities.
Existing Theories in Our Context
Most model-risk pricing schemes, ambiguity-aversion, robust-control and some parameter-learning, achieve model-risk discounts by in effect pessimistically distorting a reference belief (about model-risk), known as ’reference measure’, ’second-order prior’, or ’benchmark’.
The set of models for the state space $\mathcal{S}$ whose model is sought may be written as $\mathcal{I}_{1}(\mathcal{S})$. It is often identified with the space or a subspace of probability measures on $\mathcal{S}$. We see it simply as an index set, labelling model candidates. Potential laws governing the model-risk may be likewise collected into some set $\mathcal{I}_{2}(\mathcal{I}_{1}(\mathcal{S}))$. In our case: $\mathcal{I}_{1}(\mathcal{S})=\{b,\overline{b}\}\cong\{Q_{b},Q_{\overline{b}}\}$, one $B$-value for one model, some probability law $Q_{B}$ of state space $\mathcal{S}$, and $\mathcal{I}_{2}(\mathcal{I}_{1}(\mathcal{S}))\cong(0,1)$.
A reference model-risk belief thus may be identified with a member of $\mathcal{I}_{2}(\mathcal{I}_{1}(\mathcal{S}))$, generating learning about model-risk given the models in $\mathcal{I}_{1}(\mathcal{S})$. It is in our case some $\pi^{B}_{0}\in(0,1)$, leading to a model-risk inference process $\{\pi^{B}_{t}\}$ based on competing models $\{Q_{b},Q_{\overline{b}}\}$.
Remark 2.
Call the set $\mathcal{L}_{\mathcal{S}}:=\{\pi^{I}_{0}Q_{I}|\pi^{I}_{0}\in\mathcal{I}_{2}(\mathcal{I}_{1}(\mathcal{S})),\text{ }I\in\mathcal{I}_{1}(\mathcal{S})\}$ of reference beliefs about the law governing total state space $\mathcal{I}_{1}(\mathcal{S})\times\mathcal{S}$ the set of model-uncertainty, for later reference.
The max-min utility approach offers the simplest strategy: choose at any time the model with the worst outlook from the set $\mathcal{I}_{1}(\mathcal{S})$ of alternatives (in our setting it means pricing for the pessimistic $B$-outcome till $B$-risk resolves). It is a limit of the graded optimisations adopted by the other approaches, where pessimism is tempered by reference beliefs (e.g. Hansen and Sargent (2011) penalises distorted beliefs by their relative entropy to the reference).
Key to all is then the reference belief, by which model-risk pricing is defined and computed. In RE and some parameter-learning settings reference beliefs are known to be true, leading to classical risk-discounts. Under model-uncertainty, even if the reference beliefs happen to be true objectively, perceived uncertainty still triggers extra discounts. All worked examples in the uncertainty literature presume the reference beliefs used by their agents to be unknowingly objectively true; this in our context is unjustifiable.
At the other extreme, some parameter-learning stances see realised parameter-value as deterministic, meaning $\mathbf{p}^{B}_{0}\equiv 1\text{ or }0$ if adopted here; this would exclude many types of model-risks of practical interest. In our setting then, the norm is $\pi^{B}_{0}\not=\mathbf{p}^{B}_{0}\in(0,1)$.
Any assertion of equality would mask a basic issue facing all attempts to free classical RE frameworks from any built-in coincidence of beliefs and true laws: pricing theory outputs are ex-ante but what can be observed is ex-post, so tests and calibrations are impossible without additional theories of bias. Manifestations of this difficulty include ’observational equivalence’ (’is the large ex-post risk-premium due to bias after all?’; see e.g. Cecchetti et al. (2000)) and ’joint-hypothesis’ (’is an anomaly a sign of bias in the market or in one’s pricing theory?’; see e.g. Fama (2014)). We will illustrate how this issue resolves itself in our approach.
Finally, existing theories require as inputs the total risk as a whole, model and non-model risks together, aggregated across all the assets in a given market. This makes them unwieldy for bottom-up or cross-sectional enquiries, especially when the model-risk(s) can be micro, macro or of a variety of causes. Our method, with its isolation of model-risk pricing and RNE formulation, may prove more flexible and convenient in this context. It also reveals that unless a model-risk is ’consistent’, no interpretable economic model-risk pricing is possible.
Content and Organisation
Section 1 sets out the usual asset and data structure in continuous time, highlighting the effect of model-risk on standard risk-pricing considerations; Section 2 derives RNE asset-pricing under model-risk; Section 3 demonstrates a potential implication, followed by conclusions and discussions; technical backgrounds and proofs are collected in the Appendix.
It transpires that the RNE formulations of asset-pricing under model-risk are constrained by informational properties (Definition 1, Lemma 1 and Remark 14) and so must follow a canonical form (Proposition 1), with familiar properties (Corollary 1) and price dynamics (Section 2.2.2). One consequence is a potential link between Status Quo Bias and well-known market-anomalies, whose characteristics separate prospective (ex-ante) market risk-pricing from bias (if any) under model-uncertainty, revealing them separately (Section 3.5.1&3.5.2).
1 Setup
1.1 Usual-Case Assets and Information Structure
For clarity and to avoid clutter, consider assets with a single payoff, some bullet-payoff $Y\in\mathbb{R}$ at horizon (or maturity) $T$, with $1\ll T\leq\infty$; let value and pricing be in logarithm, and time- or risk-free discount be exogenous and nil (with a suitable choice of numéraire). Further:
  1. 1.
    Any incremental change to bullet-payoff $Y$, if at all, is public, such that the cumulative firmed-up (’cum-div’) component $\{Z_{t}\}$ of asset-worth, with $Z_{T}\equiv Y$ by definition, has the equation of motion below, subject to binary outcome $B\in\mathcal{B}:=\{+,-\}$, but otherwise based on standard Wiener process $\{w^{Z}_{t}\}$, of natural, usual, filtration $\{\mathcal{F}^{Z}_{t}\}$,
    $dZ_{t}(B)=r^{Z}_{B,t}dt+\sigma^{Z}_{t}dw^{Z}_{t},\text{ }t\in(0,T],$ (1)
    with $\{\mathcal{F}^{Z}_{t}\}$-adapted volatility $\{\sigma^{Z}_{t}\}$ and drift $\{r^{Z}_{B,t}\}$, on probability space $(\mathcal{S}^{Z}|_{T},W^{Z}_{B}|_{T})$, which is the restriction to $(0,T]$ of the usual Wiener space/measure $(\mathcal{S}^{Z},W^{Z})$ generated by $\{w^{Z}_{t}\}$, with the $B$-sure law ${W}^{Z}_{B}$ of $\{Z_{t}(B)\}$ defined via $d(W^{Z}_{-})_{t}=r^{Z}_{-,t}+\sigma^{Z}_{t}dw^{Z}_{t}$ for $B=-$, and $d(W^{Z}_{+})_{t}=r^{Z}_{+,t}dt+\sigma^{Z}_{t}dw^{Z}_{t}$ for $B=+$; write $r^{Z\Delta}_{t}:=r^{Z}_{+,t}-r^{Z}_{-,t}$.
  2. 2.
    There may be an additional $B$-informative real-vector dataflow $\{{D}_{t}(B)\}$, of similar yet $B$-conditionally independent dynamics to $\{Z_{t}(B)\}$:
    $dD_{t}(B)=r^{D}_{B,t}dt+\sigma^{D}_{t}dw^{D}_{t},$ (2)
    all as in (1) but with ’$D$’ labelling; its independence makes it price-irrelevant without model-risk (i.e. under known $B$). Combined dataflow $\{\mathbf{D}_{t}\}:=\{(Z_{t},D_{t})\}$ has state space $\mathcal{S}^{T}:=(\mathcal{S}^{Z}\times\mathcal{S}^{D})|_{T}$ and natural filtration $\{\mathbf{F}_{t}\}:=\{\mathcal{F}^{Z}_{t}\}\vee\{\mathcal{F}^{D}_{t}\}$, under product $B$-sure law $\mathbf{W}^{T}_{B}:=\mathbf{W}_{B}|_{T}$, where $\mathbf{W}_{B}:=({W}^{Z}_{B}\times{W}^{D}_{B})$.
    Remark 3.
    Current treatments ignore the make up of data $\{\mathbf{D}_{t}\}$, as if only those relevant to model-sure pricing, $\{Z_{t}\}$, exist/matter. We consider situations with data that are exclusive to model-risk (price-irrelevant once model-sure), including cases with trivial model-sure pricing data $\{Z_{t}\}$. It tackles a frequent challenge to partitioners: a crisis market unsettled by a bout of model-risk whose economics and inferential status dominate, while non-model, routine, risks remain, independently evolving or unchanging.
  3. 3.
    Given the above, continuous (predictable) inferential testing of model-risk $B$ takes place on filtered space $(\mathcal{S}^{T}_{\mathcal{B}}:=\mathcal{B}\times\mathcal{S}^{T},\text{ }\{\mathbf{F}_{t}\},\text{ }\mathbf{F}_{T};\text{ }{\pi}^{B(\cdot)}_{0}\mathbf{W}^{T}_{B(\cdot)}(\cdot))$, where ${\pi}^{B(\omega_{B})}_{0}$ is the unconditional belief in outcome $B(\omega_{B})\in\mathcal{B}$, $\omega_{B}:=(B,\omega)\in\mathcal{S}^{T}_{\mathcal{B}}$. Under the condition of $\mathbf{W}_{+}\perp\mathbf{W}_{-}$ with $\mathbf{W}_{+}|_{t}\sim\mathbf{W}_{-}|_{t}$, $\forall t<\infty$, standard testing thus resolves $B$-outcomes at guaranteed resolution-time $T(B)=\inf\big{\{}t|\mathbf{F}_{t}\ni\{B=\cdot\}\big{\}}=\infty$.
    Remark 4.
    The sequential testing of Wiener processes differing by a fixed drift is well-known. Our setup allows time-variations via absolutely continuous clock-change, making finite resolutions ($T(B)<\infty$) possible. Such issues are largely moot under economic risk-pricing (Remark 1), with $T<T(B)\leq\infty$ (see Appendix A for more).
  4. 4.
    Such testing, optimal in the sense of Neyman-Pearson Lemma, is based on likelihood-ratios (LR) $\{L^{+/-}_{t}:=\frac{d\mathbf{W}^{T}_{+}|_{t}}{d\mathbf{W}^{T}_{-}|_{t}}\}$, or equivalently, log-LR $\{l^{+/-}_{t}:=\log L^{+/-}_{t}\}$, which form Wiener processes under broad conditions; model-risk $B$ resolves iff. $\lim_{t\to T(B)}l^{+/-}_{t}=(-1)^{\mathbf{1}_{-}(B)}\cdot\infty$ (iff. $\mathbf{W}_{+}|_{T(B)}\perp\mathbf{W}_{-}|_{T(B)}$) for some $T(B)\leq\infty$. Further, standard testing is performed under the equivalent martingale measure of one of the processes (e.g that of $B=-$); henceforth let $\mathbf{W}_{-}$ be by default this base measure, under which $r^{Z}_{-,t}=0=r^{D}_{-,t}$, $\forall t$.
  5. 5.
    The above has Bayesian representations: any $L^{+/-}_{t}\in(0,\infty)$ maps bijectively and smoothly to an a posteriori belief ${\pi}^{B}_{t}\in(0,1)$. Bayes’ Rule, in terms of odds-for, $O_{f}[{\pi}^{B}_{t}]:={\pi}^{B}_{t}/\underline{\pi}^{B}_{t}$, with ()¯$\underline{(\cdot)}=1-(\cdot)$, in effect says ’new-odds $=$ old-odds $\times$ likelihood-ratios of data’:
    $O_{f}[{\pi}^{+}_{t}]=O_{f}[{\pi}^{+}_{0}]\times\exp[{l^{+/-}_{t}}]\propto{L^{+/-}_{t}}\in(0,\infty),$ (3)
    given a priori odds $O_{f}[{\pi}^{+}_{0}]\in(0,\infty)$; so inferential odds are geometric Wiener processes.
  6. 6.
    The time-$t$, $B$-sure, expected payoff $Y^{B}_{t}:={\mathbf{E}}_{B}[Y|\mathbf{F}_{t}]$ under its $B$-sure law reads ((1)):
    $Y^{B}_{t}=z_{t}+\mathbf{E}_{B}[Y-z_{t}|{\mathbf{F}}_{t}]=z_{t}+\mathbf{E}_{B}[Y-z_{t}]=z_{t}+y_{B}(T-t,t),$ (4)
[…]
  7. 7.
[…]
1.2 The FTAP Baseline under Model-Risk
Translated to our setup, FTAP states: market $\big{(}Y,\{S_{t}\};\text{ }\mathcal{S}^{T}_{\mathcal{B}},\{\mathbf{F}_{t}\},\{\mathbf{F}_{T}\};\text{ }\pi^{B}_{0}\mathbf{W}_{B}^{T}\big{)}$, with $\{\mathbf{F}_{t}\}$-adapted asset-prices $\{S_{t}\},\text{ }S_{T}\equiv Y$, is viable iff. there is an equivalent measure $\widehat{\pi^{B}_{0}\mathbf{W}^{T}_{B}}\sim{\pi}^{B}_{0}\mathbf{W}_{B}^{T}$, the RNE measure, such that $S_{t}=\hat{\mathbf{E}}[Y\big{|}\mathbf{F}_{t}]$, $\forall t<T$, where RNE expectation $\hat{\mathbf{E}}[\cdot{|}\cdot]$ may be written as:
$S_{t}=\Sigma_{B}\hat{\pi}^{B}_{t}{\hat{S}^{B}_{t}}:=\hat{\pi}^{+}_{t}\hat{S}^{+}_{t}+\hat{\pi}^{-}_{t}\hat{S}^{-}_{t},$ (5)
$\hat{S}^{B}_{t}:=\hat{\mathbf{E}}_{B}[Y\big{|}\mathbf{F}_{t}]:=\hat{\mathbf{E}}\big{[}Y\big{|}B,\mathbf{F}_{t}\big{]},$ (6)
with $\Sigma_{B}{(\cdot)^{B}}{(\cdot)^{B}}$ denoting convex combinations, such as the one under RNE beliefs $\hat{\pi}^{B}_{t}$, as calculated under Bayes’ Rule (3) given $\hat{\pi}^{B}_{0}$ and RNE log-LR $\hat{l}^{+/-}_{t}:=\log\frac{d\hat{\mathbf{W}}^{T}_{+}|_{t}}{d\hat{\mathbf{W}}^{T}_{-}|_{t}}$.
This market, defined by (1-6), is complete (Martingale Representation Theorem), and so the RNE measure, unique.
Viability by no means ensures economic validity (e.g. any pricing based on RNE measures of the form $\tilde{\pi}^{B}_{0}\tilde{\mathbf{W}}^{T}\sim{\pi}^{B}_{0}\mathbf{W}_{B}^{T}$ is blind to $B$-information); indeed, any pricing that is viable to any member of the equivalence class $[{\pi}^{B}_{0}\mathbf{W}_{B}^{T}]$ is viable to all in the class. That is, under FTAP-viability, ’anything goes’ (almost). What then characterises the processes (and the RNE measures) that do correspond to plausible economic theories? It turns out that having economically interpretable RNE price-of-model-risk is the key.
Consider risk-premium, the gap between ex-ante price and expectation: at any $t<T$, given reference model-risk belief $\{{\pi}^{B}_{t}\}$ ((3)),
$RP_{t}:=Y_{t}-S_{t}:=\Sigma_{B}{\pi}^{B}_{t}Y^{B}_{t}-S_{t},$ (7)
and its RNE decomposition (by inserting ±$\pm\Sigma_{B}\hat{\pi}^{B}_{t}{{Y}^{B}_{t}}$):
$RP_{t}=\Sigma_{B}\hat{\pi}^{B}_{t}{\hat{RP}^{B}_{t}}+B\_\widehat{RP}_{t},$ (8)
$\hat{RP}^{B}_{t}:=Y^{B}_{t}-\hat{S}^{B}_{t},$ (9)
$B\_\widehat{RP}_{t}:=RP_{t}-\Sigma_{B}\hat{\pi}^{B}_{t}{\hat{RP}^{B}_{t}}=({\pi}^{+}_{t}-\hat{\pi}^{+}_{t}){Y}^{\Delta}(t),$ (10)
where $\Sigma_{B}\hat{\pi}^{B}_{t}{\hat{RP}^{B}_{t}}$ accounts for $B$-sure (non-model) risks, and $B\_\widehat{RP}_{t}$, model-risk $B$. The latter has standard deviation $Y^{\Delta}(t)\sigma^{\pi}_{t}$, $\sigma^{\pi}_{t}:=(\pi^{+}_{t}\underline{\pi}^{+}_{t})^{\frac{1}{2}}$, which, being free of risk-pricing choices, is an intrinsic quantity. It well-defines a RNE price-of-$B$-risk:
${k}^{\hat{\pi}}_{t}:=\frac{B\_\widehat{RP}_{t}}{Y^{\Delta}(t)\sigma^{\pi}_{t}}\equiv\frac{{\pi}^{+}_{t}-\hat{\pi}^{+}_{t}}{(\pi^{+}_{t}\underline{\pi}^{+}_{t})^{\frac{1}{2}}}.$ (11)
1.3 The Economic Decomposition of Risk-Pricing under Model-Risk
Regardless of the economic pricing scheme under model-risk, classical or otherwise, the resulting asset-price $\{S_{t}\}$, with risk-premia $\{RP_{t}\}$ and RNE formulation (6-11), admits an economic decomposition of the form (16-19), provided the existence of 1) some reference model-risk beliefs $\{{\pi}^{B}_{t}\}$ and 2) $B$-sure pricing $\{S^{B}_{t}\}$ under the same scheme when $B$-sure:
$S^{B}_{t}=\check{\mathbf{E}}_{{B}}[Y{|}{\mathbf{F}}_{t}]=z_{t}+\check{\mathbf{E}}_{B}[Y-z_{t}]=z_{t}+\check{y}_{B}(T-t,t),$ (12)
$RP^{B}_{t}:=y_{B}(T-t,t)-\check{y}_{B}(T-t,t)=:RP^{B}(t)\geq 0,$ (13)
[…]
$\check{y}_{+}(T-t,t)-\check{y}_{-}(T-t,t)=:S^{\Delta}(t)>0,\text{almost surely},$ (14)
$RP^{B}(t)=\lim_{\sigma^{\pi}_{t}\to 0}\hat{RP}^{B}(t),\text{ }\text{i.e. }S^{B}_{t}=\lim_{\sigma^{\pi}_{t}\to 0}\hat{S}^{B}_{t},$ (15)
where (14) ensures consistency with asset-property Item-7, Section 1.1, and (15), internal consistency of the pricing scheme under model-risk.
Remark 5.
Usual-case model-sure pricing is simple: static noise discounted by a deterministic drift. With model-risk $B$, the only new input to economic pricing (recursive optimisation taking drivers of the expected preference-value as inputs) is the reference model-risk belief $\{{\pi}^{B}_{t}\}$.
At any $t<T$, viable pricing must satisfy $S_{t}\in(S^{-}_{t},S^{+}_{t})$, so there always exists a well-defined pricing coefficient ${A}^{B}_{t}\in(0,1)$ such that:
$S_{t}=\Sigma_{B}{A}^{B}_{t}S^{B}_{t}=S^{-}_{t}+A^{+}_{t}S^{\Delta}(t),$ (16)
$RP_{t}:=Y_{t}-S_{t}=\Sigma_{B}{\pi}^{B}_{t}{RP^{B}(t)}+B\_RP_{t},\text{ with }$ (17)
$B\_RP_{t}:=RP_{t}-\Sigma_{B}{\pi}^{B}_{t}{RP^{B}(t)}=\Sigma_{B}\pi^{B}_{t}S^{B}_{t}-S_{t}=({\pi}^{+}_{t}-A^{+}_{t}){S}^{\Delta}(t),$ (18)
$k^{A}_{t}=\frac{{\pi}^{+}_{t}-A^{+}_{t}}{(\pi^{+}_{t}\underline{\pi}^{+}_{t})^{\frac{1}{2}}}\geq 0,$ (19)
where the economic price-of-$B$-risk $\{k^{A}_{t}\}$ should be free from explicit dependence on $B$-sure pricing, the latter being based on decisions in the absence of $B$-risk.
As such, once obtained, price-of-$B$-risk $\{k^{A}_{t}\}$ may generate the coefficients $\{A^{B}_{t}\}$ via (19) for asset-pricing under $B$-risk, given consistent but otherwise arbitrary $B$-sure prices $\{{S^{\prime}}^{B}_{t}\}:=\{Y^{B}_{t}\}-\{RP^{\prime B}(t)\}$, thus separating the pricing of model and non-model risks.
This is attractive if for a useful range of risk-premia, the prices so constructed:
$\{S^{\prime}_{t}\}\equiv\{S^{(RP^{\prime}|A)}_{t}\}:=\{\Sigma_{B}A^{B}_{t}{S^{\prime}}^{B}_{t}\}=\{S^{(0|A)}_{t}\}-\text{ }\{\Sigma_{B}A^{B}_{t}{RP^{\prime}}^{B}(t)\},\text{ with }$ (20)
$\{S^{(0|A)}_{t}\}:=\{\Sigma_{B}A^{B}_{t}Y^{B}_{t}\}=\{Y^{-}_{t}\}+\{A^{+}_{t}Y^{\Delta}(t)\}=\{Y_{t}\}-\text{ }\{k^{A}_{t}\sigma^{\pi}_{t}\cdot{Y}^{\Delta}(t)\},$ (21)
are viable; both are one way or another the sum of a martingale and a drift. Observe that all applied economic asset-pricing theories, continuous and regular near risk-neutrality, are already viable written in above form, at least for sufficiently small risk-discounts.
Remark 6.
In ambiguity-aversion settings, the primary utility function, enforcing risk-aversion, can be dialled towards near neutrality (linearity) while secondary utility/variational-control continues to exert uncertainty discounts. In classical settings, including parameter-learning, where a single utility must be optimised, the same may be achieved by replacing the payoff variable with its up-to-date model-sure expectation perturbed by a small enough noise.
The real question is then the range of viability for processes (20-21), and most relevant to us, if they have convenient RNE formulations. For this, we must characterise the economic price-of-$B$-risk $\{k^{A}_{t}\}$ ((19)) more fully. Its origin suggests the following.
Definition 1.
Asset-pricing under model-risk $B$ is said to be normal if its price-of-model-risk $\{k^{A}_{t}\}$ at any $t\ll T$: 1) depends deterministically on reference model-risk belief $\{\pi^{B}_{t}\}$ in a time-homogenous manner [^2] Economic risk-pricing, solution to some optimisation-based pricing scheme, is measurable to the natural filtration of its input, ($(\{Z_{t}\},\{\pi^{B}_{t}(\mathbf{D}_{t})\})$, and so a function of its level to date (Proposition 4.9 of Breiman (1992)). Any explicit dependence on time $t$ or $Z_{t}$ itself, if at all, is presumed weak or non-existent, at least for $t\ll T$. such that the dependence is twice-differentiable and non-negative; 2) satisfies $\lim_{\sigma^{\pi}_{t}\to 0}k^{A}_{t}=0$; 3) generates viable pricing via (20-21) under an economically relevant range of risk-premium choices.
Remark 7.
Property-1&2 are shared by most if not all existing model-risk pricing theories. Property-3 will be shown to hold for usual-case assets under model-risk.
2 Canonical RNE Asset-Pricing under Model-Risk
2.1 Viable Normal Model-Risk Pricing and its RNE measure
Proposition 1.
For usual-case assets under model-risk $B$, with reference model-risk process $\{\pi^{B}_{t}\}$ based on some reference belief $\pi^{B}_{0}\mathbf{W}^{T}_{B}\in\mathcal{L}_{\mathcal{S}}$ from the set of model-uncertainty, normal pricing (Definition 1) is viable iff. it has this canonical form: at any $t<T$,
$S_{t}=S^{(RP|{\Pi})}_{t}=\Sigma_{B}\Pi^{B}_{t}S^{B}_{t}=S^{(0|{\Pi})}_{t}-\Sigma_{B}{RP}^{B}(t)\Pi^{B}_{t},$ (22)
where model-risk only pricing $\{S^{(0|{\Pi})}_{t}\}=\{\Sigma_{B}\Pi^{B}_{t}Y^{B}_{t}\}$ ((21)) is given by some RNE model-risk inference process $\{\Pi^{B}_{t}\}$ based on a RNE measure of the form $\Pi^{B}_{0}\mathbf{W}^{T}_{B}\in\mathcal{L}_{\mathcal{S}}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.
Proof. See Appendix B. Briefly, any viable model-risk only pricing $\{S^{(0|A)}_{t}\}=\{Y^{-}_{t}\}+\{A^{+}_{t}Y^{\Delta}(t)\}$ must have drift $\{Y^{\Delta}(t)dA^{+}_{t}\}$ offset $\{dY^{-}_{t}+A^{+}_{t}dY^{\Delta}(t)\}$ in expectation under some RNE measure, possible iff. it belongs to $\mathcal{L}_{\mathcal{S}}$, under which the two drift separately vanish in expectation, due to $\{A^{+}_{t}\}$ being a twice-differentiable function of model-risk inference $\{\pi^{+}_{t}\}$ driven partly/wholly by $\{dY^{-}_{t}\}$. The viability of (22), as a martingale plus a well-behaved drift, follows.
Canonical pricing thus allows the pricing of model and non-model risks to have separate RNE formulations that are each easy to identify and apply. Note however that the overall RNE measure of viable normal asset-pricing (22) under model-risk $B$, although guaranteed to exist, remains unidentifiable in general, even for simple cases.
Remark 8.
Note that canonical model-risk only pricing $\{S^{(0|{\Pi})}_{t}\}$ has identical RNE and economic price-of-model-risk ((11) and (19)). This indeed exhausts the set of price processes with this property (Lemma 1 and Remark 14, Appendix A). In fact any process with non-trivial risk-pricing of the non-model risks (i.e. $\{RP^{B}_{t}\}\not=0$) must have an RNE price-of-model-risk that is not a time-homogeneous deterministic function of any reference model-risk belief based on the set $\mathcal{L}_{\mathcal{S}}$ of model-uncertainty [^3] Any viable price process whose RNE measure is in $\mathcal{L}_{\mathcal{S}}$ must have the form $\{\Sigma_{B}A^{B}_{t}S^{B}_{t}\}=\{\Sigma_{B}\Pi^{B}_{t}Y^{B}_{t}\}$, so $\forall t<T$, $\lim_{\sigma^{\pi}_{t}\to 0}S_{t}=Y^{B}_{t}$ since $\lim_{\sigma^{\pi}_{t}\to 0}\sigma^{\Pi}_{t}=0$, but $\lim_{\sigma^{\pi}_{t}\to 0}{S}_{t}=S^{B}_{t}=Y^{B}_{t}-RP^{B}_{t}$ given (15); thus $\{RP^{B}_{t}\}=0$.. This echos the observation that parameter-learning makes RNE formulations path-dependent (Guidolin and Timmermann (2007)).
2.2 Properties of Canonical Model-Risk Pricing
2.2.1 The 1-Parameter Price-of-Model-Risk
Corollary 1.
Canonical model-risk premium $\{B\_RP_{t}\}$ and price-of-model-risk $\{k^{\Pi}_{t}\}$ are as follows:
$B\_RP_{t}$ $=({\pi}^{+}_{t}-{\Pi}^{+}_{t}){S}^{\Delta}(t),\text{ }t<T,$ (23)
$k^{\Pi}_{t}$ $=(K^{\frac{1}{2}}-K^{-\frac{1}{2}})\sigma^{\Pi}_{t},\text{ }t<T,$ (24)
with $\sigma^{\Pi}_{t}:=({\Pi}^{+}_{t}\underline{\Pi}^{+}_{t})^{\frac{1}{2}}$, and $K:=\frac{O_{f}[{\pi}^{+}_{0}]}{O_{f}[{\Pi}^{+}_{0}]}=\frac{O_{f}[{\pi}^{+}_{t}]}{O_{f}[{\Pi}^{+}_{t}]}$, a conserved quantity of risk-inference ((3)).
Remark 9.
[…]
Remark 10.
The sign of the difference process $\{{\pi}^{+}_{t}\}-\{{\Pi}^{+}_{t}\}$ is fixed ((3)); so unless model-risk impact is consistent (Item-7, Section 1.1), model-risk premia $\{B\_RP_{t}\}$ must turn negative at times: inconsistent model-risks cannot have interpretable viable economic risk-premia.
2.2.2 Connection with Familiar Price Dynamics
[…]
$S_{t}=S^{(\check{rp}|\Pi)}_{t}=S^{(0|\Pi)}_{t}-\check{rp}(t)=Y^{-}_{t}-\check{rp}(t)+S^{\Delta}(t)\Pi^{+}_{t},$ (25)
$dS_{t}(B)=dY^{-}_{t}(B)+\check{R}_{t}dt-\Pi^{+}_{j}(B)r^{Z\Delta}_{t}dt+S^{\Delta}(t)d\Pi^{+}_{t}(B).$ (26)
Given Section 1.1 and inference dynamics (A.2-A.3), we have $\{d\Pi^{+}_{t}\}$ as below, with two sources of noise, one from $\{Z_{t}\}$, given by $\{\sigma^{lZ}_{t}dw^{Z}_{t}\}$, with $\sigma^{lZ}_{t}:=r^{Z\Delta}_{t}/\sigma^{Z}_{t}$ (’signal-to-noise’ ratio), one from independent $B$-informative data $\{D_{t}\}$, modelled in parallel ((2)), which can be absorbed via $\{\sigma^{l}_{t}dw^{\prime}_{t}\}=\{\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}\}$, with $(\sigma^{l}_{t})^{2}=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$:
$d\Pi^{+}_{t}(B)=\mu^{\Pi,B}_{t}+\sigma^{\Pi,l}_{t}dw^{\prime}_{t},\text{ with},$ (27)
$\mu^{\Pi,B}_{t}:=\sigma^{\Pi,l}_{t}\sigma^{l}_{t}\Big{(}{{\mathbf{1}_{+}}}(B)-{\Pi}^{+}_{t}(B)\Big{)},$ (28)
$\sigma^{\Pi,l}_{t}:=(\sigma^{\Pi}_{t})^{2}\sigma^{l}_{t}.$ (29)
The above has a drift by design relative to reference belief $\{{\pi}^{+}_{t}\}$ (of dynamics (A.3)):
$\mu^{\Pi,\pi}_{t}:=\Sigma_{B}{\pi}^{B}_{t}{\mu^{\Pi,B}_{t}}=\sigma^{\Pi,l}_{t}\sigma^{l}_{t}\cdot(\pi^{+}_{t}-\Pi^{+}_{t})=\sigma^{\Pi,l}_{t}\sigma^{l}_{t}\cdot k^{\Pi}_{t}\sigma^{\pi}_{t}\geq 0;$ (30)
note ${\mu^{\pi,\pi}_{t}}\equiv 0\equiv{\mu^{\Pi,\Pi}_{t}}$, confirming the martingale property of inference and model-risk pricing. Further, as inference volatility $\sigma^{\Pi,l}_{t}$ (29) drives price-volatility ((26-27)), the price-of-model-risk (24) peaks as model-risk induced price-volatility peaks.
Pulling it together, the full asset-price dynamic, a $(\Pi^{B}_{0}\mathbf{W}^{T}_{B})$-martingale plus drift $\{\check{R}_{t}\}$, reads:
$dS_{t}(B)=$ $\big{[}\check{R}_{t}dt+\sigma^{Z}_{t}dw^{Z}_{t}\big{]}$ (31)
$+$ $r^{Z\Delta}_{t}dt\big{(}\mathbf{1}_{+}(B)-\Pi^{+}_{j}(B)\big{)}$ (32)
$+$ $S^{\Delta}(t)(\sigma^{\Pi}_{t})^{2}\big{[}(\sigma^{l}_{t})^{2}\big{(}\mathbf{1}_{+}(B)-\Pi^{+}_{t}(B)\big{)}+\sigma^{l}_{t}dw^{\prime}_{t}\big{]}.$ (33)
Under trivial $\{Z_{t}\}$ (firmed-up value process), the above becomes $dS_{t}(B)\propto d\Pi^{+}_{t}(B)$, making asset-pricing a pure matter of model-risk inference, which further reduces to the standard inferential test of Peskir and Shirayev (2006) under constant $\{r^{D\Delta}_{t}\}$ (signal) and $\{\sigma^{D}_{t}\}$ (noise).
Remark 11.
Practical cases often see $\check{R}_{t}\ll r^{Z\Delta}_{t}\ll S^{\Delta}(t)$: the $B$-sure dynamic (31) dominated by model-drift (32) dominated by model-inference (33); dispersion among such assets is driven by model-risk inference. It often holds also that $\text{F}^{Z/D}_{t}:=(\sigma^{lZ}_{t})^{2}/(\sigma^{lD}_{t})^{2}\ll 1$, so $(\sigma^{l}_{t})^{2}\approx(\sigma^{lD}_{t})^{2}$: prices dynamics are driven by data other than those on tangible (firmed-up) values.
Averaged under model-risk belief $\{{\pi}^{+}_{t}\}$, dominant price-dynamic (33) has a classical form:
$\sigma_{t}\bigg{(}\frac{\mu_{t}}{\sigma_{t}}+dw^{\prime}_{t}\bigg{)},$ (34)
with volatility $\sigma_{t}:=\sigma^{\Pi,l}_{t}S^{\Delta}(t)$, drift $\mu_{t}:=\mu^{\Pi,\pi}_{t}S^{\Delta}(t)$, and price-of-diffusion-risk $\mu_{t}/\sigma_{t}$ ((24)):
$\frac{\mu_{t}}{\sigma_{t}}=(K^{\frac{1}{2}}-K^{-\frac{1}{2}})\sigma^{\Pi}_{t}\sigma^{\pi}_{t}\sigma^{l}_{t}=\frac{K^{\frac{1}{2}}-K^{-\frac{1}{2}}}{S^{\Delta}(t)}\bigg{(}\frac{\sigma^{\pi}_{t}}{\sigma^{\Pi}_{t}}\bigg{)}\sigma_{t}.$ (35)
Expansion in $(K-1)$, given Remark 9, makes the classical relationship more explicit:
$\frac{\mu_{t}}{\sigma_{t}}\approx\bigg{(}\frac{K-1}{S^{\Delta}(t)}\bigg{)}\sigma_{t}\text{ and }\mu_{t}\approx\bigg{(}\frac{K-1}{S^{\Delta}(t)}\bigg{)}\sigma_{t}^{2}.$ (36)
Remark 12.
The proportionality factor characterises ex-ante pricing through an intuitive and direct metric of asset-pricing (Remark 9). How its level comes about is a matter of economic theory (e.g. risk-aversion). Regardless, our derivation links it firmly to ex-ante beliefs, making the above relevant to bias studies and measurements under model-uncertainty.
3 Bias and Risk-Pricing Measurement via Price Anomalies
3.1 The Joint-Hypothesis Issue
Let the $\{\mathbf{F}_{t}\}$-adapted objective conditional probabilities of model-risk outcomes be $\{\mathbf{p}^{B}_{t}\}$. The ex-post drift of our asset, $\{rp_{t}\}:=\{\Sigma_{B}\mathbf{p}^{B}_{t}{Y^{B}_{t}}\}-\{S_{t}\}$, assuming $B$-indifferent $B$-sure risk-premia $\{\check{rp}_{t}\}$ (Section 2.2.2) for clarity, satisfies:
$rp_{t}-\check{rp}(t)=B\_RP_{t}+(\mathbf{p}^{+}_{t}-{\pi}^{+}_{t}){S}^{\Delta}(t),\text{ }t<T.$ (37)
The LHS terms are each known/observable. The RHS terms are entangled: no conclusion on risk-pricing $\{B\_RP_{t}\}$ can be made without one on $\{\mathbf{p}^{b}_{t}\}-\{{\pi}^{b}_{t}\}$, namely, bias. For our model-risk, any realised process under a realised $B$-outcome corresponds to one run of (31-33) as if $\mathbf{p}^{b}_{t}\equiv 1\text{ or }0$. Many ’like’ observations are needed for an estimate of (37). Yet, given ’one-off’ and heterogeneous sources of model-uncertainties, how can this be done?
3.2 Assets Facing Potential Model-Change
When model-sure, our asset is a simple drift with a static noise. Upon some trigger, it may face an objective chance $\mathbf{p}^{(\cdot)}_{0}$ of entering a new state (e.g. from $B=+$ to $B=-$). Denote any new state by ’1’, any status quo, ’0’, and status quo model-sure pricing, $\{S^{\text{{\bf{0}}}}_{t}\}$.
Change-trigger may or may not follow a known law. Our interest lies in what ensues after its arrival. It amounts to treating model-risk as a compound process and ignoring the part governing how/when model-risk arises (considered known, valued and reflected in model-sure prices). The average interval between model crises, say $\lambda^{-1}$, is presumed long enough for our one-off methodology; further, consider $t\ll\lambda^{-1}$ and $t\ll T$ henceforth.
With the economic nature $sign[1{\text{{\bf{0}}}}]:=sign[Y^{1}_{t}-Y^{\bf{0}}_{t}]$ of potential change explicit, ignoring $t$-dependence of change impact $S^{\Delta}(t)$ for $t\ll T$ and $r^{Z\Delta}_{t}\ll S^{\Delta}(t)$, canonical pricing reads ((25)):
$S_{t}=S^{\text{{\bf{0}}}}_{t}+sign[1{\text{{\bf{0}}}}]S^{\Delta}\cdot\Pi^{1}_{t},\text{ }t\ll T,$ (38)
where the given RNE belief process $\{\Pi^{1}_{t}\}\sim\{\pi^{1}_{t}\}$ can be characterised by risk-pricing parameter $K:=(\frac{O_{f}[\pi^{1}_{0}]}{O_{f}[{\Pi}^{1}_{0}]})^{sign[1{\text{{\bf{0}}}}]}$ (Corollary 1); let the market be competitive, so $K\in(1,2)$ (Remark 9).
3.3 The Representation of Status Quo Bias
The objective conditional probability process $\{\mathbf{p}^{1}_{t}\}$ governing change-risk is determined by the objective but unknown unconditional law $\mathbf{p}^{1}_{0}$ of model-risk. Its relationship against the RNE $\{\Pi^{1}_{t}\}$ and reference belief process $\{\pi^{1}_{t}\}$ can be characterised through the following:
$\rho K^{sign[1{\text{{\bf{0}}}}]}\equiv\frac{O_{f}[\mathbf{p}^{1}_{0}]}{O_{f}[{\Pi}^{1}_{0}]},$ (39)
$\rho:=\frac{O_{f}[\mathbf{p}^{1}_{0}]}{O_{f}[{\pi}^{1}_{0}]};$ (40)
for instance, if a change-risk is unpriced, then $K=1$; or if it is classical, with its law $\mathbf{p}^{1}_{0}$ known and so without model-uncertainty, then $\rho=1$.
Parameter $\rho$, born of model-uncertainty, is not part of any ex-ante economic risk-discounting. The joint-hypothesis issue takes a concrete form here: observable prices reflect the product (39), and to split it into ’intention’ $K$ and ’mistake’ $\rho$ external criteria seem essential. This, as will be seen, need not be so; indeed, a high apparent bias under model-uncertainty may not be irrational either. The case of interest for us is when change is rare, $\mathbf{p}^{1}_{0}<\frac{1}{2}$, but priced as even rarer, such that (39-40) are both sizeable ($\rho\gg 1$). This ’justified strong status quo’ serves as a clear base against which ’change’ is defined and assessed (’unjustified strong status quo’, $\mathbf{p}^{1}_{0}>\frac{1}{2}$, allows easy wins to bets on ’change’ and so cannot persist).
3.4 Unconditional Model-Risk Symmetry, Dominance and Uniformity
Consider a market of a large number of assets subject to change-risks (e.g. new CEO, tort, war, recession). Let these risks be as characterised and economically priced in this market (e.g. due to their aggregate impact on market value/volatility/dispersion [^5] Dispersion is particularly intriguing in our context. As a priced factor it has been linked to market uncertainty, the business cycle and ’structural shifts’ (e.g. Demirer and Jategaonkar (2013), Kolari et al. (2021)).).
For clarity, let the unconditional all-cause probability of positive/negative economic change be even: at any time half of all potential changes in the market are positive/negative. Unconditional unevenness is addressed in Remark 17&18.
Let ’pricing intention’ $K$ ((39)) and ’apparent bias’ $\rho$ ((40)) be indifferent to the economic nature of potential change. Further, let parameter $K$, $\rho$ and change-impact $S^{\Delta}$ be uniform across the assets. Such uniformity amounts in practice to asset dispersions being driven by differences in inferential progression rather than in other aspects (Remark 11). Similarly, let all risk-inference be based on i.i.d data, with log-LR processes of fixed and uniform volatility. Note that each individual realisation of a change-outcome is considered random and independent still; while this may be a stretch for triggers such as war, it is a sensible start.
3.5 Conditional Cross-Sectional Trading
Pair-trading sorts assets into cohorts by cross-sectional statistics for long vs short positions. It amounts in our setup to conditioning by $B$-inferential events $\{\Pi^{1}_{t}=v\}$, $v\in(0,1)$, the driver of dispersion. By Bayes’ Rule (3) and definition (39), the $\{\Pi^{1}_{t}=v\}$-conditioned ex-post mean-excess $rp(\pm,v):=\pm(\mathbf{p}^{1}_{t}(v)-v)S^{\Delta}$ given the $sign[1{\text{{\bf{0}}}}]=\pm$ of potential change reads:
$rp(\pm,v)=\pm(\sigma^{v})^{2}\frac{1-(\rho K^{\pm 1})^{-1}}{v+\underline{v}(\rho K^{\pm 1})^{-1}}S^{\Delta},$ (41)
where $(\sigma^{v})^{2}:=v\underline{v}$, making price-volatility the driver of excess ((29) and (34)). Note that the above is invariant under label-switching $1\leftrightarrow{\text{{\bf{0}}}}$ (i.e. $v\leftrightarrow\underline{v}$, +$\leftrightarrow$-+ ↔ - and $\rho\leftrightarrow\rho^{-1}$).
Remark 13.
With no uncertainty, $\rho=1$, the above is the ex-ante risk-premium when $\{\Pi^{1}_{t}=v\}$ ((24)): in classical markets, all the ’strategising’ merely earns what the market deems fair.
Despite excess (41) being driven by volatility, conditioning by $\{\Pi^{1}_{t}=v\}$, $v\in(0,1)$ under strong status quo amounts to ranking by momentum (past performance): $\{\Pi^{1}_{t}=v\}$ means $\{\Pi^{1}_{t}\}$ going from $\Pi^{1}_{0}\ll\frac{1}{2}$ to $v$ and so an excess-to-date of ±$\pm(v-\Pi^{1}_{0})S^{\Delta}$ ((38)). The associated model-uncertainty ($\propto(\sigma^{v})^{2}$) is a priori low and then evolves with dataflow. The window of opportunity for pair-trading is given by (C.8), Appendix C.2.
Drifts are hard to measure (e.g. Merton (1980)). Volatility trading is a good practical alternative. Volatility events, by (26-29), take the form $\{\Pi^{1}_{t}=v\}\cup\{\Pi^{1}_{t}=\underline{v}\}$, $v\in(0,\frac{1}{2}]$, and the $\{\sigma^{\Pi}_{t}=v\underline{v}\}$-conditioned ex-post mean-excess $rp_{t}^{\sigma}(v)$ averages over the $sign[1{\text{{\bf{0}}}}]$ of potential change and over the $\{\Pi^{1}_{t}=v\}$- and $\{\Pi^{1}_{t}=\underline{v}\}$-events in the volatility cohort:
$rp_{t}^{\sigma}(v):=\frac{\sum_{\pm}\mathbb{P}_{t}(\pm,v)rp(\pm,v)+\sum_{\pm}\mathbb{P}_{t}(\pm,\underline{v})rp(\pm,\underline{v})}{\mathbb{P}_{t}(v)+\mathbb{P}_{t}(\underline{v})},$ (42)
where $\mathbb{P}_{t}(\cdot)=\sum_{\pm}\mathbb{P}_{t}(\pm,\cdot):=\sum_{sign[1{\text{{\bf{0}}}}]}\mathbb{P}_{t}\big{(}\{sign[1{\text{{\bf{0}}}}]=\cdot\}\cap\{\Pi^{1}_{t}=\cdot\}\big{)}$.
The relevant likelihoods (ratios) are determined by the underlying log-LR dynamics, whose cross-sections are Normal, making closed expressions possible. As a by-product, the ratio functions predict an association of volatility with ’negative momentum’ (Remark 17), a well-known phenomenon (e.g. Ang et al. (2006) and Wang and Xu (2015)).
The risk-reward curves of momentum trading (41) and volatility trading (42) are both concave in $v$ and vanishing as $v\underline{v}\to 0$.
3.5.1 Momentum Effect and Status Quo Bias
Conditional excess (41) exhibits Momentum, a key price-anomaly (Jegadeesh and Titman (1993)): if $\rho K^{\pm 1}\gg 1$, excess-to-date ±$\pm(v-\Pi^{1}_{0})S^{\Delta}$, due to RNE beliefs $\{\Pi^{1}_{t}\}$ rising from $\Pi^{1}_{0}\ll\frac{1}{2}$ to $v>\Pi^{1}_{0}$, persists on average in the same direction. The effect applies to $\rho K^{\pm 1}\ll 1$ also: past excess due to RNE beliefs $\{\Pi^{1}_{t}\}$ falling from $\Pi^{1}_{0}\approx 1$ to $v<\Pi^{1}_{0}$ means an average future excess of the same sign; indeed, label-switching here defines a new parameter $\rho^{-1}>1$. Momentum thus vanishes only under $\rho\approx 1$ (little model-uncertainty).
The effect grows with $\rho$, which by definition (40) measures bias against change. From (41), momentum trading earns its best realisable excess $rp^{\pm mo}_{max}$ at $v^{\pm mo}_{max}$:
$v^{\pm mo}_{max}:=\frac{1}{(\rho K^{\pm 1})^{\frac{1}{2}}+1},\text{ }rp^{\pm mo}_{max}:=\pm\frac{(\rho K^{\pm})^{\frac{1}{2}}-1}{(\rho K^{\pm})^{\frac{1}{2}}+1}S^{\Delta}.$ (43)
Peak-profitability reads $\frac{1}{2}(rp^{+mo}_{max}-rp^{-mo}_{max})$, by pairing cohorts of conditioning-momentum $v^{\pm mo}_{max}$. As a function of bias $\rho$, risk-reward curve (41) and its peak (43) imply rising rewards from rising momentum up to a point, beyond which profitability falls, as seen in Ang et al. (2006).
3.5.2 Low-Risk Effect and Status Quo Bias
Low-Risk Effect (e.g. Ang et al. (2006)) turns the classical dictum ’high risk high reward’ on its head. Momentum risk-reward (41) already hints at it: peak-reward (43) is in the low-volatility region if bias $\rho$ is sizeable. The effect is clearer under volatility-conditioning (42), whose peak-location and -size can be found by routine methods (Appendix C). The solution at $\rho=1$ (no bias) and leading-order solutions at $\rho\gg 1$ (high bias) are revealing and may be given by the same expression; the peak-location $v^{\sigma}_{max}$ and -size $rp^{\sigma}_{max}$ are:
$v^{\sigma}_{max}:=\frac{1}{\rho+1},\text{ }rp^{\sigma}_{max}:=\frac{1}{2}\frac{K-1}{K+1}S^{\Delta}.$ (44)
At $\rho=1$ the above is exact, by which $v^{\sigma}_{max}=\frac{1}{2}$, where price-volatility peaks, thus confirming ’high risk high reward’ in the absence of model-uncertainty or bias.
There is otherwise a low-risk effect: peak-reward occurs at $v^{\sigma}_{max}\approx\rho^{-1}$, corresponding to low risk if $\rho\gg 1$; peak volatility, $v=\frac{1}{2}$, is rewarded poorly ex-post: $rp^{\sigma}(\frac{1}{2})=\frac{S^{\Delta}}{2}\mathcal{O}(\frac{K-1}{\rho}-1)$. Note the separation of risk-pricing $K$ and bias $\rho$: the latter is revealed by peak-location $v^{\sigma}_{max}$, while the former, by peak-reward $rp^{\sigma}_{max}$, which is that set ex-ante for peak-risk (Remark 9). No external rules are needed to tell ’intended pricing’ from ’unintended bias’.
4 Summary and Discussion
We investigate the risk-neutral equivalent formulation of model-risk pricing. Explicit price dynamics under binary model-risks with potentially unknown laws are derived, exposing model-risk as a main source of excess value, with a key role in Status Quo Bias and anomalies such as Momentum and Low-Risk. It offers an integrated and practicable approach to these highly relevant and interconnected topics.
We make use of a feature of standard hypothesis testing (Lemma 1) that ’informationally equivalent’ tests must be essentially identical. This makes risk-neutral equivalent formulations under model-risk path-dependent and intractable. On the other hand, it limits viable economic asset-pricing (Remark 1) under model-risk to just one form (Proposition 1), leading to an intuitive risk-pricing formula with intuitive properties (Corollary 1).
The sole risk-pricing parameter is associated with a conservation law of inference and is a product of ’intention’ (preference) and ’mistake’ (difference between objective and subjective laws). Ex-ante (’intended’) model-risk pricing, classical or with ambiguity-aversion/robust-control, can be captured by our parameter $K\in[1,\infty)$ (39), with $K=1$ corresponding to ’no model-risk discounts’. Its derivation reveals that it sets the intended gain-loss ratio of the asset at peak model-risk, suggesting $K\in[1,2)$ in competitive markets (Remark 9).
The difference between true model-risk law and the reference model-risk belief on which asset-pricing is based can be captured by our parameter $\rho\in(0,\infty)$ (40), with $\rho=1$ corresponding to ’no difference’ (the classical setting), and $\rho\not=1$, ’bias’. It gives expression to Status Quo Bias and allows its effects to be disentangled from risk-pricing (Section 3.5.1-3.5.2).
The above is interesting also in light of the inertia axiom of Bewley (2002), which says, in effect, that to update an existing decision by incorporating a new state subject to Knightian uncertainty, the status quo is kept unless an alternative that is better in all scenarios exists. For our model-risk, it suggests inertial pricing, one that is piece-wise free of model-risk pricing, jumping between alternative models when appropriate, despite data that can resolve model-risk via regular inference. Asset-prices under max-min utility already have this characteristic, albeit with a built-in pessimism absent in inertial pricing. Both are viable in our setup, as models are given by probability laws that are equivalent before risk-resolution. The theoretical and practical aspects of this topic will be examined in a follow-up study.
Our approach may be extended to model-risks characterisable as $(N+1)$-tuples of independent binary outcomes ${B}^{N+1}:=(B_{0}B_{1}...B_{N})$, $N\in\mathbb{N}$, through iteration. It may be applied first to the $2^{N}$ pairs of ${B}^{N+1}$-sure, elementary, price processes, then to the resulting $2^{N-i}$ pairs of ${B}^{N+1-i}$-sure composites, $i=1,...,N$, each discounting the $i$-tuple model-risk $(B_{0}B_{1}...B_{i-1})$.
APPENDIX
A Properties of Inferential Hypothesis Testing
  1. 1.
    Standard Continuous-Time Inference. Any continuous process with independent increments are essentially Wiener. By definition (Item-4, Section 1.1), the log-LR process under i.i.d data, in terms of some constant $\sigma^{l}>0$ (the ’signal-to-noise’ of data), reads:
    $dl^{+/-}_{\tau}(B)=(-1)^{\mathbf{1}_{-}(B)}\frac{(\sigma^{l})^{2}}{2}d\tau+\sigma^{l}dw_{\tau},$ (A.1)
    with $\{w_{\tau}\}$ denoting a standard Wiener process. Time-variation may be modelled via an absolutely continuous clock-change $t\mapsto\tau(t)$. The image process $\{l^{+/-}_{t}\}:=\{l^{+/-}_{\tau(t)}\}$, again Wiener and with absolutely continuous time-varying parameters (e.g. Kallsen (2006)), along with inferential beliefs, an Ito process of it ((3)), read as follows:
    $dl^{+/-}_{t}(B)=\mu^{l^{+/-}}_{t}(B)dt+\sigma^{l}_{t}dw_{t},\text{ }\text{ }\mu^{l^{+/-}}_{t}(B):=(-1)^{\mathbf{1}_{-}(B)}\frac{(\sigma^{l}_{t})^{2}}{2};$ (A.2)
    $\frac{d\pi^{+}_{t}(B)}{(\sigma^{\pi}_{t})^{2}}=\big{(}\mathbf{1}_{+}(B)-\pi^{+}_{t}(B)\big{)}(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw_{t}.$ (A.3)
  2. 2.
    Regular Inference. Its defining feature is the predictable resolution of risky outcomes. In usual situations (e.g. i.i.d data), risk-resolution occurs at infinity. Clock-change allows continuous-time regular inference to have finite resolution-times, by compactifying uniform regular inference to bring resolution $T_{\tau}(B)=\infty$ forward to some $T(B):=\tau^{-1}(\infty)<\infty$. The resulting log-LR process diverges as $t\to T(B)$, but the associated beliefs and so asset-pricing remain finite and continuous under Bayes’ Rule (3).
  3. 3.
    Complete Markets and Unhedgeable Model-Risks. Market completeness excludes unpredictable resolutions (price-jumps) and so demands regular inference. In this setting, to be unhedgeable and thus priced economically (Remark 1), model-risk $B$ cannot be resolvable before horizon $T$ (so $T(B)>T$); this is the default case for model-risk studies. The mixed case of probabilistic resolution, ($\mathbb{P}\big{(}\{T(B)\leq t+s\}|F_{t}\big{)}\in(0,1)$, $\forall t<T$, $\forall s\in(0,T-t]$, which may occur in continuous-time complete markets so long as $T(B)$ is predictable, involves treating the unhedgeable part of the model-risk only.
  4. 4.
    Non-Resolution and Novikov’s Condition. For any log-LR process based on equivalent Wiener measures on the Wiener space $\mathcal{S}$ of total data, written as $\hat{\mathbf{W}}_{B}\sim\mathbf{W}_{B}$ in the financial context, one being some RNE version of the other, the Radon-Nikodym Theorem ensures $\lim_{t\to\infty}\log\frac{d\hat{\mathbf{W}}_{B}|_{t}(\cdot)}{d\mathbf{W}_{B}|_{t}(\cdot)}<\infty$ almost surely, and hence a finite total variance ((A.2)). That is, the non-resolution of $B$-outcomes (Item-4, Section 1.1) equates to (A.2-A.3) satisfying Novikov’s Condition, and vice versa.
  5. 5.
    Adjacency. Given any pair of log-LR process $\{l^{+/-}_{t}\}$ and $\{\hat{l}^{+/-}_{t}\}$ of respective defining laws $\{\mathbf{W}_{+},\mathbf{W}_{-}\}$ and $\{\hat{\mathbf{W}}_{+},\hat{\mathbf{W}}_{-}\}$ such that equivalence $\hat{\mathbf{W}}_{B}\sim\mathbf{W}_{B}$ holds on the space $\mathcal{S}$ of total data, then they almost surely can differ at most by a constant as $t\to T(B)$, since:
    $\hat{l}^{+/-}_{t}-l^{+/-}_{t}=\log{\frac{d\hat{\mathbf{W}}_{+}|_{t}}{d\mathbf{W}_{+}|_{t}}}-\log{\frac{d\hat{\mathbf{W}}_{-}|_{t}}{d\mathbf{W}_{-|_{t}}}},\text{ }\forall t\in\mathbb{R}^{+},$ (A.4)
    where the RHS is convergent under equivalence.
    Lemma 1.
    No adjacent log-LR processes, unless identical up to a constant difference, can be informationally equivalent, that is, be deterministically and time-homogenously related to one another, nor can any inferential processes derived from them.
    Proof. It is a feature of inference (and due to Bayes’ Rule). It can be verified concretely via Ito’s Lemma here, given (A.2) and restrictions to twice-differentiable maps between processes (as in Proposition 1). Any such mapping $g:{l}^{+/-}_{t}\in\mathbb{R}\mapsto\hat{l}^{+/-}_{t}=g(l^{+/-}_{t})\in\mathbb{R}$, $t\in(0,T(B))$, must solve the ordinary differential equation:
    $g^{\prime\prime}=-(-1)^{\mathbf{1}_{+}(B)}(g^{\prime}-1)g^{\prime}.$ (A.5)
    If $B=+$, then $g^{\prime}e^{-g}=g^{\prime}(0)e^{-g(0)}e^{-\mathbf{Id}}$ and so $g=g(0)-\log[g^{\prime}(0)e^{-\mathbf{Id}}+1-g^{\prime}(0)]$; divergence $\lim_{t\to T(+)}\hat{l}^{+/-}_{t}=\infty$ requires $g^{\prime}(0)=1$; If $B=-$, then $g=g(0)+\log[g^{\prime}(0)e^{\mathbf{Id}}+1-g^{\prime}(0)]$ and $\lim_{t\to T(-)}\hat{l}^{+/-}_{t}=-\infty$ imply the same.
    Remark 14.
    Given identical LR up to a constant factor, inference process $\{\hat{\pi}^{+}_{t}\}$, a $(\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B})$-martingale by definition, is also a $(\hat{\pi}^{B}_{0}\mathbf{W}_{B})$-martingale ((3)), meaning $\hat{\mathbf{W}}_{B}=\mathbf{W}_{B}$ (Martingale Representation Theorem and so completeness of the Wiener basis): thus, only tests based on identical underlying beliefs can be informationally equivalent.
B PROOF OF PROPOSITION 1
Proof. Part I. Viable model-risk only pricing $\{S^{(0|A)}_{t}\}:=\{\Sigma_{B}A^{B}_{t}Y^{B}_{t}\}=\{Y^{-}_{t}\}+\{A^{+}_{t}Y^{\Delta}_{t}\}$, given twice-differentiable coefficients $\{A^{+}_{t}(\pi^{+}_{t})\}$, must have the canonical form.
Under reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$, with $d(\mathbf{W}^{T}_{-})_{t}=\sigma^{Z}_{t}dw^{Z}_{t}+\sigma^{D}_{t}dw^{D}_{t}$, $\{dw^{Z}_{t}\}$ and $\{dw^{Z}_{t}\}$ being independent standard Wiener processes (Item-2&4, Section 1.1), we have:
$dS^{(0|A)}_{t}=\big{[}($ $\mathbf{1}_{+}-A^{+}_{t})r^{Z\Delta}_{t}dt\text{ }+\text{ }\sigma^{Z}_{t}dw^{Z}_{t}\big{]}\text{ }+\text{ }$ (B.1)
$(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)\partial_{\pi}A^{+}_{t}\cdot\bigg{[}($ $\mathbf{1}_{+}-\pi^{+}_{t})(\sigma^{l}_{t})^{2}dt+(\sigma^{\pi}_{t})^{2}\frac{\partial_{\pi}^{2}A^{+}_{t}}{2\partial_{\pi}A^{+}_{t}}(\sigma^{l}_{t})^{2}dt\text{ }+\text{ }\sigma^{l}_{t}dw^{\prime}_{t}\bigg{]},$ (B.2)
with model-drift $r^{Z\Delta}_{t}dt=-{d}Y^{\Delta}(t)$ and $(\sigma^{l}_{t})^{2}:=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$, where $\sigma^{lZ}_{t}:={r^{Z\Delta}_{t}}/{\sigma^{Z}_{t}}$ (’signal-to-noise’ ratio) stems from data $\{Z_{t}(B)\}$ (Peskir and Shirayev (2006)), and $\sigma^{lD}_{t}$ likewise, from $B$-conditionally independent data $\{D_{t}(B)\}$, and so $\sigma^{l}_{t}dw^{\prime}_{t}:=\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}$.
Without risk-pricing, $\{A^{+}_{t}\}=\{\pi^{+}_{t}\}$, we have trivial viability under reference belief $\pi^{B}_{0}\mathbf{W}^{T}_{B}$:
$dS^{(0|\pi)}_{t}=\big{[}(\mathbf{1}_{+}-\pi^{+}_{t})r^{Z\Delta}_{t}dt+\sigma^{Z}_{t}dw^{Z}_{t}\big{]}+(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)\big{[}(\mathbf{1}_{+}-\pi^{+}_{t})(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw^{\prime}_{t}\big{]}.$ (B.3)
With viable model-risk only pricing, under some RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}\sim\pi^{B}_{0}\mathbf{W}_{B}^{T}$: we have $\{S^{(0|A)}_{t}\}=\{\Sigma_{B}\hat{\pi}^{B}_{t}\hat{S}^{B}_{t}\}$ and the following relationship, by (8-11) and (19),
$\pi^{+}_{t}-A^{+}_{t}=k^{{A}}_{t}\sigma^{\pi}_{t}=k^{\hat{\pi}}_{t}\sigma^{\pi}_{t}+\frac{\Sigma_{B}\hat{\pi}^{B}_{t}\hat{RP}^{B}_{t}}{{Y}^{\Delta}(t)}=(\pi^{+}_{t}-\hat{\pi}^{+}_{t})+\frac{\Sigma_{B}\hat{\pi}^{B}_{t}\hat{RP}^{B}_{t}}{{Y}^{\Delta}(t)}.$ (B.4)
Remark 15.
Reference model-risk beliefs $\{{\pi}^{B}_{t}\}$ and any RNE $\{\hat{\pi}^{B}_{t}\}$ are ’adjacent’ (Item-5, Appendix A), implying that at any fixed $t<T$ for $\sigma^{\pi}_{t}\ll\frac{1}{2}$, by (23-24) when $\{{\hat{\pi}}^{B}_{t}\}=\{\Pi^{B}_{t}\}$, we have $\mathcal{O}({\pi}^{B}_{t}-{\hat{\pi}}^{B}_{t})=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and so $\mathcal{O}({k}^{\hat{\pi}}_{t})=\mathcal{O}(\sigma^{\pi}_{t})$ for any RNE price-of-model-risk ((11)). In turn, by Lemma 1, to balance the LHS and RHS of (B.4), we have: $\mathcal{O}({\hat{RP}^{B}_{t}})=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and $\mathcal{O}({k}^{A}_{t})=\mathcal{O}(\sigma^{\pi}_{t})$, as well as $\mathcal{O}(\partial_{\pi}A^{B}_{t})=\mathcal{O}(\partial^{2}_{\pi}A^{B}_{t})=\mathcal{O}(1)$ and $\mathcal{O}(d({\hat{RP}^{B}_{t})})=\mathcal{O}(d(\sigma^{\pi}_{t})^{2})=\mathcal{O}((\sigma^{\pi}_{t})^{4})$.
Consider first the case of trivial $\{D_{t}\}$. The drift of $\{dS^{(0|A)}_{t}\}$ is nil in expectation under some RNE measure $\hat{\pi}^{B}_{0}\hat{W}_{B}^{Z}|_{T}$, where $d(\hat{W}^{Z}_{B})_{t}=\hat{r}_{B,\text{ }t}dt+\sigma^{Z}_{t}dw^{Z}_{t}$, $\hat{r}_{B,\text{ }t}dt=-d(\hat{RP}^{B}_{t})$, $t<T$.
Being the RNE measure, the $(\hat{\pi}^{B}_{0}\hat{W}_{B}^{Z}|_{T})$-expectation of (B.1) must offset that of (B.2). Write that of (B.1) as $\nu_{t}:=(\hat{\pi}^{+}_{t}-A^{+}_{t})r^{Z\Delta}_{t}-\Sigma_{B}\hat{\pi}^{B}_{t}\hat{r}_{B,\text{ }t}$ and note $\hat{\pi}^{+}_{t}-\pi^{+}_{t}=(\hat{\pi}^{+}_{t}-A^{+}_{t})-(\pi^{+}_{t}-A^{+}_{t})$, viability thus requires:
$\frac{-\nu_{t}}{r^{Z\Delta}_{t}}=(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)\partial_{\pi}A^{+}_{t}\cdot\frac{(\sigma^{lZ}_{t})^{2}}{r^{Z\Delta}_{t}}\bigg{[}(\sigma^{\pi}_{t})^{2}\frac{\partial_{\pi}^{2}A^{+}_{t}}{2\partial_{\pi}A^{+}_{t}}-({\pi}^{+}_{t}-A^{+}_{t})-\frac{-\nu_{t}}{r^{Z\Delta}_{t}}\bigg{]}.$ (B.5)
By Remark 15, we have $\mathcal{O}(RHS)=\mathcal{O}((\sigma^{\pi}_{t})^{4})$ and $\mathcal{O}(LHS)=\mathcal{O}((\sigma^{\pi}_{t})^{2})$, $\forall\sigma^{\pi}_{t}\ll\frac{1}{2}$, so nil-drift is possible only if each side of the above vanishes on its own, meaning:
$\frac{\partial_{\pi}^{2}A^{+}_{t}}{2\partial_{\pi}A^{+}_{t}}=\frac{{\pi}^{+}_{t}-A^{+}_{t}}{(\sigma^{\pi}_{t})^{2}}.$ (B.6)
This has solutions of the form $\{A^{+}_{t}\}=\{\Pi^{+}_{t}\}$ only, where $\{\Pi^{+}_{t}\}$ is any model-risk inference based on any measure $\Pi^{B}_{0}W_{B}^{Z}|_{T}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$. The same can be obtained from the LHS of (B.5).
Adding $B$-conditionally independent data $\{D_{t}\}$ under law $W_{B}^{D}|_{T}$, the argument above applies unchanged to the dynamics of $\{dS^{(0|A)}_{t}\}$, (B.1-B.2), under reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$ and some RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}$ such that (B.4) holds and $d(\hat{\mathbf{W}}_{B}^{T})_{t}=\hat{r}_{B,\text{ }t}dt+d(\mathbf{W}_{-}^{T})_{t}$, $\hat{{r}}_{B,\text{ }t}dt=-d(\hat{{RP}}^{B}_{t})$. To conclude, the RNE measure $\{S^{(0|A)}_{t}\}$ must have the canonical form: $\Pi^{B}_{0}\mathbf{W}_{B}^{T}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.
Part II. Viability of canonical normal pricing $\{S^{(RP|{\Pi})}_{t}\}=\{{S}^{(0|{\Pi})}_{t}\}-\{\Sigma_{B}{RP}^{B}(t)\Pi^{B}_{t}\}$.
Under the martingale measure $\Pi^{B}_{0}\mathbf{W}_{B}^{T}$ of process $\{{S}^{(0|{\Pi})}_{t}\}$, process $\{S^{(RP|\Pi)}_{t}\}$ is a martingale of form (B.3) plus the extra drift $R^{B}_{t}dt=-{d}(RP^{B}(t))$ that stems from $B$-sure risk-pricing:
$dS^{(RP|\Pi)}_{t}=\big{[}(\mathbf{1}_{+}-\Pi^{+}_{t})\check{r}^{Z\Delta}_{t}dt+\sigma^{Z}_{t}dw^{Z}_{t}\big{]}+S^{\Delta}(t)(\sigma^{\Pi}_{t})^{2}\big{[}(\mathbf{1}_{+}-\Pi^{+}_{t})(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw^{\prime}_{t}\big{]}+R^{(\cdot)}_{t}dt,$ (B.7)
where $\check{r}^{Z\Delta}_{t}dt:=-{d}{}S^{\Delta}(t)$, with $S^{\Delta}(t)\equiv Y^{\Delta}(t)-RP^{+}(t)+RP^{-}(t)$. As such, it is viable so long as this extra drift $\{R^{B}_{t}dt\}$ can be ’absorbed’ into the known $(\Pi^{B}_{0}\mathbf{W}_{B}^{T})$-martingale $\{d{S}^{(0|{\Pi})}_{t}\}$, that is, whenever
$\frac{{R^{\pm}_{t}}/{S^{\Delta}(t)}}{{\sigma^{Z}_{t}}/{S^{\Delta}(t)}+{(\sigma^{\Pi}_{t})^{2}{\sigma^{lD}_{t}}\sqrt{1+\text{F}^{Z/D}_{t}}}}$
is square-integrable up to horizon $T$ (Novikov’s Condition), where, for convenience, we have written $(\sigma^{l}_{t})^{2}=(\sigma^{lD}_{t})^{2}(1+\text{F}^{Z/D}_{t})$, with $\text{F}^{Z/D}_{t}:=(\sigma^{lZ}_{t})^{2}/(\sigma^{lD}_{t})^{2}$.
C THE MOMENTUM AND LOW-RISK FORMULAE
For the assets of Section 3.2-3.4, all the distributions required stem from the underlying log-LR process $\{l^{1/{\text{{\bf{0}}}}}_{t}\}$, of parameter $\sigma^{l}$, for which log-LR levels $l^{1/{\text{{\bf{0}}}}}_{t}$ follow the Normal distribution:
$\mathcal{N}\bigg{(}\mu^{l}_{[t]}(B)=\frac{(-1)^{B+1}}{2}(\sigma^{l})^{2}t,\text{ }(\sigma^{l}_{[t]})^{2}=(\sigma^{l})^{2}t\bigg{)},\text{ }B\in\{0,1\}.$ (C.1)
The distributions of RNE beliefs $\{\Pi^{1}_{t}\}$ and so of prices ((38)) then derive via change-of-variable:
$l=H_{\Pi\pm}+\log\circ\text{ }O_{f}[v],$ (C.2)
$H_{\Pi\pm}:=\log\circ\text{ }O_{f}[\underline{\Pi}^{1}_{0\pm}]=H_{\mathbf{p}}+\log(\rho K^{\pm 1}),$ (C.3)
$H_{\mathbf{p}}:=\log\circ\text{ }O_{f}[\underline{\mathbf{p}}^{1}_{0}],$ (C.4)
for $\Pi^{1}_{t}=v$ given $l^{1/{\text{{\bf{0}}}}}_{t}=l$, with $dl=(v\underline{v})^{-1}dv$ ((3)) and $O_{f}[\Pi^{1}_{0\pm}]=O_{f}[\mathbf{p}^{1}_{0}]/(\rho K^{\pm 1})$ ((39)) given $sign[1{\text{{\bf{0}}}}]=\pm$. The $H$-variables above are inferential milestones: $H_{\Pi\pm}$ is the log-LR hurdle for event $\{\Pi^{1}_{t}\geq\frac{1}{2}\}$, and $H_{\mathbf{p}}$, for $\{\mathbf{p}^{1}_{t}\geq\frac{1}{2}\}$.
C.1 Tracking Inferential Progress and Bias Dominance
The degree of $B$-certainty as data accumulate may be assessed in the usual way given Normal distributions. The objective degree is tracked by some $|C^{\mathbf{p}}_{t}|$, where, with $t_{\mathbf{p}}:=2H_{\mathbf{p}}/(\sigma^{l})^{2}$,
$C^{\mathbf{p}}_{t}(B):=\frac{H_{\mathbf{p}}-\mu^{l}_{[t]}(B)}{\sigma^{l}_{[t]}}=\frac{1}{2}\sigma^{l}_{[t]}\cdot\big{(}{(-1)^{B}}+\frac{t_{\mathbf{p}}}{t}\big{)}.$ (C.5)
Certainty is high for $\sigma^{l}_{[t]}\ll 1$ (little data) or $\sigma^{l}_{[t]}\gg 1$ (lots of data), with $|C^{\mathbf{p}}_{t}|$ bottoming at 00 if and when $H_{\mathbf{p}}=\mu^{l}_{[t]}(B)$, that is, if and when $t=(-1)^{B+1}t_{\mathbf{p}}$ has a solution.
For the price-implied RNE inference process $\{\Pi^{1}_{t}\}$, the same indicator, denoted $|C^{\Pi}_{t\pm}|$, reads:
$C^{\Pi}_{t\pm}(B):=$ $\frac{H_{\Pi\pm}-\mu^{l}_{[t]}(B)}{\sigma^{l}_{[t]}}=\frac{1}{2}\sigma^{l}_{[t]}\cdot\big{(}{(-1)^{B}}+\frac{t_{\Pi\pm}}{t}\big{)}$ (C.6)
$=$ $\frac{1}{2}\sigma^{l}_{[t]}\cdot(\frac{t_{\rho}}{t}\pm\frac{t_{K}}{t})+C^{\mathbf{p}}_{t}(B),$ (C.7)
[…]
Bias creates a ’burden of proof’ $t_{\rho}$, and $t_{\rho}/t$ tracks how it is overcome. We consider cases of positive objective hurdle $H_{\mathbf{p}}>0$ and bias domination $H_{\Pi\pm}-H_{\mathbf{p}}\gg 0$, so $t_{\mathbf{p}}>0$ and $t_{\rho}\gg t_{K}$, where ’objective burden’ $t_{\mathbf{p}}$ is presumed ’non-extreme’ for non-trivial inference/excess.
C.2 The Window of Opportunity
Excess-profit opportunities occur when data become sufficient for the conditional probabilities of change to be meaningful but insufficient for subjective inference to overcome bias:
$\big{\{}\frac{t_{\mathbf{p}}}{t}\ll 1\big{\}}\text{ }\bigcap\text{ }\big{\{}\frac{t_{\rho}}{t}\gg 1\big{\}};$ (C.8)
the first demand means data dominance over objective hurdle and the second, bias over data.
Remark 16.
Consider for example profitable events $\{O_{f}[{\Pi}^{1}_{t}]=\rho^{-1}\}$ or $\{O_{f}[{\Pi}^{1}_{t\pm u}]=\rho^{-1}\}$ ((44)), with $t$ fixed and inside the above window, and $u>0$ large enough for the second event to be outside. Their respective probability densities are $\propto(t)^{-1/2}e^{-|C^{\mathbf{p}}_{t}|^{2}/2}$ and $\propto(t\pm u)^{-1/2}e^{-|C^{\mathbf{p}}_{(t\pm u)}|^{2}/2}$. Thus, in-window events dominate, with a log-LR of $\propto\big{(}\log[1\pm u/t]+|C^{\mathbf{p}}_{(t\pm u)}|^{2}-|C^{\mathbf{p}}_{t}|^{2}\big{)}\gg 1$, which diverges as $(1-u/t)\to 0$ or $(1+u/t)\to\infty$, with $t$ fixed.
C.3 Momentum and Volatility Mixtures
The probabilities relevant to volatility-conditioned mean-excess (42) obey the rule:
$\mathbb{P}_{t}(\Pi^{1}_{t}=v)\mathbb{P}_{t}(\pm|\text{ }\Pi^{1}_{t}=v)\equiv\mathbb{P}_{t}(\pm)\mathbb{P}_{t}(\Pi^{1}_{t}=v|\text{ }\pm),$ (C.9)
with $\mathbb{P}_{t}(\Pi^{1}_{t}=v|\text{ }\pm)$ from Normal distribution (C.1) via (C.2). The likelihood-ratio of $\{\Pi^{1}_{t}=\underline{v}\}$ vs $\{\Pi^{1}_{t}=v\}$, $v\in(0,\frac{1}{2}]$, given the $sign[1{\text{{\bf{0}}}}]=\pm$ of potential change, is:
$R_{t|\pm}(\underline{v}/v)(B):=\frac{\mathbb{P}_{t}(\underline{v}|\pm)}{\mathbb{P}_{t}(v|\pm)}(B)=({\underline{v}}/{v})^{-(t_{\Pi\pm}/{t})-(-1)^{B}},\text{ }B\in\{0,1\}.$ (C.10)
Note that $\{\Pi^{1}_{t}=v\}$-events dominate as long as bias does (i.e. $t_{\rho}/{t}\gg 1$) regardless; its dependence on actual $B$-outcomes is weak.
To compute (42) we also need the mix-ratio function $M_{t|v}:={\mathbb{P}_{t}(+|v)}/{\mathbb{P}_{t}(-|v)}$ of the event-set $\{\Pi^{1}_{t}=v\}$, $v\in(0,1)$, between member-events with $\{sign[1{\text{{\bf{0}}}}]=+\}$ and with $\{sign[1{\text{{\bf{0}}}}]=-\}$.
By (C.9) we have $M_{t|v}={\mathbb{P}_{t}(v|+)}/{\mathbb{P}_{t}(v|-)}$, since $\mathbb{P}_{t}(\pm)=\frac{1}{2}$ by setup. Hence:
$M_{t|v}(B)$ $=(\frac{\rho v}{\underline{v}})^{-t_{K}/t}\cdot K^{-(t_{\mathbf{p}}/t)-(-1)^{B}},$ (C.11)
and likewise the volatility-conditioned ratio $M_{t|v\underline{v}}:={\mathbb{P}_{t}(+|v\underline{v})}/{\mathbb{P}_{t}(-|v\underline{v})},\text{ }v\in(0,\frac{1}{2}]$:
$M_{t|v\underline{v}}=M_{t|v}\frac{1+R_{t|+}(\underline{v}/v)}{1+R_{t|-}(\underline{v}/v)}.$ (C.12)
Note the source of conditionality: risk-pricing $K$; if $K=1$, conditional mixes equal the unconditional mix; if $K\in(1,2)$, they are perturbations of the unconditional one.
Remark 17.
Both mix-ratios (C.11-C.12) are monotone declining, and peak-volatility ($v=\frac{1}{2}$) brings $M_{t|\frac{1}{2}}=M_{t|\frac{1}{2}\frac{1}{2}}\approx\rho^{-t_{K}/t}=\exp[{-(t_{\rho}/t)\log K}]$. That is, for $\rho\gg 1$, in the window (C.8) of opportunity, when trading events are most abundant, the observed mix at high volatility can be highly negative vs the unconditional level. Mix returns to its unconditional state as $t\to\infty$ when model-risk resolves. Any uneven unconditional background merely introduces a constant factor in (C.11-C.12); all above statements remain valid.
C.4 Peak-Reward Location and Size for Volatility-Conditioned Trading
Focusing first on the $\{\Pi^{1}_{t}=v\}$-contributions to volatility-conditioned mean-excess (42), given mix-function (C.11) and condition (C.8), it has a unique optimum, (44), with $\mathcal{O}(\frac{(K-1)H_{\mathbf{p}}}{(\sigma^{l})^{2}t})$-errors; the accuracy of solution (44) improves with data accumulation.
Further, given (C.10) and under condition (C.8), the $\{\Pi^{1}_{t}=\underline{v}\}$-contributions to (42) are no more than $\mathcal{O}(\rho^{-t_{\rho}/{t}})$. That is, the leading order solution (44) is not affected by these contributions while $t_{\rho}/t\gg 1$, i.e. while bias dominates.
Remark 18.
For uneven unconditional-mix, $\mathbb{P}_{t}(\pm)\not=\frac{1}{2}$, the same applies. Compared to even solution (44), its peak moves left (right), and size, down (up), for unevenness that tilts negatively (positively); peak-location is capped by $v^{+mo}_{max}$ ((43)), and peak-size, by $rp^{+mo}_{max}$ ((43)).
Without uncertainty or bias, $\rho=1$, both the $\{\Pi^{1}_{t}=v\}$- and $\{\Pi^{1}_{t}=\underline{v}\}$-part of (42) are invariant under $v\leftrightarrow\underline{v}$ switching ((41) and (C.10-C.12)), so their derivatives both vanish at $v=\frac{1}{2}$, thus confirming ’high risk high reward’ under classical conditions.

### v6

@@
 Existing approaches to asset-pricing under model-uncertainty adapt classical utility-maximisation frameworks and seek theoretical comprehensiveness.
-We move toward practice by considering binary model-uncertainties and by switching attention from ’preference’ to ’constraints’.
-This decomposes economic asset-pricing into the viable pricing of model-risk and non-model risk separately such that the former has a unique and intuitive formula with convenient properties.
-Its parameter, dynamically conserved under model-risk inference, allows an integrated representation of ex-ante risk-pricing and bias, with ex-post price-effects that can be disentangled, through well-known anomalies, Momentum and Low-Risk, whose risk-reward curves acquire a new significance: peak-rewards measure ex-ante risk-pricing, and peak-locations, bias.
+We move toward practice by considering binary model-risks and by emphasizing ’constraints’ over ’preference’.
+This decomposes viable economic asset-pricing into that of model and non-model risks separately, leading to a unique and convenient model-risk pricing formula.
+Its parameter, a dynamically conserved constant of model-risk inference, allows an integrated representation of ex-ante risk-pricing and bias such that their ex-post impacts are disentangled via well-known anomalies, Momentum and Low-Risk, whose risk-reward patterns acquire a fresh significance: peak-reward reveals ex-ante risk-premia, and peak-location, bias.
 JEL Classification: C11, C12, C22, D81, G02, G12
@@
 (1991)), with a prominent role in Knightian Decision Theory (Bewley (2002)), making our framework a tool for the exploration of these concepts in practical and concrete terms.
-Existing works adapting classical theories to incorporate model-risk come under ambiguity-aversion or robust-control, where agents perform recursive optimisation, against goals such as max-min utility, optimal smooth-ambiguity or variational preference, over the set of model choices (e.g.
+Existing works adapting classical theories to incorporate model-risk come under the heading of ambiguity-aversion or robust-control, where agents perform recursive optimisation, against goals such as max-min utility, optimal smooth-ambiguity or variational preference, over the set of model choices (e.g.
 Ju and Miao (2012) or Machina and Siniscalchi (2014)).
-Related is parameter learning (e.g.
+Related also is parameter learning (e.g.
 Guidolin and Timmermann (2007), Collin-Dufresne et al.
-(2016)), carried out usually under a definitive model-risk law, without uncertainty.
-Existing focus is the nature of preference/utility under model-risk and model-risk discount in a general equilibrium.
-Our focus is how model-risk pricing may be achieved or characterised ’minimally’, with fewest assumptions and parameters.
+(2016)), often performed under some known model-risk law.
+Existing focus is the nature of preference/utility under model-risk and the resulting general equilibrium.
+We focus here instead on how model-risk pricing might be done or characterised ’minimally’, in terms of assumptions and parameters.
 A natural start is thus the Risk-Neutral Equivalent (RNE) formulation, given the First Theorem of Asset Pricing (FTAP) [^1] Early max-min utility struggled with FTAP-compliance (e.g.
@@
 (2021))..
-However, viability, No Arbitrage in the FTAP sense, is too weak for most pricing tasks: there are far more viable processes than economically valid ones.
-Moreover, learning about model-risk makes RNE probabilities path-dependent, and so intractable, even in simple cases (proved formally here, but already observed in Guidolin and Timmermann (2007)).
-One way to address such obstacles is to add incremental conditions to the FTAP baseline till RNE formulations that do provide economic asset-pricing can be characterised.
-It inverts the routine of ’theory making followed by viability checking’, and switches attention from ’agent preference’ to ’structural constraints’.
-It culminates in the decomposition of viable economic asset-pricing into separately viable components and the identification of a unique and intuitive RNE model-risk pricing formula, whose sole parameter, a constant of motion of the model-risk inferential dynamics, allows an integrated treatment of ex-ante risk-premia and bias such that their ex-post price-effects, only jointly observable, can be disentangled.
+The RNE approach to model-risk has been seldom discussed, for two good reasons.
+Firstly, viability, No Arbitrage in the FTAP sense, is too weak for most purposes: there are far more viable price processes than economically valid ones.
+Secondly, parameter learning makes RNE probabilities path-dependent and intractable, even in simple cases, as observed by Guidolin and Timmermann (2007).
+We overcome these difficulties by tracing their origins and examining the consequences of the constraints they impose on asset-pricing.
+The key turns out to be whether model-risk pricing is a standard Ito process of the model-risk inference process, or equivalently, whether the economic price-of-model-risk has an identical RNE form.
+It leads to a unique and intuitive model-risk pricing formula, whose parameter, a constant of motion of model-risk inference, allows an integrated treatment of ex-ante risk-premia and bias such that their ex-post effects, only jointly observable, may be disentangled.
+The RNE approach can apply at the aggregate or the individual-asset level, thereby potentially expanding the data universe helpful to and addressable by model-risk/parameter-learning studies, which are currently closed to stock-level or cross-sectional data.
 Scope and Premise
-We take the existence of representative agents/beliefs for granted, and use the language of rational (Bayesian) inference to mean no more than adherence to the laws of probabilities; likewise, ’rational’ and ’expectation’ signify the same without implying ’objective correctness’, unlike in classical Rational Expectation (RE) frameworks.
-Three types of probabilities are thus involved: 1) objective ones governing events; 2) subjective ones representing inferential beliefs about the objective world; 3) RNE ones for pricing.
-Let our model-risk $B$, of micro or macro economic origin, be binary, with outcome $b$ or $\overline{b}\not=b$, and our $B$-sure models, elementary in the sense of being model-risk free.
-Such $B$-sure processes may represent the familiar, which by classical RE principles must be well-modelled already.
-In our context therefore, only change, breaks from routines, may create model-risk (e.g.
-new CEO, R&D results, or regulatory/governmental/geopolitical events).
-Many real-world asset valuation challenges fall into this category.
-Generalisation to model-risks that are non-binary is outlined towards the end, by involving composite $B$-sure processes, which contain discounts for other model-risks.
+The existence of representative agents/beliefs is taken for granted, so the language of rational (Bayesian) inference is used, to mean no more than adherence to the laws of probabilities; likewise, ’rational’ and ’expectation’ signify the same without implying ’objective truthfulness’, unlike in classical Rational Expectation (RE) frameworks.
+We thus work with three types of probabilities: 1) the objectively true ones governing events; 2) the subjective ones representing inferential beliefs about the objective world; and 3) the RNE ones for pricing.
+Let our model-risk $B$ have two outcomes, $b$ or $\overline{b}\not=b$, and our $B$-sure models, elementary (’model-risk free’).
+Such $B$-sure processes represent ’routines’ that by classical RE principles are well-modelled; only change creates model-risk (e.g.
+new CEO, R&D results, macroeconomic, regulatory or geopolitical events).
+Non-binary situations that can be handled within the binary framework emerge quite naturally, as will be demonstrated in an application addressing typical questions encountered in practice.
 Remark 1.
@@
 Our results apply only to the economically priced: any deemed worthy of being assigned at each moment a risk-premium through some recursive optimisation scheme taking said risk as an input.
-Our model-risk shall be considered ’one-off’.
+Our model-risk shall be considered one-off.
 How it ’flares up’ and the pricing of ’flare-risk’ remains implicit; that is, we study episodic ’model crises’, one at a time.
@@
 Correspondingly, let horizons be sufficiently long for the meaningful development of model-risk inference, and let asset-pricing be time-homogeneous (’identical risk-levels identically priced’) during such developments (i.e.
-while the model-risk remains active and relevant).
+while the given model-risk remains active and relevant).
 The law of our model-risk, the unconditional distribution of $B$-outcomes, may be unknown.
-If the true unconditional probability, denoted $\mathbf{p}^{B}_{0}$ throughout, is known, model-detection (learning) and asset-pricing remain classical, with familiar conclusions.
+If the objectively true unconditional probability, denoted $\mathbf{p}^{B}_{0}$ throughout, is known, model-detection (learning) and asset-pricing remain classical, with familiar conclusions.
 If unknown, the same may involve non-classical cognitive and economic elements.
 Nonetheless, asset-pricing may be formulated in terms of classical inference, with non-classical aspects reflected through preference and discount mechanisms; this is the usual, and our, strategy.
-Lastly, data about model-risks (e.g.
-weekly/monthly indicators of economic regimes) tend to be more frequent and impactful than data about ’routine’ risks (e.g.
+Lastly, data on model-risk (e.g.
+weekly/monthly indicators of economic regimes) tend to be more frequent and impactful than data on ’routine’ risks (e.g.
 biannual dividend).
@@
 The setup as outlined above makes model-risk a prominent feature of asset-pricing, a major source of excess value, and so a focus of market effort and activities.
-Existing Theories in Our Context
+From the Perspectives of Standard Model-Risk Studies
 Most model-risk pricing schemes, ambiguity-aversion, robust-control and some parameter-learning, achieve model-risk discounts by in effect pessimistically distorting a reference belief (about model-risk), known as ’reference measure’, ’second-order prior’, or ’benchmark’.
@@
 Remark 2.
-Call the set $\mathcal{L}_{\mathcal{S}}:=\{\pi^{I}_{0}Q_{I}|\pi^{I}_{0}\in\mathcal{I}_{2}(\mathcal{I}_{1}(\mathcal{S})),\text{ }I\in\mathcal{I}_{1}(\mathcal{S})\}$ of reference beliefs about the law governing total state space $\mathcal{I}_{1}(\mathcal{S})\times\mathcal{S}$ the set of model-uncertainty, for later reference.
-The max-min utility approach offers the simplest strategy: choose at any time the model with the worst outlook from the set $\mathcal{I}_{1}(\mathcal{S})$ of alternatives (in our setting it means pricing for the pessimistic $B$-outcome till $B$-risk resolves).
-It is a limit of the graded optimisations adopted by the other approaches, where pessimism is tempered by reference beliefs (e.g.
-Hansen and Sargent (2011) penalises distorted beliefs by their relative entropy to the reference).
+The set $\mathcal{L}_{\mathcal{S}}:=\{\pi^{I}_{0}Q_{I}|\pi^{I}_{0}\in\mathcal{I}_{2}(\mathcal{I}_{1}(\mathcal{S})),\text{ }I\in\mathcal{I}_{1}(\mathcal{S})\}$ of reference beliefs about the law of total state space $\mathcal{I}_{1}(\mathcal{S})\times\mathcal{S}$ may be called ’the set of model-uncertainty’, for convenience.
+The max-min utility approach offers the simplest solution: choose at any time the model with the worst outlook from the set $\mathcal{I}_{1}(\mathcal{S})$ of alternatives (meaning in our setting choosing the pessimistic $B$-outcome till $B$-risk resolves).
+It is a limit of the graded optimisations used by others, where pessimism is tempered by reference beliefs (e.g.
+Hansen and Sargent (2011) penalises distorted beliefs by their relative entropy to a chosen benchmark).
 Key to all is then the reference belief, by which model-risk pricing is defined and computed.
@@
 Under model-uncertainty, even if the reference beliefs happen to be true objectively, perceived uncertainty still triggers extra discounts.
-All worked examples in the uncertainty literature presume the reference beliefs used by their agents to be unknowingly objectively true; this in our context is unjustifiable.
+All worked examples in the ambiguity-aversion/robust-control literature presume the reference beliefs of their agents to be unknowingly objectively true; this in our context is hard to justify.
 At the other extreme, some parameter-learning stances see realised parameter-value as deterministic, meaning $\mathbf{p}^{B}_{0}\equiv 1\text{ or }0$ if adopted here; this would exclude many types of model-risks of practical interest.
@@
 Any assertion of equality would mask a basic issue facing all attempts to free classical RE frameworks from any built-in coincidence of beliefs and true laws: pricing theory outputs are ex-ante but what can be observed is ex-post, so tests and calibrations are impossible without additional theories of bias.
-Manifestations of this difficulty include ’observational equivalence’ (’is the large ex-post risk-premium due to bias after all?’; see e.g.
+Manifestations of this ’impasse’ include ’observational equivalence’ (’is the high realised risk-premium observed due to bias after all?’; see e.g.
 Cecchetti et al.
@@
 Fama (2014)).
-We will illustrate how this issue resolves itself in our approach.
-Finally, existing theories require as inputs the total risk as a whole, model and non-model risks together, aggregated across all the assets in a given market.
-This makes them unwieldy for bottom-up or cross-sectional enquiries, especially when the model-risk(s) can be micro, macro or of a variety of causes.
-Our method, with its isolation of model-risk pricing and RNE formulation, may prove more flexible and convenient in this context.
-It also reveals that unless a model-risk is ’consistent’, no interpretable economic model-risk pricing is possible.
+We show how this issue may resolve itself in our approach.
+Finally, existing theories require ’total risk’ as a whole as inputs, model and non-model risks together, aggregated across all relevant assets in a given market.
+They are unwieldy for bottom-up (individual-asset) or cross-sectional enquiries, especially when the model-risk(s) can be micro-economic or heterogeneous in origin.
+Our method, with its isolation of model-risk pricing and RNE formulation, is more flexible and convenient for such applications.
 Content and Organisation
-Section 1 sets out the usual asset and data structure in continuous time, highlighting the effect of model-risk on standard risk-pricing considerations; Section 2 derives RNE asset-pricing under model-risk; Section 3 demonstrates a potential implication, followed by conclusions and discussions; technical backgrounds and proofs are collected in the Appendix.
-It transpires that the RNE formulations of asset-pricing under model-risk are constrained by informational properties (Definition 1, Lemma 1 and Remark 14) and so must follow a canonical form (Proposition 1), with familiar properties (Corollary 1) and price dynamics (Section 2.2.2).
-One consequence is a potential link between Status Quo Bias and well-known market-anomalies, whose characteristics separate prospective (ex-ante) market risk-pricing from bias (if any) under model-uncertainty, revealing them separately (Section 3.5.1&3.5.2).
+Section 1 sets out the usual asset and data structure in continuous time, highlighting the effect of model-risk on standard risk-pricing considerations; Section 2 derives RNE asset-pricing under model-risk; Section 3 demonstrates a potential application, followed by conclusions and discussions; technical backgrounds and proofs are collected in the Appendix.
+We find that the RNE formulations of asset-pricing under model-risk are constrained by economic, informational, necessities (Definition 1 and Lemma 1) and so have a canonical form (Proposition 1), resulting in familiar properties (Corollary 1) and dynamics (Section 2.2.2).
+One implication is a potential link between Status Quo Bias and well-known anomalies, whose characteristic risk-reward curves reveal and separate ex-ante risk-premia from bias under model-uncertainty (Section 3.4.1&3.4.2).
 1 Setup
-1.1 Usual-Case Assets and Information Structure
-For clarity and to avoid clutter, consider assets with a single payoff, some bullet-payoff $Y\in\mathbb{R}$ at horizon (or maturity) $T$, with $1\ll T\leq\infty$; let value and pricing be in logarithm, and time- or risk-free discount be exogenous and nil (with a suitable choice of numéraire).
+1.1 The Usual-Case Asset and Information Structure
+For clarity, consider an usual-case asset: it has just one, bullet, payoff, $Y\in\mathbb{R}$, at some horizon or maturity $T$ much longer than unit-time, so $1\ll T<\infty$; let value and price be in logarithm, and time- and risk-free discount exogenous and nil (in a suitable numéraire).
 Further:
   1. 1.
-    Any incremental change to bullet-payoff $Y$, if at all, is public, such that the cumulative firmed-up (’cum-div’) component $\{Z_{t}\}$ of asset-worth, with $Z_{T}\equiv Y$ by definition, has the equation of motion below, subject to binary outcome $B\in\mathcal{B}:=\{+,-\}$, but otherwise based on standard Wiener process $\{w^{Z}_{t}\}$, of natural, usual, filtration $\{\mathcal{F}^{Z}_{t}\}$,
+    Any incremental change (if at all) to bullet-payoff $Y$ is such that the cumulative firmed-up (’cum-div’) asset-worth $\{Z_{t}\}$, with $Z_{T}\equiv Y$, follows the dynamics below, based on a standard Wiener process $\{w^{Z}_{t}\}$, with natural filtration $\{\mathcal{F}^{Z}_{t}\}$, subject to binary outcomes $\{B=b\}$ (e.g.
+’high’ or ’low’) labelled by $b\in\mathcal{B}:=\{+,-\}$,
     $dZ_{t}(B)=r^{Z}_{B,t}dt+\sigma^{Z}_{t}dw^{Z}_{t},\text{ }t\in(0,T],$ (1)
-    with $\{\mathcal{F}^{Z}_{t}\}$-adapted volatility $\{\sigma^{Z}_{t}\}$ and drift $\{r^{Z}_{B,t}\}$, on probability space $(\mathcal{S}^{Z}|_{T},W^{Z}_{B}|_{T})$, which is the restriction to $(0,T]$ of the usual Wiener space/measure $(\mathcal{S}^{Z},W^{Z})$ generated by $\{w^{Z}_{t}\}$, with the $B$-sure law ${W}^{Z}_{B}$ of $\{Z_{t}(B)\}$ defined via $d(W^{Z}_{-})_{t}=r^{Z}_{-,t}+\sigma^{Z}_{t}dw^{Z}_{t}$ for $B=-$, and $d(W^{Z}_{+})_{t}=r^{Z}_{+,t}dt+\sigma^{Z}_{t}dw^{Z}_{t}$ for $B=+$; write $r^{Z\Delta}_{t}:=r^{Z}_{+,t}-r^{Z}_{-,t}$.
+    where its $B$-sure drifts $\{r^{Z}_{B,t}\}$ and $B$-independent volatilities $\{\sigma^{Z}_{t}\}$ are both $\{\mathcal{F}^{Z}_{t}\}$-predictable.
+The underlying $B$-sure probability space $(\mathcal{S}^{Z}|_{T},W^{Z}_{B}|_{T})$ is the restriction to $(0,T]$ of the Wiener space/measure $(\mathcal{S}^{Z},W^{Z}_{B})$ associated with process $\{Z_{t}(B)\}$, that is, $dW^{Z}_{B,t}=dZ_{t}(B)$.
+The above allows the sequential testing of Wiener processes of differing drifts [^2] Text-book sequential tests (e.g.
+Peskir and Shirayev (2006)) concern processes of fixed drift and volatility.
+Our setup allows time-variations achieved through an absolutely continuous clock-change, or equivalently, $\{\mathcal{F}^{Z}_{t}\}$-predictable drifts and volatilities.
+For more background see Appendix A..
   2. 2.
-    There may be an additional $B$-informative real-vector dataflow $\{{D}_{t}(B)\}$, of similar yet $B$-conditionally independent dynamics to $\{Z_{t}(B)\}$:
-    $dD_{t}(B)=r^{D}_{B,t}dt+\sigma^{D}_{t}dw^{D}_{t},$ (2)
-    all as in (1) but with ’$D$’ labelling; its independence makes it price-irrelevant without model-risk (i.e.
-under known $B$).
-Combined dataflow $\{\mathbf{D}_{t}\}:=\{(Z_{t},D_{t})\}$ has state space $\mathcal{S}^{T}:=(\mathcal{S}^{Z}\times\mathcal{S}^{D})|_{T}$ and natural filtration $\{\mathbf{F}_{t}\}:=\{\mathcal{F}^{Z}_{t}\}\vee\{\mathcal{F}^{D}_{t}\}$, under product $B$-sure law $\mathbf{W}^{T}_{B}:=\mathbf{W}_{B}|_{T}$, where $\mathbf{W}_{B}:=({W}^{Z}_{B}\times{W}^{D}_{B})$.
+    The economic impact of model-risk $B$ thus derives from drift difference $r^{Z\Delta}_{t}:=r^{Z}_{+,t}-r^{Z}_{-,t}$; let it be consistent: $Y^{\Delta}(t):=\int_{t}^{T}r^{Z\Delta}_{s}ds>0$ almost surely, at any $t<T$.
+This is intuitive, also necessary for viable and economically valid risk-pricing, as will be seen (Remark 8).
+  3. 3.
+    There may be a purely $B$-informative real-vector dataflow $\{{D}_{t}(B)\}$ that is $B$-conditionally independent but otherwise similar to $\{Z_{t}(B)\}$, based on a standard Wiener process $\{w^{D}_{t}\}$,
+    $dD_{t}(B)=r^{D}_{B,t}dt+\sigma^{D}_{t}dw^{D}_{t}.$ (2)
+    It by definition is price-irrelevant in the absence of model-risk (when $B$-sure).
+Combined, total dataflow $\{\mathbf{D}_{t}\}:=\{(Z_{t},D_{t})\}$ has state space $\mathcal{S}^{T}:=(\mathcal{S}^{Z}\times\mathcal{S}^{D})|_{T}$ and natural filtration $\{\mathbf{F}_{t}\}:=\{\mathcal{F}^{Z}_{t}\}\vee\{\mathcal{F}^{D}_{t}\}$, under the product $B$-sure law $\mathbf{W}^{T}_{B}:=\mathbf{W}_{B}|_{T}$, with $\mathbf{W}_{B}:={W}^{Z}_{B}\times{W}^{D}_{B}$.
     Remark 3.
-    Current treatments ignore the make up of data $\{\mathbf{D}_{t}\}$, as if only those relevant to model-sure pricing, $\{Z_{t}\}$, exist/matter.
-We consider situations with data that are exclusive to model-risk (price-irrelevant once model-sure), including cases with trivial model-sure pricing data $\{Z_{t}\}$.
-It tackles a frequent challenge to partitioners: a crisis market unsettled by a bout of model-risk whose economics and inferential status dominate, while non-model, routine, risks remain, independently evolving or unchanging.
-  3. 3.
-    Given the above, continuous (predictable) inferential testing of model-risk $B$ takes place on filtered space $(\mathcal{S}^{T}_{\mathcal{B}}:=\mathcal{B}\times\mathcal{S}^{T},\text{ }\{\mathbf{F}_{t}\},\text{ }\mathbf{F}_{T};\text{ }{\pi}^{B(\cdot)}_{0}\mathbf{W}^{T}_{B(\cdot)}(\cdot))$, where ${\pi}^{B(\omega_{B})}_{0}$ is the unconditional belief in outcome $B(\omega_{B})\in\mathcal{B}$, $\omega_{B}:=(B,\omega)\in\mathcal{S}^{T}_{\mathcal{B}}$.
-Under the condition of $\mathbf{W}_{+}\perp\mathbf{W}_{-}$ with $\mathbf{W}_{+}|_{t}\sim\mathbf{W}_{-}|_{t}$, $\forall t<\infty$, standard testing thus resolves $B$-outcomes at guaranteed resolution-time $T(B)=\inf\big{\{}t|\mathbf{F}_{t}\ni\{B=\cdot\}\big{\}}=\infty$.
-    Remark 4.
-    The sequential testing of Wiener processes differing by a fixed drift is well-known.
-Our setup allows time-variations via absolutely continuous clock-change, making finite resolutions ($T(B)<\infty$) possible.
-Such issues are largely moot under economic risk-pricing (Remark 1), with $T<T(B)\leq\infty$ (see Appendix A for more).
+    Existing studies ignore potential structures within dataflow ($\{\mathbf{D}_{t}\}$), as if only data relevant to tangible, firmed-up value, $\{Z_{t}\}$, exist/matter.
+By considering data that are exclusive to model-risk (price-irrelevant once model-sure), including when tangible-value data, $\{Z_{t}\}$, are absent or trivial, we tackle a frequent challenge facing partitioners: a crisis market unsettled by a bout of model-risk whose economics dominate, while non-model, routine, risks remain, independently evolving or unchanging.
   4. 4.
-    Such testing, optimal in the sense of Neyman-Pearson Lemma, is based on likelihood-ratios (LR) $\{L^{+/-}_{t}:=\frac{d\mathbf{W}^{T}_{+}|_{t}}{d\mathbf{W}^{T}_{-}|_{t}}\}$, or equivalently, log-LR $\{l^{+/-}_{t}:=\log L^{+/-}_{t}\}$, which form Wiener processes under broad conditions; model-risk $B$ resolves iff.
-$\lim_{t\to T(B)}l^{+/-}_{t}=(-1)^{\mathbf{1}_{-}(B)}\cdot\infty$ (iff.
-$\mathbf{W}_{+}|_{T(B)}\perp\mathbf{W}_{-}|_{T(B)}$) for some $T(B)\leq\infty$.
-Further, standard testing is performed under the equivalent martingale measure of one of the processes (e.g that of $B=-$); henceforth let $\mathbf{W}_{-}$ be by default this base measure, under which $r^{Z}_{-,t}=0=r^{D}_{-,t}$, $\forall t$.
+[…]
   5. 5.
-    The above has Bayesian representations: any $L^{+/-}_{t}\in(0,\infty)$ maps bijectively and smoothly to an a posteriori belief ${\pi}^{B}_{t}\in(0,1)$.
-Bayes’ Rule, in terms of odds-for, $O_{f}[{\pi}^{B}_{t}]:={\pi}^{B}_{t}/\underline{\pi}^{B}_{t}$, with ()¯$\underline{(\cdot)}=1-(\cdot)$, in effect says ’new-odds $=$ old-odds $\times$ likelihood-ratios of data’:
+    Tests based on likelihood-ratios (LR) $L^{+/-}_{t}:=\frac{d\mathbf{W}^{T}_{+}|_{t}}{d\mathbf{W}^{T}_{-}|_{t}}$, or log-LR $l^{+/-}_{t}:=\log L^{+/-}_{t}$, $\forall t<T$, are optimal in the sense of Neyman-Pearson Lemma.
+Model-risk $B$ resolves iff.
+$\exists T_{\mathcal{B}}\leq\infty$ such that $\lim_{t\to T_{\mathcal{B}}}l^{+/-}_{t}=(-1)^{\mathbf{1}_{-}(B)}\cdot\infty$ (iff.
+$\mathbf{W}_{+}|_{T_{\mathcal{B}}}\perp\mathbf{W}_{-}|_{T_{\mathcal{B}}}$).
+Further, such inferential tests are performed under the equivalent martingale measure of one of the processes (e.g that of $B=$-).
+Henceforth let $\mathbf{W}_{-}$ be this base measure, under which $r^{Z}_{-,t}=0=r^{D}_{-,t}$, $\forall t$.
+  6. 6.
+    The above has a Bayesian representation: any $L^{+/-}_{t}\in(0,\infty)$ maps bijectively and smoothly to an a posteriori belief ${\pi}^{B}_{t}\in(0,1)$.
+Bayes’ Rule, in terms of odds-for, $O_{f}[{\pi}^{B}_{t}]:={\pi}^{B}_{t}/\underline{\pi}^{B}_{t}$, with $\underline{(\cdot)}=1-(\cdot)$, says in effect ’new-odds == old-odds $\times$ likelihood-ratios of data’:
     $O_{f}[{\pi}^{+}_{t}]=O_{f}[{\pi}^{+}_{0}]\times\exp[{l^{+/-}_{t}}]\propto{L^{+/-}_{t}}\in(0,\infty),$ (3)
-    given a priori odds $O_{f}[{\pi}^{+}_{0}]\in(0,\infty)$; so inferential odds are geometric Wiener processes.
-  6. 6.
-    The time-$t$, $B$-sure, expected payoff $Y^{B}_{t}:={\mathbf{E}}_{B}[Y|\mathbf{F}_{t}]$ under its $B$-sure law reads ((1)):
-    $Y^{B}_{t}=z_{t}+\mathbf{E}_{B}[Y-z_{t}|{\mathbf{F}}_{t}]=z_{t}+\mathbf{E}_{B}[Y-z_{t}]=z_{t}+y_{B}(T-t,t),$ (4)
+    given a priori odds $O_{f}[{\pi}^{+}_{0}]\in(0,\infty)$.
+That is, inferential odds are exponentiated log-LR processes; the latter, $\{l^{+/-}_{t}\}$, are Wiener processes under broad conditions ((A.2-A.3)).
+  7. 7.
+    Given $Z_{t}=z_{t}$, $t<T$, the $B$-sure expected payoff $Y^{B}_{t}:={\mathbf{E}}_{B}[Y|\mathbf{F}_{t}]\equiv{\mathbf{E}}_{B}[Y|\mathcal{F}^{Z}_{t}]$ reads ((1-2)):
+    $Y^{B}_{t}=z_{t}+\mathbf{E}_{B}[Y-z_{t}|{\mathcal{F}^{Z}_{t}}]=z_{t}+\mathbf{E}_{B}[Y-z_{t}]=z_{t}+y_{B}(T-t,t),$ (4)
+    where $y_{B}(T-t,t)-y_{B}(T-t-t^{\prime},t+t^{\prime})=\mathbf{E}_{B}[\int_{t}^{t+t^{\prime}}dZ_{s}(B)]$, $\forall t^{\prime}\leq T-t$, and so $y_{B}(T-t,t)=\int_{t}^{T}{r}^{Z}_{B,s}ds$, as by definition $\{Y^{B}_{t}\}$ is a martingale and $y_{B}(0,T)\equiv 0$.
+Note also that this relates to model-risk impact $Y^{\Delta}(t)$ via $Y^{\Delta}(t)\equiv Y^{+}_{t}-Y^{-}_{t}\equiv{y}_{+}(T-t,t)-{y}_{-}(T-t,t)$.
+1.2 The FTAP Baseline under Model-Risk
 […]
-  7. 7.
-[…]
-1.2 The FTAP Baseline under Model-Risk
-Translated to our setup, FTAP states: market $\big{(}Y,\{S_{t}\};\text{ }\mathcal{S}^{T}_{\mathcal{B}},\{\mathbf{F}_{t}\},\{\mathbf{F}_{T}\};\text{ }\pi^{B}_{0}\mathbf{W}_{B}^{T}\big{)}$, with $\{\mathbf{F}_{t}\}$-adapted asset-prices $\{S_{t}\},\text{ }S_{T}\equiv Y$, is viable iff.
-there is an equivalent measure $\widehat{\pi^{B}_{0}\mathbf{W}^{T}_{B}}\sim{\pi}^{B}_{0}\mathbf{W}_{B}^{T}$, the RNE measure, such that $S_{t}=\hat{\mathbf{E}}[Y\big{|}\mathbf{F}_{t}]$, $\forall t<T$, where RNE expectation $\hat{\mathbf{E}}[\cdot{|}\cdot]$ may be written as:
+Such FTAP-viable asset-pricing under model-risk $B$ may be written always as follows:
 $S_{t}=\Sigma_{B}\hat{\pi}^{B}_{t}{\hat{S}^{B}_{t}}:=\hat{\pi}^{+}_{t}\hat{S}^{+}_{t}+\hat{\pi}^{-}_{t}\hat{S}^{-}_{t},$ (5)
 $\hat{S}^{B}_{t}:=\hat{\mathbf{E}}_{B}[Y\big{|}\mathbf{F}_{t}]:=\hat{\mathbf{E}}\big{[}Y\big{|}B,\mathbf{F}_{t}\big{]},$ (6)
-with $\Sigma_{B}{(\cdot)^{B}}{(\cdot)^{B}}$ denoting convex combinations, such as the one under RNE beliefs $\hat{\pi}^{B}_{t}$, as calculated under Bayes’ Rule (3) given $\hat{\pi}^{B}_{0}$ and RNE log-LR $\hat{l}^{+/-}_{t}:=\log\frac{d\hat{\mathbf{W}}^{T}_{+}|_{t}}{d\hat{\mathbf{W}}^{T}_{-}|_{t}}$.
-This market, defined by (1-6), is complete (Martingale Representation Theorem), and so the RNE measure, unique.
+where $\Sigma_{B}{(\cdot)^{B}}{(\cdot)^{B}}$ denotes averages weighted by $B$-outcomes, such as by RNE beliefs $\hat{\pi}^{B}_{t}$ under Bayes’ Rule (3) given $\hat{\pi}^{B}_{0}$ and RNE log-LR $\hat{l}^{+/-}_{t}:=\log\frac{d\hat{\mathbf{W}}^{T}_{+}|_{t}}{d\hat{\mathbf{W}}^{T}_{-}|_{t}}$.
 Viability by no means ensures economic validity (e.g.
-any pricing based on RNE measures of the form $\tilde{\pi}^{B}_{0}\tilde{\mathbf{W}}^{T}\sim{\pi}^{B}_{0}\mathbf{W}_{B}^{T}$ is blind to $B$-information); indeed, any pricing that is viable to any member of the equivalence class $[{\pi}^{B}_{0}\mathbf{W}_{B}^{T}]$ is viable to all in the class.
-That is, under FTAP-viability, ’anything goes’ (almost).
-What then characterises the processes (and the RNE measures) that do correspond to plausible economic theories?
-It turns out that having economically interpretable RNE price-of-model-risk is the key.
-Consider risk-premium, the gap between ex-ante price and expectation: at any $t<T$, given reference model-risk belief $\{{\pi}^{B}_{t}\}$ ((3)),
-$RP_{t}:=Y_{t}-S_{t}:=\Sigma_{B}{\pi}^{B}_{t}Y^{B}_{t}-S_{t},$ (7)
-and its RNE decomposition (by inserting ±$\pm\Sigma_{B}\hat{\pi}^{B}_{t}{{Y}^{B}_{t}}$):
+any pricing based on RNE measures of the form $\tilde{\pi}^{B}_{0}\tilde{\mathbf{W}}^{T}\sim{\pi}^{B}_{0}\mathbf{W}_{B}^{T}$ is blind to $B$-informative data); indeed, any pricing that is viable to any member of the equivalence class ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}]$ is viable to all in the class.
+In this sense, ’anything goes’ under FTAP-viability.
+What characterises the processes and the RNE measures that do correspond to plausible economic theories?
+For answers, consider risk-premium, the gap between ex-ante pricing and expectation, non-negative under economic asset-pricing: at $\forall t<T$, given reference model-risk belief $\{{\pi}^{B}_{t}\}$,
+$RP_{t}:=Y_{t}-S_{t}=\Sigma_{B}{\pi}^{B}_{t}Y^{B}_{t}-S_{t}\geq 0.$ (7)
+It has RNE decomposition (by inserting $\pm\Sigma_{B}\hat{\pi}^{B}_{t}{{Y}^{B}_{t}}$):
 $RP_{t}=\Sigma_{B}\hat{\pi}^{B}_{t}{\hat{RP}^{B}_{t}}+B\_\widehat{RP}_{t},$ (8)
@@
 ${k}^{\hat{\pi}}_{t}:=\frac{B\_\widehat{RP}_{t}}{Y^{\Delta}(t)\sigma^{\pi}_{t}}\equiv\frac{{\pi}^{+}_{t}-\hat{\pi}^{+}_{t}}{(\pi^{+}_{t}\underline{\pi}^{+}_{t})^{\frac{1}{2}}}.$ (11)
+Decomposition (8-11) exists whenever the RNE measure exists, that is, for all viable pricing schemes.
+However, its elements, $\{\hat{S}^{B}_{t}\}$ and $\{{k}^{\hat{\pi}}_{t}\}$, need not correspond to economically interpretable entities.
+It turns out that establishing such a correspondence, albeit indirectly, is the key to finding economically valid asset-prices (out of mere viable ones).
 1.3 The Economic Decomposition of Risk-Pricing under Model-Risk
-Regardless of the economic pricing scheme under model-risk, classical or otherwise, the resulting asset-price $\{S_{t}\}$, with risk-premia $\{RP_{t}\}$ and RNE formulation (6-11), admits an economic decomposition of the form (16-19), provided the existence of 1) some reference model-risk beliefs $\{{\pi}^{B}_{t}\}$ and 2) $B$-sure pricing $\{S^{B}_{t}\}$ under the same scheme when $B$-sure:
+Just as any viable price process admits a RNE decomposition of the embedded risk-premia, any economic pricing scheme under model-risk, classical or otherwise, with asset-price $\{S_{t}\}$, risk-premia $\{RP_{t}\}$ and RNE formulation (6-11), admits an economic decomposition, (16-19).
+It is unique and well-defined provided the existence of 1) a reference model-risk belief process $\{{\pi}^{B}_{t}\}$; 2) viable $B$-sure pricing $\{S^{B}_{t}\}$ produced under the same scheme when $B$-sure:
 $S^{B}_{t}=\check{\mathbf{E}}_{{B}}[Y{|}{\mathbf{F}}_{t}]=z_{t}+\check{\mathbf{E}}_{B}[Y-z_{t}]=z_{t}+\check{y}_{B}(T-t,t),$ (12)
-$RP^{B}_{t}:=y_{B}(T-t,t)-\check{y}_{B}(T-t,t)=:RP^{B}(t)\geq 0,$ (13)
+$RP^{B}_{t}:=Y^{B}_{t}-S^{B}_{t}=y_{B}(T-t,t)-\check{y}_{B}(T-t,t)=:RP^{B}(t)\geq 0,$ (13)
+where $\check{\mathbf{E}}_{{B}}[\cdot|\cdot]$ denotes expectations under its $B$-sure RNE measure $\check{\mathbf{W}}^{T}_{B}\sim{\mathbf{W}}^{T}_{B}$, leading to RNE outlook $\check{y}_{B}(T-t,t)=\int_{t}^{T}\check{r}_{B,s}ds$ as in (4), and $B$-sure risk-premium $RP^{B}(t)$ satisfying: $\forall t\leq T$,
+$\check{y}_{+}(T-t,t)-\check{y}_{-}(T-t,t)=:S^{\Delta}(t)>0,\text{ almost surely},$ (14)
+$S^{B}_{t}=\lim_{\sigma^{\pi}_{t}\to 0}\hat{S}^{B}_{t},\text{ }\text{i.e. }RP^{B}(t)=\lim_{\sigma^{\pi}_{t}\to 0}\hat{RP}^{B}(t),$ (15)
+where (14) ensures consistency with asset-property Item-2, Section 1.1, while (15), the internal consistency of the pricing scheme under model-risk $B$.
+Remark 4.
+Model-sure usual-case pricing is simple: static noise compensated by a predictable drift.
+Adding model-risk, the only new input to economic pricing (recursive optimisation taking drivers of the expected preference-value as inputs) is the reference model-risk belief $\{{\pi}^{B}_{t}\}$.
 […]
-$\check{y}_{+}(T-t,t)-\check{y}_{-}(T-t,t)=:S^{\Delta}(t)>0,\text{almost surely},$ (14)
-$RP^{B}(t)=\lim_{\sigma^{\pi}_{t}\to 0}\hat{RP}^{B}(t),\text{ }\text{i.e. }S^{B}_{t}=\lim_{\sigma^{\pi}_{t}\to 0}\hat{S}^{B}_{t},$ (15)
-where (14) ensures consistency with asset-property Item-7, Section 1.1, and (15), internal consistency of the pricing scheme under model-risk.
-Remark 5.
-Usual-case model-sure pricing is simple: static noise discounted by a deterministic drift.
-With model-risk $B$, the only new input to economic pricing (recursive optimisation taking drivers of the expected preference-value as inputs) is the reference model-risk belief $\{{\pi}^{B}_{t}\}$.
-At any $t<T$, viable pricing must satisfy $S_{t}\in(S^{-}_{t},S^{+}_{t})$, so there always exists a well-defined pricing coefficient ${A}^{B}_{t}\in(0,1)$ such that:
 $S_{t}=\Sigma_{B}{A}^{B}_{t}S^{B}_{t}=S^{-}_{t}+A^{+}_{t}S^{\Delta}(t),$ (16)
@@
 $k^{A}_{t}=\frac{{\pi}^{+}_{t}-A^{+}_{t}}{(\pi^{+}_{t}\underline{\pi}^{+}_{t})^{\frac{1}{2}}}\geq 0,$ (19)
-where the economic price-of-$B$-risk $\{k^{A}_{t}\}$ should be free from explicit dependence on $B$-sure pricing, the latter being based on decisions in the absence of $B$-risk.
-As such, once obtained, price-of-$B$-risk $\{k^{A}_{t}\}$ may generate the coefficients $\{A^{B}_{t}\}$ via (19) for asset-pricing under $B$-risk, given consistent but otherwise arbitrary $B$-sure prices $\{{S^{\prime}}^{B}_{t}\}:=\{Y^{B}_{t}\}-\{RP^{\prime B}(t)\}$, thus separating the pricing of model and non-model risks.
-This is attractive if for a useful range of risk-premia, the prices so constructed:
-$\{S^{\prime}_{t}\}\equiv\{S^{(RP^{\prime}|A)}_{t}\}:=\{\Sigma_{B}A^{B}_{t}{S^{\prime}}^{B}_{t}\}=\{S^{(0|A)}_{t}\}-\text{ }\{\Sigma_{B}A^{B}_{t}{RP^{\prime}}^{B}(t)\},\text{ with }$ (20)
-$\{S^{(0|A)}_{t}\}:=\{\Sigma_{B}A^{B}_{t}Y^{B}_{t}\}=\{Y^{-}_{t}\}+\{A^{+}_{t}Y^{\Delta}(t)\}=\{Y_{t}\}-\text{ }\{k^{A}_{t}\sigma^{\pi}_{t}\cdot{Y}^{\Delta}(t)\},$ (21)
-are viable; both are one way or another the sum of a martingale and a drift.
-Observe that all applied economic asset-pricing theories, continuous and regular near risk-neutrality, are already viable written in above form, at least for sufficiently small risk-discounts.
+which parallels RNE decomposition (8-11), but pricing coefficients $\{A^{B}_{t}\}\in(0,1)$, unlike $\{\hat{\pi}^{B}_{t}\}$, need not be a conditional-probability process.
+We show shortly that they in fact must be.
+The economic price-of-$B$-risk $\{k^{A}_{t}\}$ should have no explicit dependencies on $B$-sure prices, the latter being decisions in the absence of $B$-risk.
+Thus, once devised, process $\{k^{A}_{t}\}$ can generate asset-prices under $B$-risk via (19) and (16) given any consistent but otherwise arbitrary $B$-sure pricing $\{{S}^{B}_{t}\}\equiv\{Y^{B}_{t}\}-\{RP^{B}(t)\}$ ((13)), thus separating the pricing of model and non-model risks.
+This is attractive if for a useful range of risk-discount choices, the asset-pricing so achieved,
+$\{S^{(RP|A)}_{t}\}:=\{\Sigma_{B}A^{B}_{t}{S}^{B}_{t}\}=\{S^{(0|A)}_{t}\}\text{ }-\text{ }\{\Sigma_{B}A^{B}_{t}{RP}^{B}(t)\},\text{ with }$ (20)
+$\{S^{(0|A)}_{t}\}:=\{\Sigma_{B}A^{B}_{t}Y^{B}_{t}\}=\{Y^{-}_{t}\}+\{A^{+}_{t}Y^{\Delta}(t)\}=\{Y_{t}\}\text{ }-\text{ }\{k^{A}_{t}\sigma^{\pi}_{t}\cdot{Y}^{\Delta}(t)\},$ (21)
+is viable.
+Observe that both above can be the sum of a martingale and a well-behaved drift, and that all applied economic pricing theories, continuous and regular near risk-neutrality, are already viable when written in this form at least for small enough risk-discounts.
 Remark 5.
-In ambiguity-aversion settings, the primary utility function, enforcing risk-aversion, can be dialled towards near neutrality (linearity) while secondary utility/variational-control continues to exert uncertainty discounts.
-In classical settings, including parameter-learning, where a single utility must be optimised, the same may be achieved by replacing the payoff variable with its up-to-date model-sure expectation perturbed by a small enough noise.
-The real question is then the range of viability for processes (20-21), and most relevant to us, if they have convenient RNE formulations.
-For this, we must characterise the economic price-of-$B$-risk $\{k^{A}_{t}\}$ ((19)) more fully.
-Its origin suggests the following.
+In ambiguity-aversion settings, the primary utility function, enforcing risk-aversion, and the secondary utility/variational-control, enforcing uncertainty-aversion, can be dialled independently and arbitrarily close to neutrality (linearity), to achieve the above separation.
+In classical and some parameter-learning settings, with a single utility, the above obtains by replacing payoffs with their model-sure expectations perturbed by suitably small noise.
+The real question is then the range of viability for (20-21) and if they have tractable RNE formulations. For this, let us characterise the economic price-of-$B$-risk $\{k^{A}_{t}\}$ better:
 Definition 1.
-Asset-pricing under model-risk $B$ is said to be normal if its price-of-model-risk $\{k^{A}_{t}\}$ at any $t\ll T$: 1) depends deterministically on reference model-risk belief $\{\pi^{B}_{t}\}$ in a time-homogenous manner [^2] Economic risk-pricing, solution to some optimisation-based pricing scheme, is measurable to the natural filtration of its input, ($(\{Z_{t}\},\{\pi^{B}_{t}(\mathbf{D}_{t})\})$, and so a function of its level to date (Proposition 4.9 of Breiman (1992)).
-Any explicit dependence on time $t$ or $Z_{t}$ itself, if at all, is presumed weak or non-existent, at least for $t\ll T$.
-such that the dependence is twice-differentiable and non-negative; 2) satisfies $\lim_{\sigma^{\pi}_{t}\to 0}k^{A}_{t}=0$; 3) generates viable pricing via (20-21) under an economically relevant range of risk-premium choices.
+Asset-pricing under model-risk $B$ is said to be usual if its price-of-$B$-risk $k^{A}_{t}$, (19), is a time-homogenous twice-differentiable function of its model-risk belief [^3] Economic risk-pricing, as the solution to some optimisation scheme, is measurable to the natural filtration of its input $(\{Z_{t}\},\{\pi^{B}_{t}(\mathbf{D}_{t})\})$; so it is at any time $t$ a function of $(Z_{t},\pi^{B}_{t}(\mathbf{D}_{t}))$ (Proposition 4.9 of Breiman (1992)).
+Any explicit dependence on $t$ or $Z_{t}$, if at all, is presumed here weak or non-existent, at least for $t\ll T$.
+$\pi^{B}_{t}$, $\forall t\ll T$.
 Remark 6.
-Property-1&2 are shared by most if not all existing model-risk pricing theories. Property-3 will be shown to hold for usual-case assets under model-risk.
+This makes $\{k^{A}_{t}\}$ a time-homogeneous Ito process of risk-inference $\{\pi^{B}_{t}\}$, a common feature of economic risk-pricing theories: priced risks are captured as state-variables so that the resulting risk-pricing is their function and hence path-independent with respect to them.
+It will be seen (via Proposition 1 and Corollary 1) that this is responsible for all the ’nice’ properties of $\{k^{A}_{t}\}$: fixed-signed, with $\lim_{\sigma^{\pi}_{t}\to 0}k^{A}_{t}=0$ (ensuring (15)), and generating prices (20-21) that follow familiar dynamics and allow a wide range of viable risk-premium choices.
 2 Canonical RNE Asset-Pricing under Model-Risk
-2.1 Viable Normal Model-Risk Pricing and its RNE measure
+2.1 Viable, Usual, Model-Risk Pricing
 Proposition 1.
-For usual-case assets under model-risk $B$, with reference model-risk process $\{\pi^{B}_{t}\}$ based on some reference belief $\pi^{B}_{0}\mathbf{W}^{T}_{B}\in\mathcal{L}_{\mathcal{S}}$ from the set of model-uncertainty, normal pricing (Definition 1) is viable iff.
-it has this canonical form: at any $t<T$,
+Under model-risk $B$ and model-risk inference $\{\pi^{B}_{t}\}$ based on a reference belief $\pi^{B}_{0}\mathbf{W}^{T}_{B}\in\mathcal{L}_{\mathcal{S}}$ that belongs to the set $\mathcal{L}_{\mathcal{S}}$ of model-uncertainty (Remark 2), any usual pricing (Definition 1) of any usual-case asset is viable iff.
+it has the canonical form below: $\forall t<T$,
 $S_{t}=S^{(RP|{\Pi})}_{t}=\Sigma_{B}\Pi^{B}_{t}S^{B}_{t}=S^{(0|{\Pi})}_{t}-\Sigma_{B}{RP}^{B}(t)\Pi^{B}_{t},$ (22)
-where model-risk only pricing $\{S^{(0|{\Pi})}_{t}\}=\{\Sigma_{B}\Pi^{B}_{t}Y^{B}_{t}\}$ ((21)) is given by some RNE model-risk inference process $\{\Pi^{B}_{t}\}$ based on a RNE measure of the form $\Pi^{B}_{0}\mathbf{W}^{T}_{B}\in\mathcal{L}_{\mathcal{S}}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.
+where model-risk only asset-pricing $S^{(0|{\Pi})}_{t}=\Sigma_{B}\Pi^{B}_{t}Y^{B}_{t}$ (see (21)) is given by a RNE model-risk inference process $\{\Pi^{B}_{t}\}$ under a RNE measure of the form $\Pi^{B}_{0}\mathbf{W}^{T}_{B}\in\mathcal{L}_{\mathcal{S}}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.
 Proof.
-See Appendix B.
-Briefly, any viable model-risk only pricing $\{S^{(0|A)}_{t}\}=\{Y^{-}_{t}\}+\{A^{+}_{t}Y^{\Delta}(t)\}$ must have drift $\{Y^{\Delta}(t)dA^{+}_{t}\}$ offset $\{dY^{-}_{t}+A^{+}_{t}dY^{\Delta}(t)\}$ in expectation under some RNE measure, possible iff.
-it belongs to $\mathcal{L}_{\mathcal{S}}$, under which the two drift separately vanish in expectation, due to $\{A^{+}_{t}\}$ being a twice-differentiable function of model-risk inference $\{\pi^{+}_{t}\}$ driven partly/wholly by $\{dY^{-}_{t}\}$.
-The viability of (22), as a martingale plus a well-behaved drift, follows.
-Canonical pricing thus allows the pricing of model and non-model risks to have separate RNE formulations that are each easy to identify and apply.
-Note however that the overall RNE measure of viable normal asset-pricing (22) under model-risk $B$, although guaranteed to exist, remains unidentifiable in general, even for simple cases.
+Appendix B.
+Briefly, viable pricing $S^{(0|A)}_{t}=Y^{-}_{t}+A^{+}_{t}Y^{\Delta}(t)$ must see $dY^{-}_{t}+A^{+}_{t}dY^{\Delta}(t)$ offset $Y^{\Delta}(t)dA^{+}_{t}$ in expectation under a RNE measure, possible iff.
+the two expectations vanish separately due to $\{A^{+}_{t}\}$ being an Ito process of $\{\pi^{+}_{t}\}$, thus limiting the RNE to the form stated.
+The viability of (22) then follows from the ’good behaviour’ of $\{\Sigma_{B}{RP}^{B}(t)\Pi^{B}_{t}\}$
+Canonical model-risk only pricing $\{S^{(0|{\Pi})}_{t}\}$, with $\{RP^{B}_{t}\}=0$, has identical economic and RNE price-of-model-risk ((19) and (11)); indeed any pricing with this property must be model-risk only (with $\{RP^{B}_{t}\}=0$) [^4] Any such pricing must have its RNE measure in $\mathcal{L}_{\mathcal{S}}$ and so have the form $\{\Sigma_{B}A^{B}_{t}S^{B}_{t}\}=\{\Sigma_{B}\Pi^{B}_{t}Y^{B}_{t}\}$.
+Hence, $\forall t<T$, $\lim_{\sigma^{\pi}_{t}\to 0}S_{t}=Y^{B}_{t}$ since $\lim_{\sigma^{\pi}_{t}\to 0}\sigma^{\Pi}_{t}=0$, but $\lim_{\sigma^{\pi}_{t}\to 0}{S}_{t}=S^{B}_{t}=Y^{B}_{t}-RP^{B}_{t}$ given (15); thus $\{RP^{B}_{t}\}=0$..
+Given Definition 1 and Remark 6, this means that viable prices (16) that have non-trivial $B$-sure risk-pricing (with $\{RP^{B}_{t}\}\not=0$) must have a RNE price-of-model-risk that is path-dependent with respect to any model-risk process $\{\pi^{B}_{t}\}$ derived from the set $\mathcal{L}_{\mathcal{S}}$ of model-uncertainty.
+This formalises the observation of Guidolin and Timmermann (2007) that learning makes RNE probabilities path-dependent and so intractable.
+The source of this difficulty is basically (15) (see footnote-4), given Lemma 1 and the fact that the operation of $B$-conditioning and of risk-neutralising (reference beliefs such as $\pi^{B}_{0}\mathbf{W}^{T}_{B}$) do not commute in general.
+Canonical pricing (22) does not remove this intrinsic issue but avoids it by separating the RNE pricing of model and non-model risks, each by itself easy and tractable.
+The total RNE measure of (22), though guaranteed, remains intractable.
+2.2 Properties of Canonical Model-Risk Pricing
+2.2.1 The 1-Parameter Price-of-Model-Risk Formula
+Corollary 1.
+Canonical price-of-model-risk $\{k^{\Pi}_{t}\}:=\{\frac{{\pi}^{+}_{t}-\Pi^{+}_{t}}{\sigma^{\pi}_{t}}\}$ (see (19) or (11)) satisfies:
+$k^{\Pi}_{t}=(K^{\frac{1}{2}}-K^{-\frac{1}{2}})\sigma^{\Pi}_{t},\text{ }\forall t<\infty,$ (23)
+with $\sigma^{\Pi}_{t}:=({\Pi}^{+}_{t}\underline{\Pi}^{+}_{t})^{\frac{1}{2}}$ and $K:=\frac{O_{f}[{\pi}^{+}_{0}]}{O_{f}[{\Pi}^{+}_{0}]}=\frac{O_{f}[{\pi}^{+}_{t}]}{O_{f}[{\Pi}^{+}_{t}]}$, a conserved quantity.
 Remark 7.
-Note that canonical model-risk only pricing $\{S^{(0|{\Pi})}_{t}\}$ has identical RNE and economic price-of-model-risk ((11) and (19)).
-This indeed exhausts the set of price processes with this property (Lemma 1 and Remark 14, Appendix A).
-In fact any process with non-trivial risk-pricing of the non-model risks (i.e.
-$\{RP^{B}_{t}\}\not=0$) must have an RNE price-of-model-risk that is not a time-homogeneous deterministic function of any reference model-risk belief based on the set $\mathcal{L}_{\mathcal{S}}$ of model-uncertainty [^3] Any viable price process whose RNE measure is in $\mathcal{L}_{\mathcal{S}}$ must have the form $\{\Sigma_{B}A^{B}_{t}S^{B}_{t}\}=\{\Sigma_{B}\Pi^{B}_{t}Y^{B}_{t}\}$, so $\forall t<T$, $\lim_{\sigma^{\pi}_{t}\to 0}S_{t}=Y^{B}_{t}$ since $\lim_{\sigma^{\pi}_{t}\to 0}\sigma^{\Pi}_{t}=0$, but $\lim_{\sigma^{\pi}_{t}\to 0}{S}_{t}=S^{B}_{t}=Y^{B}_{t}-RP^{B}_{t}$ given (15); thus $\{RP^{B}_{t}\}=0$..
-This echos the observation that parameter-learning makes RNE formulations path-dependent (Guidolin and Timmermann (2007)).
-2.2 Properties of Canonical Model-Risk Pricing
-2.2.1 The 1-Parameter Price-of-Model-Risk
-Corollary 1.
-Canonical model-risk premium $\{B\_RP_{t}\}$ and price-of-model-risk $\{k^{\Pi}_{t}\}$ are as follows:
-$B\_RP_{t}$ $=({\pi}^{+}_{t}-{\Pi}^{+}_{t}){S}^{\Delta}(t),\text{ }t<T,$ (23)
-$k^{\Pi}_{t}$ $=(K^{\frac{1}{2}}-K^{-\frac{1}{2}})\sigma^{\Pi}_{t},\text{ }t<T,$ (24)
-with $\sigma^{\Pi}_{t}:=({\Pi}^{+}_{t}\underline{\Pi}^{+}_{t})^{\frac{1}{2}}$, and $K:=\frac{O_{f}[{\pi}^{+}_{0}]}{O_{f}[{\Pi}^{+}_{0}]}=\frac{O_{f}[{\pi}^{+}_{t}]}{O_{f}[{\Pi}^{+}_{t}]}$, a conserved quantity of risk-inference ((3)).
+This follows from Bayes’ Rule (3).
+Price-of-model-risk $k^{\Pi}_{t}$ is positive iff.
+$K>1$.
+It takes value $k_{\frac{1}{2}}:=(K-1)/(K+1)$ at peak risk $\sigma^{\pi}_{t}=\frac{1}{2}$, where model-risk pricing offers a risk-premium of $\frac{1}{2}k_{\frac{1}{2}}{S}^{\Delta}(t)$ (see (18)) and a gain-to-loss ratio [^5] Gain-to-loss with respect to $B$-outcomes is always $O_{f}[{\Pi}^{+}_{t}]^{-1}$, but $O_{f}[{\pi}^{b}_{t}]=1$ at peak $B$-risk, hence the result.
+of $(1+k_{\frac{1}{2}})/(1-k_{\frac{1}{2}})\equiv K$.
+To any buyers (sellers), the higher (lower) $K$ is, the better; it is natural in theory and practice to set $K\in(1,2)$ for ’competitively priced’ model-risk, with $K-1\not\gg 1$ in any case.
 Remark 8.
-[…]
-Remark 10.
-The sign of the difference process $\{{\pi}^{+}_{t}\}-\{{\Pi}^{+}_{t}\}$ is fixed ((3)); so unless model-risk impact is consistent (Item-7, Section 1.1), model-risk premia $\{B\_RP_{t}\}$ must turn negative at times: inconsistent model-risks cannot have interpretable viable economic risk-premia.
-2.2.2 Connection with Familiar Price Dynamics
-[…]
-$S_{t}=S^{(\check{rp}|\Pi)}_{t}=S^{(0|\Pi)}_{t}-\check{rp}(t)=Y^{-}_{t}-\check{rp}(t)+S^{\Delta}(t)\Pi^{+}_{t},$ (25)
-$dS_{t}(B)=dY^{-}_{t}(B)+\check{R}_{t}dt-\Pi^{+}_{j}(B)r^{Z\Delta}_{t}dt+S^{\Delta}(t)d\Pi^{+}_{t}(B).$ (26)
-Given Section 1.1 and inference dynamics (A.2-A.3), we have $\{d\Pi^{+}_{t}\}$ as below, with two sources of noise, one from $\{Z_{t}\}$, given by $\{\sigma^{lZ}_{t}dw^{Z}_{t}\}$, with $\sigma^{lZ}_{t}:=r^{Z\Delta}_{t}/\sigma^{Z}_{t}$ (’signal-to-noise’ ratio), one from independent $B$-informative data $\{D_{t}\}$, modelled in parallel ((2)), which can be absorbed via $\{\sigma^{l}_{t}dw^{\prime}_{t}\}=\{\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}\}$, with $(\sigma^{l}_{t})^{2}=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$:
-$d\Pi^{+}_{t}(B)=\mu^{\Pi,B}_{t}+\sigma^{\Pi,l}_{t}dw^{\prime}_{t},\text{ with},$ (27)
-$\mu^{\Pi,B}_{t}:=\sigma^{\Pi,l}_{t}\sigma^{l}_{t}\Big{(}{{\mathbf{1}_{+}}}(B)-{\Pi}^{+}_{t}(B)\Big{)},$ (28)
+The difference process $\{{\pi}^{+}_{t}\}-\{{\Pi}^{+}_{t}\}=\{k^{\Pi}_{t}\sigma^{\pi}_{t}\}$ has a fixed sign.
+Unless the model-risk is consistent (see Item-2, Section 1.1, and (14)), viable model-risk premia $\{B\_RP_{t}\}$ (see (18)) must be negative sometimes and so correspond to no known economic risk-pricing theories.
+2.2.2 Connection with Familiar Price-Dynamics
+For better focus, let any $B$-sure risk-premia be $B$-independent henceforth, to be written as $\check{rp}(t)=\int_{t}^{T}\check{R}_{s}ds>0$, with $\check{R}_{t}$ denoting the $B$-sure risk-premium drift at any $t<T$; note then $Y^{\Delta}(t)=S^{\Delta}(t)=\int_{t}^{T}r^{Z\Delta}_{s}ds$ as a result.
+Canonical pricing (22) becomes:
+$S_{t}=S^{(0|\Pi)}_{t}-\check{rp}(t)=Y^{-}_{t}+S^{\Delta}(t)\Pi^{+}_{t}-\check{rp}(t),$ (24)
+$dS_{t}(B)=dY^{-}_{t}(B)+S^{\Delta}(t)d\Pi^{+}_{t}(B)-r^{Z\Delta}_{t}\Pi^{+}_{j}(B)dt+\check{R}_{t}dt.$ (25)
+By setup (Section 1.1) and inference dynamics (A.2-A.3), process $\{d\Pi^{+}_{t}\}$ has two sources of noise, one from $\{Z_{t}\}$, given by $\{\sigma^{lZ}_{t}dw^{Z}_{t}\}$, with $\sigma^{lZ}_{t}:=r^{Z\Delta}_{t}/\sigma^{Z}_{t}\ll 1$ (’signal-to-noise’), one from independent $B$-informative data $\{D_{t}\}$, modelled in parallel ((2)), which can be absorbed through $(\sigma^{l}_{t})^{2}:=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$ and $\sigma^{l}_{t}dw_{t}:=\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}$; then:
+$d\Pi^{+}_{t}(B)=\mu^{\Pi,B}_{t}dt+\sigma^{\Pi,l}_{t}dw_{t},\text{ with},$ (26)
+$\mu^{\Pi,B}_{t}:=\sigma^{\Pi,l}_{t}\sigma^{l}_{t}\cdot\big{(}{{\mathbf{1}_{+}}}(B)-{\Pi}^{+}_{t}(B)\big{)},$ (27)
 $\sigma^{\Pi,l}_{t}:=(\sigma^{\Pi}_{t})^{2}\sigma^{l}_{t}.$ (28)
-The above has a drift by design relative to reference belief $\{{\pi}^{+}_{t}\}$ (of dynamics (A.3)):
+The above has a drift by design relative to reference belief $\{{\pi}^{+}_{t}\}$ (with dynamics (A.3)):
 $\mu^{\Pi,\pi}_{t}:=\Sigma_{B}{\pi}^{B}_{t}{\mu^{\Pi,B}_{t}}=\sigma^{\Pi,l}_{t}\sigma^{l}_{t}\cdot(\pi^{+}_{t}-\Pi^{+}_{t})=\sigma^{\Pi,l}_{t}\sigma^{l}_{t}\cdot k^{\Pi}_{t}\sigma^{\pi}_{t}\geq 0;$ (29)
-note ${\mu^{\pi,\pi}_{t}}\equiv 0\equiv{\mu^{\Pi,\Pi}_{t}}$, confirming the martingale property of inference and model-risk pricing.
-Further, as inference volatility $\sigma^{\Pi,l}_{t}$ (29) drives price-volatility ((26-27)), the price-of-model-risk (24) peaks as model-risk induced price-volatility peaks.
+note ${\mu^{\pi,\pi}_{t}}\equiv 0\equiv{\mu^{\Pi,\Pi}_{t}}$, reflecting the martingale property of inference and viable pricing.
+Further, as inference volatility $\sigma^{\Pi,l}_{t}$ (28) drives price-volatility (see (25-26)), price-of-model-risk $\{k^{\Pi}_{t}\}$ (by (23)) peaks where price-volatility peaks, ceteris paribus.
 Pulling it together, the full asset-price dynamic, a $(\Pi^{B}_{0}\mathbf{W}^{T}_{B})$-martingale plus drift $\{\check{R}_{t}\}$, reads:
 $dS_{t}(B)=$ $\big{[}\check{R}_{t}dt+\sigma^{Z}_{t}dw^{Z}_{t}\big{]}$ (30)
-$+$ $r^{Z\Delta}_{t}dt\big{(}\mathbf{1}_{+}(B)-\Pi^{+}_{j}(B)\big{)}$ (32)
-$+$ $S^{\Delta}(t)(\sigma^{\Pi}_{t})^{2}\big{[}(\sigma^{l}_{t})^{2}\big{(}\mathbf{1}_{+}(B)-\Pi^{+}_{t}(B)\big{)}+\sigma^{l}_{t}dw^{\prime}_{t}\big{]}.$ (33)
-Under trivial $\{Z_{t}\}$ (firmed-up value process), the above becomes $dS_{t}(B)\propto d\Pi^{+}_{t}(B)$, making asset-pricing a pure matter of model-risk inference, which further reduces to the standard inferential test of Peskir and Shirayev (2006) under constant $\{r^{D\Delta}_{t}\}$ (signal) and $\{\sigma^{D}_{t}\}$ (noise).
+$+$ $\big{[}\big{(}\mathbf{1}_{+}(B)-\Pi^{+}_{j}(B)\big{)}\cdot r^{Z\Delta}_{t}dt\big{]}$ (31)
+$+$ $S^{\Delta}(t)(\sigma^{\Pi}_{t})^{2}\big{[}\big{(}\mathbf{1}_{+}(B)-\Pi^{+}_{t}(B)\big{)}\cdot(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw_{t}\big{]}.$ (32)
 Remark 9.
-Practical cases often see $\check{R}_{t}\ll r^{Z\Delta}_{t}\ll S^{\Delta}(t)$: the $B$-sure dynamic (31) dominated by model-drift (32) dominated by model-inference (33); dispersion among such assets is driven by model-risk inference.
-It often holds also that $\text{F}^{Z/D}_{t}:=(\sigma^{lZ}_{t})^{2}/(\sigma^{lD}_{t})^{2}\ll 1$, so $(\sigma^{l}_{t})^{2}\approx(\sigma^{lD}_{t})^{2}$: prices dynamics are driven by data other than those on tangible (firmed-up) values.
-Averaged under model-risk belief $\{{\pi}^{+}_{t}\}$, dominant price-dynamic (33) has a classical form:
-$\sigma_{t}\bigg{(}\frac{\mu_{t}}{\sigma_{t}}+dw^{\prime}_{t}\bigg{)},$ (34)
+Practical cases often see $B$-sure dynamics (30) and model-drift (31) dominated by model-inference (32): $\check{R}_{t}\ll S^{\Delta}(t)$ and $r^{Z\Delta}_{t}\ll S^{\Delta}(t)$; indeed firmed-up values $\{Z_{t}\}$ are often nil or fixed so that asset-pricing is just a matter of model-risk inference, with $\{dS_{t}\}$ reduced to the inferential test equation of Peskir and Shirayev (2006) (for constant $\{r^{D\Delta}_{t}\}$ and $\{\sigma^{D}_{t}\}$).
+Further, it often holds that relative variance $\text{F}^{Z/D}_{t}:=(\sigma^{lZ}_{t})^{2}/(\sigma^{lD}_{t})^{2}\ll 1$ and so $(\sigma^{l}_{t})^{2}\approx(\sigma^{lD}_{t})^{2}$: model-risk is the prime price-mover, not tangible value.
+Dispersion among such assets is thus driven mainly by their differing underlying model-risk inference processes.
+Averaged under model-risk belief $\{{\pi}^{+}_{t}\}$, dominant price-dynamic (32) has this familiar form:
+$\sigma_{t}\bigg{(}\frac{\mu_{t}}{\sigma_{t}}dt+dw_{t}\bigg{)},$ (33)
 with volatility $\sigma_{t}:=\sigma^{\Pi,l}_{t}S^{\Delta}(t)$, drift $\mu_{t}:=\mu^{\Pi,\pi}_{t}S^{\Delta}(t)$, and price-of-diffusion-risk $\mu_{t}/\sigma_{t}$ ((23)):
 $\frac{\mu_{t}}{\sigma_{t}}=(K^{\frac{1}{2}}-K^{-\frac{1}{2}})\sigma^{\Pi}_{t}\sigma^{\pi}_{t}\sigma^{l}_{t}=\frac{K^{\frac{1}{2}}-K^{-\frac{1}{2}}}{S^{\Delta}(t)}\bigg{(}\frac{\sigma^{\pi}_{t}}{\sigma^{\Pi}_{t}}\bigg{)}\sigma_{t}.$ (34)
-Expansion in $(K-1)$, given Remark 9, makes the classical relationship more explicit:
+Expansion in (K1)(K-1), given Remark 7, recovers the standard and ubiquitous price-of-risk vs price-volatility relationship:
 $\frac{\mu_{t}}{\sigma_{t}}\approx\bigg{(}\frac{K-1}{S^{\Delta}(t)}\bigg{)}\sigma_{t}\text{ and }\mu_{t}\approx\bigg{(}\frac{K-1}{S^{\Delta}(t)}\bigg{)}\sigma_{t}^{2}.$ (35)
@@
 risk-aversion).
-Regardless, our derivation links it firmly to ex-ante beliefs, making the above relevant to bias studies and measurements under model-uncertainty.
+Regardless, our derivation links it firmly to ex-ante beliefs, making the above relevant to bias studies and measurement under model-uncertainty.
 3 Bias and Risk-Pricing Measurement via Price Anomalies
 3.1 The Joint-Hypothesis Issue
-Let the $\{\mathbf{F}_{t}\}$-adapted objective conditional probabilities of model-risk outcomes be $\{\mathbf{p}^{B}_{t}\}$.
-The ex-post drift of our asset, $\{rp_{t}\}:=\{\Sigma_{B}\mathbf{p}^{B}_{t}{Y^{B}_{t}}\}-\{S_{t}\}$, assuming $B$-indifferent $B$-sure risk-premia $\{\check{rp}_{t}\}$ (Section 2.2.2) for clarity, satisfies:
+To focus on main issues, as in Section 2.2.2, consider usual-case assets under model-risk $B$ with $B$-independent $B$-sure risk-premia $\{\check{rp}(t)\}$.
+Let $\{\mathbf{p}^{B}_{t}\}$ be the $\{\mathbf{F}_{t}\}$-adapted true conditional probabilities of model-risk outcomes, given the true and known $B$-sure laws $\mathbf{W}^{T}_{B}$ and the true but potentially unknown unconditional law $\mathbf{p}^{B}_{0}$ of model-risk $B$.
+Its ex-post price-drift to be realised on average $\{rp_{t}\}:=\{\Sigma_{B}\mathbf{p}^{B}_{t}{Y^{B}_{t}}\}-\{S_{t}\}$ satisfies:
 $rp_{t}-\check{rp}(t)=B\_RP_{t}+(\mathbf{p}^{+}_{t}-{\pi}^{+}_{t}){S}^{\Delta}(t),\text{ }t<T.$ (36)
-The LHS terms are each known/observable.
-The RHS terms are entangled: no conclusion on risk-pricing $\{B\_RP_{t}\}$ can be made without one on $\{\mathbf{p}^{b}_{t}\}-\{{\pi}^{b}_{t}\}$, namely, bias.
-For our model-risk, any realised process under a realised $B$-outcome corresponds to one run of (31-33) as if $\mathbf{p}^{b}_{t}\equiv 1\text{ or }0$.
-Many ’like’ observations are needed for an estimate of (37).
-Yet, given ’one-off’ and heterogeneous sources of model-uncertainties, how can this be done?
-3.2 Assets Facing Potential Model-Change
-When model-sure, our asset is a simple drift with a static noise.
-Upon some trigger, it may face an objective chance $\mathbf{p}^{(\cdot)}_{0}$ of entering a new state (e.g.
-from $B=+$ to $B=-$).
-Denote any new state by ’1’, any status quo, ’0’, and status quo model-sure pricing, $\{S^{\text{{\bf{0}}}}_{t}\}$.
-Change-trigger may or may not follow a known law.
-Our interest lies in what ensues after its arrival.
-It amounts to treating model-risk as a compound process and ignoring the part governing how/when model-risk arises (considered known, valued and reflected in model-sure prices).
-The average interval between model crises, say $\lambda^{-1}$, is presumed long enough for our one-off methodology; further, consider $t\ll\lambda^{-1}$ and $t\ll T$ henceforth.
-With the economic nature $sign[1{\text{{\bf{0}}}}]:=sign[Y^{1}_{t}-Y^{\bf{0}}_{t}]$ of potential change explicit, ignoring $t$-dependence of change impact $S^{\Delta}(t)$ for $t\ll T$ and $r^{Z\Delta}_{t}\ll S^{\Delta}(t)$, canonical pricing reads ((25)):
-$S_{t}=S^{\text{{\bf{0}}}}_{t}+sign[1{\text{{\bf{0}}}}]S^{\Delta}\cdot\Pi^{1}_{t},\text{ }t\ll T,$ (38)
-where the given RNE belief process $\{\Pi^{1}_{t}\}\sim\{\pi^{1}_{t}\}$ can be characterised by risk-pricing parameter $K:=(\frac{O_{f}[\pi^{1}_{0}]}{O_{f}[{\Pi}^{1}_{0}]})^{sign[1{\text{{\bf{0}}}}]}$ (Corollary 1); let the market be competitive, so $K\in(1,2)$ (Remark 9).
-3.3 The Representation of Status Quo Bias
-The objective conditional probability process $\{\mathbf{p}^{1}_{t}\}$ governing change-risk is determined by the objective but unknown unconditional law $\mathbf{p}^{1}_{0}$ of model-risk.
-Its relationship against the RNE $\{\Pi^{1}_{t}\}$ and reference belief process $\{\pi^{1}_{t}\}$ can be characterised through the following:
-$\rho K^{sign[1{\text{{\bf{0}}}}]}\equiv\frac{O_{f}[\mathbf{p}^{1}_{0}]}{O_{f}[{\Pi}^{1}_{0}]},$ (39)
-$\rho:=\frac{O_{f}[\mathbf{p}^{1}_{0}]}{O_{f}[{\pi}^{1}_{0}]};$ (40)
-for instance, if a change-risk is unpriced, then $K=1$; or if it is classical, with its law $\mathbf{p}^{1}_{0}$ known and so without model-uncertainty, then $\rho=1$.
-Parameter $\rho$, born of model-uncertainty, is not part of any ex-ante economic risk-discounting.
-The joint-hypothesis issue takes a concrete form here: observable prices reflect the product (39), and to split it into ’intention’ $K$ and ’mistake’ $\rho$ external criteria seem essential.
-This, as will be seen, need not be so; indeed, a high apparent bias under model-uncertainty may not be irrational either.
-The case of interest for us is when change is rare, $\mathbf{p}^{1}_{0}<\frac{1}{2}$, but priced as even rarer, such that (39-40) are both sizeable ($\rho\gg 1$).
-This ’justified strong status quo’ serves as a clear base against which ’change’ is defined and assessed (’unjustified strong status quo’, $\mathbf{p}^{1}_{0}>\frac{1}{2}$, allows easy wins to bets on ’change’ and so cannot persist).
-3.4 Unconditional Model-Risk Symmetry, Dominance and Uniformity
-Consider a market of a large number of assets subject to change-risks (e.g.
-new CEO, tort, war, recession).
-Let these risks be as characterised and economically priced in this market (e.g.
-due to their aggregate impact on market value/volatility/dispersion [^5] Dispersion is particularly intriguing in our context.
-As a priced factor it has been linked to market uncertainty, the business cycle and ’structural shifts’ (e.g.
+The LHS terms are each observable or known, but the RHS are joint: no study of risk-pricing $\{B\_RP_{t}\}$ can be done without one on bias $\{\mathbf{p}^{b}_{t}\}-\{{\pi}^{b}_{t}\}$.
+The above, in standard model-risk study settings, with the ’asset’ set to ’equity index’ and ’bias’ to nil, expresses the usual attribution of ’excess historical equity-return’ (LHS) to ’model-risk premium’ (RHS).
+Our focus is bottom-up and when bias is not necessarily nil.
+3.2 Bias and Risk-Pricing Parameters under Model-Uncertainty
+An usual-case asset, if model-sure, is classical (a drift with a static noise).
+Upon the triggering of model-risk, it may enter a new state (e.g.
+from $B=+$ to $B=$-) with true probability $\mathbf{p}^{(\cdot)}_{0}$: denote the new state by ’1’, status quo, ’0’, and status quo model-sure pricing, $\{S^{\text{{\bf{0}}}}_{t}\}$.
+The triggering events may or may not follow known laws.
+The process of how/when they arise is ignored, assuming that the average time between model crises, say some $\lambda^{-1}\gg 1$, is so long that each can be considered one-off.
+The period of interest is right after the arrival of a trigger but long before the next, so $t\ll\lambda^{-1}$, and long before horizon/maturity, so $t\ll T$.
+Canonical asset-pricing (24), with the direction $sign[1{\text{{\bf{0}}}}]:=sign[Y^{1}_{t}-Y^{\bf{0}}_{t}]$ of potential change explicit and the $t$-dependence of impact $S^{\Delta}(t)$ dropped given $t\ll T$ (and so $S^{\Delta}(t)\gg r^{Z\Delta}_{t}$), reads:
+$S_{t}=S^{\text{{\bf{0}}}}_{t}+sign[1{\text{{\bf{0}}}}]S^{\Delta}\cdot\Pi^{1}_{t},$ (37)
+where the RNE belief process $\{\Pi^{1}_{t}\}\sim\{\pi^{1}_{t}\}$ is characterised by parameter $K:=(\frac{O_{f}[\pi^{1}_{0}]}{O_{f}[{\Pi}^{1}_{0}]})^{sign[1{\text{{\bf{0}}}}]}$ (Corollary 1); let the risk-pricing be competitive, so $K\in(1,2)$ (Remark 7).
+The relationship between the true conditional probabilities $\{\mathbf{p}^{1}_{t}\}$ (given the true unconditional law $\mathbf{p}^{1}_{0}$), reference beliefs $\{\pi^{1}_{t}\}$ and RNE beliefs $\{\Pi^{1}_{t}\}$, all $\{\mathbf{F}_{t}\}$-adapted, can be characterised by:
+$\rho:=\frac{O_{f}[\mathbf{p}^{1}_{0}]}{O_{f}[{\pi}^{1}_{0}]},\text{ and}$ (38)
+$\rho K^{sign[1{\text{{\bf{0}}}}]}\equiv\frac{O_{f}[\mathbf{p}^{1}_{0}]}{O_{f}[{\Pi}^{1}_{0}]}.$ (39)
+If the change-risk is diversifiable, so unpriced, then $K=1$; if it is classical, so its law $\mathbf{p}^{1}_{0}$ is known, then $\rho=1$.
+Parameter $\rho$ stems from model-uncertainty and is thus unknown ex-ante.
+The joint-hypothesis problem takes a concrete form: observed prices reflect product (39), so external criteria seem necessary to separate ’intention’ $K$ from ’mistake’ $\rho$.
+We show that this need not be so.
+The case of interest for us is when change is unlikely, $\mathbf{p}^{1}_{0}<\frac{1}{2}$, but is priced as rare so that (39-38) are both sizeable ($\rho\gg 1$).
+Such a strong status quo forms a clear base against which ’change’ is defined and assessed.
+Unjustified strong status quo (i.e.
+$\mathbf{p}^{1}_{0}\geq\frac{1}{2}$ and $\rho\gg 1$) would allow easy wins for ’bet on change’ and so should not persist.
+3.3 An Ideal Model-Uncertainty Driven Market
+Consider a model-uncertainty driven market, of a large number of assets, each facing some change-risk $B$, whose nature may vary from asset to asset (e.g.
+new CEO, tort, war) but whose true $B$-sure laws are known.
+Let the $B$-risks be priced (i.e.
+correlated with the market’s stochastic-discount-factor), for their effects on aggregate value/volatility/cross-sectional dispersion [^6] Dispersion within members of an equity index, as a priced factor, has been linked to market uncertainty, the business cycle and structural shifts (e.g.
 Demirer and Jategaonkar (2013), Kolari et al.
-(2021)).).
-For clarity, let the unconditional all-cause probability of positive/negative economic change be even: at any time half of all potential changes in the market are positive/negative.
-Unconditional unevenness is addressed in Remark 17&18.
-Let ’pricing intention’ $K$ ((39)) and ’apparent bias’ $\rho$ ((40)) be indifferent to the economic nature of potential change.
-Further, let parameter $K$, $\rho$ and change-impact $S^{\Delta}$ be uniform across the assets.
-Such uniformity amounts in practice to asset dispersions being driven by differences in inferential progression rather than in other aspects (Remark 11).
-Similarly, let all risk-inference be based on i.i.d data, with log-LR processes of fixed and uniform volatility.
-Note that each individual realisation of a change-outcome is considered random and independent still; while this may be a stretch for triggers such as war, it is a sensible start.
+(2021))..
+Further, let them be economically priced, so that by Proposition 1 their pricing is canonical with respect to some reference inferential belief process of the relevant $B$-risk.
+Finally, let the realisation and detection of individual change-outcomes be stochastic and independent (even for macroeconomic triggers such as war or recession).
+Note that such a market, when seen as a whole, is an asset with non-binary model-uncertainty; this is so even when there is only one binary macroeconomic model-risk trigger.
+For brevity, let the unconditional all-cause probability of positive/negative economic change be symmetrical: at any time half of all potential changes, of any origin, are positive/negative (unconditional asymmetry is addressed in Remark 14&15).
+Let ’pricing intention’ $K$ ((39)) and ’apparent bias’ $\rho$ ((38)) be indifferent to the economic direction and origin of change-risk, and the size of $K$, $\rho$ and risk-impact $S^{\Delta}$, uniform across the assets.
+Such uniformity in practice amounts to asset dispersions being driven by differences in the paths of model-risk inference rather than in any other aspect (Remark 9).
+Similarly, let all risk-inference be based on log-LR processes that are homogenous in time and uniform across.
 3.4 Conditional Cross-Sectional Trading
 Pair-trading sorts assets into cohorts by cross-sectional statistics for long vs short positions.
-It amounts in our setup to conditioning by $B$-inferential events $\{\Pi^{1}_{t}=v\}$, $v\in(0,1)$, the driver of dispersion.
-By Bayes’ Rule (3) and definition (39), the $\{\Pi^{1}_{t}=v\}$-conditioned ex-post mean-excess $rp(\pm,v):=\pm(\mathbf{p}^{1}_{t}(v)-v)S^{\Delta}$ given the $sign[1{\text{{\bf{0}}}}]=\pm$ of potential change reads:
+It means in our setup conditioning by $B$-inferential events $\{\Pi^{1}_{t}=v\}$, $v\in(0,1)$, the driver of such statistics.
+Under (37-39), given the $sign[1{\text{{\bf{0}}}}]=\pm$ of potential change, the $\{\Pi^{1}_{t}=v\}$-conditioned ex-post mean-excess $rp(\pm,v):=\pm(\mathbf{p}^{1}_{t}(v)-v)S^{\Delta}$ unpacks (by Bayes’ Rule (3)) to:
 $rp(\pm,v)=\pm(\sigma^{v})^{2}\frac{1-(\rho K^{\pm 1})^{-1}}{v+\underline{v}(\rho K^{\pm 1})^{-1}}S^{\Delta},$ (40)
-where $(\sigma^{v})^{2}:=v\underline{v}$, making price-volatility the driver of excess ((29) and (34)).
+where $(\sigma^{v})^{2}:=v\underline{v}$, essentially making price-volatility (see (25-28)) the driver of excess.
 Note that the above is invariant under label-switching $1\leftrightarrow{\text{{\bf{0}}}}$ (i.e.
-$v\leftrightarrow\underline{v}$, +$\leftrightarrow$-+ ↔ - and $\rho\leftrightarrow\rho^{-1}$).
+$v\leftrightarrow\underline{v}$, $\leftrightarrow-$ and $\rho\leftrightarrow\rho^{-1}$).
 Remark 11.
-With no uncertainty, $\rho=1$, the above is the ex-ante risk-premium when $\{\Pi^{1}_{t}=v\}$ ((24)): in classical markets, all the ’strategising’ merely earns what the market deems fair.
-Despite excess (41) being driven by volatility, conditioning by $\{\Pi^{1}_{t}=v\}$, $v\in(0,1)$ under strong status quo amounts to ranking by momentum (past performance): $\{\Pi^{1}_{t}=v\}$ means $\{\Pi^{1}_{t}\}$ going from $\Pi^{1}_{0}\ll\frac{1}{2}$ to $v$ and so an excess-to-date of ±$\pm(v-\Pi^{1}_{0})S^{\Delta}$ ((38)).
-The associated model-uncertainty ($\propto(\sigma^{v})^{2}$) is a priori low and then evolves with dataflow.
-The window of opportunity for pair-trading is given by (C.8), Appendix C.2.
-Drifts are hard to measure (e.g.
-Merton (1980)).
-Volatility trading is a good practical alternative.
-Volatility events, by (26-29), take the form $\{\Pi^{1}_{t}=v\}\cup\{\Pi^{1}_{t}=\underline{v}\}$, $v\in(0,\frac{1}{2}]$, and the $\{\sigma^{\Pi}_{t}=v\underline{v}\}$-conditioned ex-post mean-excess $rp_{t}^{\sigma}(v)$ averages over the $sign[1{\text{{\bf{0}}}}]$ of potential change and over the $\{\Pi^{1}_{t}=v\}$- and $\{\Pi^{1}_{t}=\underline{v}\}$-events in the volatility cohort:
+With no uncertainty, $\rho=1$, the above is by (23) the ex-ante risk-premium given $\{\Pi^{1}_{t}=v\}$: in classical markets all such strategising merely earns what the market deems fair.
+Conditioning, or in practice, ranking, by events $\{\Pi^{1}_{t}=v\}$, $v\in(0,1)$, amounts to, under strong status quo, ranking by momentum (past performance): $\{\Pi^{1}_{t}=v\}$ means $\{\Pi^{1}_{t}\}$ going from $\Pi^{1}_{0}\ll\frac{1}{2}$ to $v$ and so an excess-to-date of $\pm(v-\Pi^{1}_{0})S^{\Delta}$ ((37)).
+The associated model-uncertainty ($\propto(\sigma^{v})^{2}$) is a priori low (’strong status quo’) and evolves with dataflow.
+For efficient pair-trading, the window of opportunity is given by (C.8), Appendix C.2.
+Historical drifts are hard to measure in practice (e.g.
+Merton (1980)), so volatility trading can be an alternative.
+Volatility events $\{\sigma^{\Pi}_{t}=\cdot\}$ have the form $\{\Pi^{1}_{t}=v\}\cup\{\Pi^{1}_{t}=\underline{v}\}$, $v\in(0,\frac{1}{2}]$, by (25-28).
+Such $\{\sigma^{\Pi}_{t}=v\underline{v}\}$-conditioned ex-post mean-excess $rp_{t}^{\sigma}(v)$ is an average over the $sign[1{\text{{\bf{0}}}}]$ of potential change and over the $\{\Pi^{1}_{t}=v\}$- and $\{\Pi^{1}_{t}=\underline{v}\}$-events in the volatility cohort:
 $rp_{t}^{\sigma}(v):=\frac{\sum_{\pm}\mathbb{P}_{t}(\pm,v)rp(\pm,v)+\sum_{\pm}\mathbb{P}_{t}(\pm,\underline{v})rp(\pm,\underline{v})}{\mathbb{P}_{t}(v)+\mathbb{P}_{t}(\underline{v})},$ (41)
@@
 The relevant likelihoods (ratios) are determined by the underlying log-LR dynamics, whose cross-sections are Normal, making closed expressions possible.
-As a by-product, the ratio functions predict an association of volatility with ’negative momentum’ (Remark 17), a well-known phenomenon (e.g.
+As a by-product, these probability ratios predict an association of volatility with ’negative momentum’ (Remark 14), a well-known phenomenon (e.g.
 Ang et al.
 (2006) and Wang and Xu (2015)).
-The risk-reward curves of momentum trading (41) and volatility trading (42) are both concave in $v$ and vanishing as $v\underline{v}\to 0$.
+Observe that both the risk-reward of momentum (40) and volatility trading (41) are driven by $(\sigma^{v})^{2}$ essentially, and are thus concave as a function of $v$ and vanishing as $(\sigma^{v})^{2}\to 0$.
 3.4.1 Momentum Effect and Status Quo Bias
-Conditional excess (41) exhibits Momentum, a key price-anomaly (Jegadeesh and Titman (1993)): if $\rho K^{\pm 1}\gg 1$, excess-to-date ±$\pm(v-\Pi^{1}_{0})S^{\Delta}$, due to RNE beliefs $\{\Pi^{1}_{t}\}$ rising from $\Pi^{1}_{0}\ll\frac{1}{2}$ to $v>\Pi^{1}_{0}$, persists on average in the same direction.
-The effect applies to $\rho K^{\pm 1}\ll 1$ also: past excess due to RNE beliefs $\{\Pi^{1}_{t}\}$ falling from $\Pi^{1}_{0}\approx 1$ to $v<\Pi^{1}_{0}$ means an average future excess of the same sign; indeed, label-switching here defines a new parameter $\rho^{-1}>1$.
-Momentum thus vanishes only under $\rho\approx 1$ (little model-uncertainty).
-The effect grows with $\rho$, which by definition (40) measures bias against change. From (41), momentum trading earns its best realisable excess $rp^{\pm mo}_{max}$ at $v^{\pm mo}_{max}$:
-$v^{\pm mo}_{max}:=\frac{1}{(\rho K^{\pm 1})^{\frac{1}{2}}+1},\text{ }rp^{\pm mo}_{max}:=\pm\frac{(\rho K^{\pm})^{\frac{1}{2}}-1}{(\rho K^{\pm})^{\frac{1}{2}}+1}S^{\Delta}.$ (43)
-Peak-profitability reads $\frac{1}{2}(rp^{+mo}_{max}-rp^{-mo}_{max})$, by pairing cohorts of conditioning-momentum $v^{\pm mo}_{max}$.
-As a function of bias $\rho$, risk-reward curve (41) and its peak (43) imply rising rewards from rising momentum up to a point, beyond which profitability falls, as seen in Ang et al.
+Conditional excess (40) exhibits Momentum, the well-known anomaly (Jegadeesh and Titman (1993)): for $\rho\gg 1$, any excess-to-date $\pm(v-\Pi^{1}_{0})S^{\Delta}$, due to RNE beliefs $\{\Pi^{1}_{t}\}$ rising from $\Pi^{1}_{0}\ll\frac{1}{2}$ to $v>\Pi^{1}_{0}$, persists on average in the same direction.
+The effect grows with $\rho$, which by its definition (38) is an indicator of bias against change.
+Momentum applies also to $\rho\ll 1$: past excess due to RNE beliefs $\{\Pi^{1}_{t}\}$ falling from $\Pi^{1}_{0}\approx 1$ to $v<\Pi^{1}_{0}$ means an average future excess of the same sign; indeed, label-switching defines a new parameter $\rho^{-1}>1$.
+Momentum thus vanishes only when $\rho=1$ (e.g.
+when there is no model-uncertainty).
+Momentum trading (40) earns its best realisable excess $rp^{\pm mo}_{max}$ at $v^{\pm mo}_{max}$:
+$v^{\pm mo}_{max}:=\frac{1}{(\rho K^{\pm 1})^{\frac{1}{2}}+1},\text{ }rp^{\pm mo}_{max}:=\pm\frac{(\rho K^{\pm 1})^{\frac{1}{2}}-1}{(\rho K^{\pm 1})^{\frac{1}{2}}+1}S^{\Delta}.$ (42)
+Peak-profitability reads $\frac{1}{2}(rp^{+mo}_{max}-rp^{-mo}_{max})$, by pairing cohorts of conditioning-momenta $v^{\pm mo}_{max}$.
+As a function of bias $\rho$, risk-reward (40) and its peak (42) imply rising rewards from rising momentum up to a point, beyond which profitability falls, as observed in Ang et al.
 (2006).
 3.4.2 Low-Risk Effect and Status Quo Bias
-Low-Risk Effect (e.g.
+Low-Risk Effect (see e.g.
 Ang et al.
 (2006)) turns the classical dictum ’high risk high reward’ on its head.
-Momentum risk-reward (41) already hints at it: peak-reward (43) is in the low-volatility region if bias $\rho$ is sizeable.
-The effect is clearer under volatility-conditioning (42), whose peak-location and -size can be found by routine methods (Appendix C).
-The solution at $\rho=1$ (no bias) and leading-order solutions at $\rho\gg 1$ (high bias) are revealing and may be given by the same expression; the peak-location $v^{\sigma}_{max}$ and -size $rp^{\sigma}_{max}$ are:
+Momentum risk-reward (40) already hints at it: peak-reward (42) is in the low-volatility zone for large bias $\rho\gg 1$.
+The effect is stronger under volatility-conditioning (41), whose peak-location and -size can be found by usual methods (Appendix C).
+The solution at $\rho=1$ (no bias) and the leading-order solutions at $\rho\gg 1$ (high bias) are revealing and happen to have the same expression; the peak-location $v^{\sigma}_{max}$ and -size $rp^{\sigma}_{max}$ are given by:
 $v^{\sigma}_{max}:=\frac{1}{\rho+1},\text{ }rp^{\sigma}_{max}:=\frac{1}{2}\frac{K-1}{K+1}S^{\Delta}.$ (43)
 At $\rho=1$ the above is exact, by which $v^{\sigma}_{max}=\frac{1}{2}$, where price-volatility peaks, thus confirming ’high risk high reward’ in the absence of model-uncertainty or bias.
-There is otherwise a low-risk effect: peak-reward occurs at $v^{\sigma}_{max}\approx\rho^{-1}$, corresponding to low risk if $\rho\gg 1$; peak volatility, $v=\frac{1}{2}$, is rewarded poorly ex-post: $rp^{\sigma}(\frac{1}{2})=\frac{S^{\Delta}}{2}\mathcal{O}(\frac{K-1}{\rho}-1)$.
-Note the separation of risk-pricing $K$ and bias $\rho$: the latter is revealed by peak-location $v^{\sigma}_{max}$, while the former, by peak-reward $rp^{\sigma}_{max}$, which is that set ex-ante for peak-risk (Remark 9).
-No external rules are needed to tell ’intended pricing’ from ’unintended bias’.
+There is a low-risk effect otherwise: peak-reward occurs at $v^{\sigma}_{max}\approx\rho^{-1}$, while peak volatility ($v=\frac{1}{2}$) is poorly rewarded ex-post: $rp^{\sigma}(\frac{1}{2})=\frac{S^{\Delta}}{2}\mathcal{O}(\frac{K-1}{\rho}-1)$.
+Note the separation: bias $\rho$ is revealed by peak-location $v^{\sigma}_{max}$, and risk-pricing $K$, by peak-reward $rp^{\sigma}_{max}$, which is exactly that set ex-ante for peak-risk (Remark 7).
+No external speculation is needed to tell ’intended pricing’ from ’unintended bias’.
 4 Summary and Discussion
-We investigate the risk-neutral equivalent formulation of model-risk pricing.
-Explicit price dynamics under binary model-risks with potentially unknown laws are derived, exposing model-risk as a main source of excess value, with a key role in Status Quo Bias and anomalies such as Momentum and Low-Risk.
-It offers an integrated and practicable approach to these highly relevant and interconnected topics.
-We make use of a feature of standard hypothesis testing (Lemma 1) that ’informationally equivalent’ tests must be essentially identical.
-This makes risk-neutral equivalent formulations under model-risk path-dependent and intractable.
-On the other hand, it limits viable economic asset-pricing (Remark 1) under model-risk to just one form (Proposition 1), leading to an intuitive risk-pricing formula with intuitive properties (Corollary 1).
-The sole risk-pricing parameter is associated with a conservation law of inference and is a product of ’intention’ (preference) and ’mistake’ (difference between objective and subjective laws).
-Ex-ante (’intended’) model-risk pricing, classical or with ambiguity-aversion/robust-control, can be captured by our parameter $K\in[1,\infty)$ (39), with $K=1$ corresponding to ’no model-risk discounts’.
-Its derivation reveals that it sets the intended gain-loss ratio of the asset at peak model-risk, suggesting $K\in[1,2)$ in competitive markets (Remark 9).
-The difference between true model-risk law and the reference model-risk belief on which asset-pricing is based can be captured by our parameter $\rho\in(0,\infty)$ (40), with $\rho=1$ corresponding to ’no difference’ (the classical setting), and $\rho\not=1$, ’bias’.
-It gives expression to Status Quo Bias and allows its effects to be disentangled from risk-pricing (Section 3.5.1-3.5.2).
+We study the risk-neutral equivalent formulation of model-risk pricing. Explicit price dynamics under binary model-risks with potentially unknown laws are derived.
+The results show that model-risk can be a main source of excess value, with a key role in Status Quo Bias and anomalies such as Momentum and Low-Risk.
+They offer an integrated and practicable approach to these highly relevant and interconnected topics.
+We arrive at the above by making use of the informational constraint imposed by ’usual’ economic asset-pricing (Remark 1 and Definition 1), to show that viable asset-pricing under model-risk may take just one canonical form (Lemma 1 and Proposition 1), which leads to a unique and intuitive model-risk pricing formula with convenient properties (Corollary 1).
+The sole risk-pricing parameter is associated with a conserved quantity of inferential dynamics and is a product of ’intention’ (preference) and ’mistake’ (difference between objective and subjective laws).
+Ex-ante (’intended’) pricing, classical or with ambiguity-aversion/robust-control, can be captured by our parameter $K\in[1,\infty)$, (39), with $K=1$ corresponding to ’no model-risk discounts’.
+Its derivation reveals that it sets the intended gain-to-loss ratio of the asset at peak model-risk, suggesting $K\in[1,2)$ in competitive markets (Remark 7).
+The difference between true model-risk law and the reference model-risk belief on which asset-pricing is based can be captured by our parameter $\rho\in(0,\infty)$, (38), with $\rho=1$ corresponding to ’no difference’ (the classical setting), and $\rho\not=1$, ’bias’.
+It gives expression to Status Quo Bias and allows its effects to be disentangled from risk-pricing (Section 3.4.1-3.4.2).
 The above is interesting also in light of the inertia axiom of Bewley (2002), which says, in effect, that to update an existing decision by incorporating a new state subject to Knightian uncertainty, the status quo is kept unless an alternative that is better in all scenarios exists.
@@
 Asset-prices under max-min utility already have this characteristic, albeit with a built-in pessimism absent in inertial pricing.
-Both are viable in our setup, as models are given by probability laws that are equivalent before risk-resolution.
+Both are viable in our setup, as ’models’ are represented by probability laws that are equivalent before any risk-resolution (that is predictable).
 The theoretical and practical aspects of this topic will be examined in a follow-up study.
-Our approach may be extended to model-risks characterisable as $(N+1)$-tuples of independent binary outcomes ${B}^{N+1}:=(B_{0}B_{1}...B_{N})$, $N\in\mathbb{N}$, through iteration.
-It may be applied first to the $2^{N}$ pairs of ${B}^{N+1}$-sure, elementary, price processes, then to the resulting $2^{N-i}$ pairs of ${B}^{N+1-i}$-sure composites, $i=1,...,N$, each discounting the $i$-tuple model-risk $(B_{0}B_{1}...B_{i-1})$.
 APPENDIX
@@
     Standard Continuous-Time Inference.
-Any continuous process with independent increments are essentially Wiener.
-By definition (Item-4, Section 1.1), the log-LR process under i.i.d data, in terms of some constant $\sigma^{l}>0$ (the ’signal-to-noise’ of data), reads:
+Given the setup and Item-5 of Section 1.1, the log-LR process of time-homogeneous Wiener data processes has the form below (Peskir and Shirayev (2006)), in terms of some constant $\sigma^{l}>0$ characterising the signal-to-noise ratio of the given data,
     $dl^{+/-}_{\tau}(B)=(-1)^{\mathbf{1}_{-}(B)}\frac{(\sigma^{l})^{2}}{2}d\tau+\sigma^{l}dw_{\tau},$ (A.1)
     with $\{w_{\tau}\}$ denoting a standard Wiener process.
-Time-variation may be modelled via an absolutely continuous clock-change $t\mapsto\tau(t)$.
-The image process $\{l^{+/-}_{t}\}:=\{l^{+/-}_{\tau(t)}\}$, again Wiener and with absolutely continuous time-varying parameters (e.g.
-Kallsen (2006)), along with inferential beliefs, an Ito process of it ((3)), read as follows:
-    $dl^{+/-}_{t}(B)=\mu^{l^{+/-}}_{t}(B)dt+\sigma^{l}_{t}dw_{t},\text{ }\text{ }\mu^{l^{+/-}}_{t}(B):=(-1)^{\mathbf{1}_{-}(B)}\frac{(\sigma^{l}_{t})^{2}}{2};$ (A.2)
+Any time-variation may be introduced via an absolutely continuous clock-change $t\mapsto\tau(t)$.
+The image process $\{l^{+/-}_{t}\}:=\{l^{+/-}_{\tau(t)}\}$ is again Wiener, but with continuous time-varying parameters (e.g.
+Kallsen (2006)).
+It and its resulting inferential belief process under Bayes’ Rule (3) read: in terms of some standard Wiener process $\{w_{t}\}$,
+    $dl^{+/-}_{t}(B)=(-1)^{\mathbf{1}_{-}(B)}\frac{(\sigma^{l}_{t})^{2}}{2}+\sigma^{l}_{t}dw_{t};$ (A.2)
     $\frac{d\pi^{+}_{t}(B)}{(\sigma^{\pi}_{t})^{2}}=\big{(}\mathbf{1}_{+}(B)-\pi^{+}_{t}(B)\big{)}(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw_{t}.$ (A.3)
@@
 Its defining feature is the predictable resolution of risky outcomes.
-In usual situations (e.g.
-i.i.d data), risk-resolution occurs at infinity.
-Clock-change allows continuous-time regular inference to have finite resolution-times, by compactifying uniform regular inference to bring resolution $T_{\tau}(B)=\infty$ forward to some $T(B):=\tau^{-1}(\infty)<\infty$.
-The resulting log-LR process diverges as $t\to T(B)$, but the associated beliefs and so asset-pricing remain finite and continuous under Bayes’ Rule (3).
+In standard situations, with i.i.d data, risk-resolution occurs only at infinity.
+Predictable finite-time resolution can be modelled via a clock-change that brings resolution forward to some $T_{\mathcal{B}}\leq\infty$.
+The resulting log-LR process diverges as $t\to T_{\mathcal{B}}$, but the associated beliefs and so asset-pricing remain finite and continuous at $T_{\mathcal{B}}$ under Bayes’ Rule (3).
   3. 3.
-    Complete Markets and Unhedgeable Model-Risks.
-Market completeness excludes unpredictable resolutions (price-jumps) and so demands regular inference.
-In this setting, to be unhedgeable and thus priced economically (Remark 1), model-risk $B$ cannot be resolvable before horizon $T$ (so $T(B)>T$); this is the default case for model-risk studies.
-The mixed case of probabilistic resolution, ($\mathbb{P}\big{(}\{T(B)\leq t+s\}|F_{t}\big{)}\in(0,1)$, $\forall t<T$, $\forall s\in(0,T-t]$, which may occur in continuous-time complete markets so long as $T(B)$ is predictable, involves treating the unhedgeable part of the model-risk only.
+    Non-Resolution and Novikov’s Condition.
+For any log-LR process based on equivalent Wiener measures on the Wiener space $\mathcal{S}$ of total data, written as $\hat{\mathbf{W}}_{B}\sim\mathbf{W}_{B}$ in the financial context, one being some RNE version of the other, the Radon-Nikodym Theorem ensures $\lim_{t\to\infty}\log\frac{d\hat{\mathbf{W}}_{B}|_{t}(\cdot)}{d\mathbf{W}_{B}|_{t}(\cdot)}<\infty$ almost surely and hence a finite total variance for the associated log-LR process of form (A.2).
+The resolution of $B$-outcomes (Item-5, Section 1.1) thus equates to the associated LR process (the exponentiation of (A.2)), or equivalently, (A.3), violating Novikov’s Condition, and vice versa.
   4. 4.
-    Non-Resolution and Novikov’s Condition.
-For any log-LR process based on equivalent Wiener measures on the Wiener space $\mathcal{S}$ of total data, written as $\hat{\mathbf{W}}_{B}\sim\mathbf{W}_{B}$ in the financial context, one being some RNE version of the other, the Radon-Nikodym Theorem ensures $\lim_{t\to\infty}\log\frac{d\hat{\mathbf{W}}_{B}|_{t}(\cdot)}{d\mathbf{W}_{B}|_{t}(\cdot)}<\infty$ almost surely, and hence a finite total variance ((A.2)).
-That is, the non-resolution of $B$-outcomes (Item-4, Section 1.1) equates to (A.2-A.3) satisfying Novikov’s Condition, and vice versa.
-  5. 5.
     Adjacency.
-Given any pair of log-LR process $\{l^{+/-}_{t}\}$ and $\{\hat{l}^{+/-}_{t}\}$ of respective defining laws $\{\mathbf{W}_{+},\mathbf{W}_{-}\}$ and $\{\hat{\mathbf{W}}_{+},\hat{\mathbf{W}}_{-}\}$ such that equivalence $\hat{\mathbf{W}}_{B}\sim\mathbf{W}_{B}$ holds on the space $\mathcal{S}$ of total data, then they almost surely can differ at most by a constant as $t\to T(B)$, since:
+Given any pair of log-LR process $\{l^{+/-}_{t}\}$ and $\{\hat{l}^{+/-}_{t}\}$ of respective defining laws $\{\mathbf{W}_{+},\mathbf{W}_{-}\}$ and $\{\hat{\mathbf{W}}_{+},\hat{\mathbf{W}}_{-}\}$ such that equivalence $\hat{\mathbf{W}}_{B}\sim\mathbf{W}_{B}$ holds on the space $\mathcal{S}$ of total data, then they almost surely can differ at most by a constant as $t\to\infty$, since:
     $\hat{l}^{+/-}_{t}-l^{+/-}_{t}=\log{\frac{d\hat{\mathbf{W}}_{+}|_{t}}{d\mathbf{W}_{+}|_{t}}}-\log{\frac{d\hat{\mathbf{W}}_{-}|_{t}}{d\mathbf{W}_{-|_{t}}}},\text{ }\forall t\in\mathbb{R}^{+},$ (A.4)
-    where the RHS is convergent under equivalence.
+    where the RHS is convergent under equivalence; this is the case even when each of $\{l^{+/-}_{t}\}$ and $\{\hat{l}^{+/-}_{t}\}$ is resolving (i.e. divergent in total variance and so total drift).
     Lemma 1.
-    No adjacent log-LR processes, unless identical up to a constant difference, can be informationally equivalent, that is, be deterministically and time-homogenously related to one another, nor can any inferential processes derived from them.
-    Proof.
-It is a feature of inference (and due to Bayes’ Rule).
-It can be verified concretely via Ito’s Lemma here, given (A.2) and restrictions to twice-differentiable maps between processes (as in Proposition 1).
-Any such mapping $g:{l}^{+/-}_{t}\in\mathbb{R}\mapsto\hat{l}^{+/-}_{t}=g(l^{+/-}_{t})\in\mathbb{R}$, $t\in(0,T(B))$, must solve the ordinary differential equation:
+    No two resolving log-LR processes, unless identical up to a constant difference, can be deterministically and time-homogenously related to one another.
+[…]
     $g^{\prime\prime}=-(-1)^{\mathbf{1}_{+}(B)}(g^{\prime}-1)g^{\prime}.$ (A.5)
-    If $B=+$, then $g^{\prime}e^{-g}=g^{\prime}(0)e^{-g(0)}e^{-\mathbf{Id}}$ and so $g=g(0)-\log[g^{\prime}(0)e^{-\mathbf{Id}}+1-g^{\prime}(0)]$; divergence $\lim_{t\to T(+)}\hat{l}^{+/-}_{t}=\infty$ requires $g^{\prime}(0)=1$; If $B=-$, then $g=g(0)+\log[g^{\prime}(0)e^{\mathbf{Id}}+1-g^{\prime}(0)]$ and $\lim_{t\to T(-)}\hat{l}^{+/-}_{t}=-\infty$ imply the same.
-    Remark 14.
-    Given identical LR up to a constant factor, inference process $\{\hat{\pi}^{+}_{t}\}$, a $(\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B})$-martingale by definition, is also a $(\hat{\pi}^{B}_{0}\mathbf{W}_{B})$-martingale ((3)), meaning $\hat{\mathbf{W}}_{B}=\mathbf{W}_{B}$ (Martingale Representation Theorem and so completeness of the Wiener basis): thus, only tests based on identical underlying beliefs can be informationally equivalent.
+    If $B=+$, then $g^{\prime}e^{-g}=g^{\prime}(0)e^{-g(0)}e^{-\mathbf{Id}}$, so $g=g(0)-\log[g^{\prime}(0)e^{-\mathbf{Id}}+1-g^{\prime}(0)]$; we then have $g^{\prime}(0)=1$ as $\{\hat{l}^{+/-}_{t}\}$ and $\{{l}^{+/-}_{t}\}$ are resolving.
+If $B=$-, so $g=g(0)+\log[g^{\prime}(0)e^{\mathbf{Id}}+1-g^{\prime}(0)]$, then almost-sure divergence $\lim_{t\to\infty}\hat{l}^{+/-}_{t}=-\infty=\lim_{t\to\infty}{l}^{+/-}_{t}$ requires the same.
 B PROOF OF PROPOSITION 1
-Proof.
-Part I.
-Viable model-risk only pricing $\{S^{(0|A)}_{t}\}:=\{\Sigma_{B}A^{B}_{t}Y^{B}_{t}\}=\{Y^{-}_{t}\}+\{A^{+}_{t}Y^{\Delta}_{t}\}$, given twice-differentiable coefficients $\{A^{+}_{t}(\pi^{+}_{t})\}$, must have the canonical form.
-Under reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$, with $d(\mathbf{W}^{T}_{-})_{t}=\sigma^{Z}_{t}dw^{Z}_{t}+\sigma^{D}_{t}dw^{D}_{t}$, $\{dw^{Z}_{t}\}$ and $\{dw^{Z}_{t}\}$ being independent standard Wiener processes (Item-2&4, Section 1.1), we have:
-$dS^{(0|A)}_{t}=\big{[}($ $\mathbf{1}_{+}-A^{+}_{t})r^{Z\Delta}_{t}dt\text{ }+\text{ }\sigma^{Z}_{t}dw^{Z}_{t}\big{]}\text{ }+\text{ }$ (B.1)
-$(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)\partial_{\pi}A^{+}_{t}\cdot\bigg{[}($ $\mathbf{1}_{+}-\pi^{+}_{t})(\sigma^{l}_{t})^{2}dt+(\sigma^{\pi}_{t})^{2}\frac{\partial_{\pi}^{2}A^{+}_{t}}{2\partial_{\pi}A^{+}_{t}}(\sigma^{l}_{t})^{2}dt\text{ }+\text{ }\sigma^{l}_{t}dw^{\prime}_{t}\bigg{]},$ (B.2)
-with model-drift $r^{Z\Delta}_{t}dt=-{d}Y^{\Delta}(t)$ and $(\sigma^{l}_{t})^{2}:=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$, where $\sigma^{lZ}_{t}:={r^{Z\Delta}_{t}}/{\sigma^{Z}_{t}}$ (’signal-to-noise’ ratio) stems from data $\{Z_{t}(B)\}$ (Peskir and Shirayev (2006)), and $\sigma^{lD}_{t}$ likewise, from $B$-conditionally independent data $\{D_{t}(B)\}$, and so $\sigma^{l}_{t}dw^{\prime}_{t}:=\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}$.
+[…]
+Under reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$, with $d\mathbf{W}^{T}_{-,t}=\sigma^{Z}_{t}dw^{Z}_{t}+\sigma^{D}_{t}dw^{D}_{t}$, $\{dw^{Z}_{t}\}$ and $\{dw^{Z}_{t}\}$ being independent standard Wiener processes (Item-3&5, Section 1.1), we have:
+[…]
+$(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)(A^{+}_{t})^{\prime}_{\pi}\bigg{[}($ $\mathbf{1}_{+}-\pi^{+}_{t})\cdot(\sigma^{l}_{t})^{2}dt+(\sigma^{\pi}_{t})^{2}\frac{(A^{+}_{t})^{\prime\prime}_{\pi}}{2(A^{+}_{t})^{\prime}_{\pi}}\cdot(\sigma^{l}_{t})^{2}dt\text{ }+\text{ }\sigma^{l}_{t}dw_{t}\bigg{]},$ (B.2)
+with model-drift $r^{Z\Delta}_{t}dt=-{d}Y^{\Delta}(t)$ (Item-2, Section 1.1) and $(\sigma^{l}_{t})^{2}:=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$, where $\sigma^{lZ}_{t}:={r^{Z\Delta}_{t}}/{\sigma^{Z}_{t}}$ (’signal-to-noise’) stems from data $\{Z_{t}(B)\}$ (Peskir and Shirayev (2006)), with $\sigma^{lD}_{t}$ likewise from $B$-conditionally independent data $\{D_{t}(B)\}$, and so $\sigma^{l}_{t}dw_{t}:=\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}$.
 Without risk-pricing, $\{A^{+}_{t}\}=\{\pi^{+}_{t}\}$, we have trivial viability under reference belief $\pi^{B}_{0}\mathbf{W}^{T}_{B}$:
-$dS^{(0|\pi)}_{t}=\big{[}(\mathbf{1}_{+}-\pi^{+}_{t})r^{Z\Delta}_{t}dt+\sigma^{Z}_{t}dw^{Z}_{t}\big{]}+(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)\big{[}(\mathbf{1}_{+}-\pi^{+}_{t})(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw^{\prime}_{t}\big{]}.$ (B.3)
-With viable model-risk only pricing, under some RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}\sim\pi^{B}_{0}\mathbf{W}_{B}^{T}$: we have $\{S^{(0|A)}_{t}\}=\{\Sigma_{B}\hat{\pi}^{B}_{t}\hat{S}^{B}_{t}\}$ and the following relationship, by (8-11) and (19),
+$dS^{(0|\pi)}_{t}=\big{[}(\mathbf{1}_{+}-\pi^{+}_{t})\cdot r^{Z\Delta}_{t}dt+\sigma^{Z}_{t}dw^{Z}_{t}\big{]}+(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)\big{[}(\mathbf{1}_{+}-\pi^{+}_{t})\cdot(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw_{t}\big{]}.$ (B.3)
+With viable model-risk only pricing, under some RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}\sim\pi^{B}_{0}\mathbf{W}_{B}^{T}$, we have $\{S^{(0|A)}_{t}\}=\{\Sigma_{B}\hat{\pi}^{B}_{t}\hat{S}^{B}_{t}\}$, and the following, by (8-11) and (19),
 $\pi^{+}_{t}-A^{+}_{t}=k^{{A}}_{t}\sigma^{\pi}_{t}=k^{\hat{\pi}}_{t}\sigma^{\pi}_{t}+\frac{\Sigma_{B}\hat{\pi}^{B}_{t}\hat{RP}^{B}_{t}}{{Y}^{\Delta}(t)}=(\pi^{+}_{t}-\hat{\pi}^{+}_{t})+\frac{\Sigma_{B}\hat{\pi}^{B}_{t}\hat{RP}^{B}_{t}}{{Y}^{\Delta}(t)}.$ (B.4)
 Remark 12.
-Reference model-risk beliefs $\{{\pi}^{B}_{t}\}$ and any RNE $\{\hat{\pi}^{B}_{t}\}$ are ’adjacent’ (Item-5, Appendix A), implying that at any fixed $t<T$ for $\sigma^{\pi}_{t}\ll\frac{1}{2}$, by (23-24) when $\{{\hat{\pi}}^{B}_{t}\}=\{\Pi^{B}_{t}\}$, we have $\mathcal{O}({\pi}^{B}_{t}-{\hat{\pi}}^{B}_{t})=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and so $\mathcal{O}({k}^{\hat{\pi}}_{t})=\mathcal{O}(\sigma^{\pi}_{t})$ for any RNE price-of-model-risk ((11)).
-In turn, by Lemma 1, to balance the LHS and RHS of (B.4), we have: $\mathcal{O}({\hat{RP}^{B}_{t}})=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and $\mathcal{O}({k}^{A}_{t})=\mathcal{O}(\sigma^{\pi}_{t})$, as well as $\mathcal{O}(\partial_{\pi}A^{B}_{t})=\mathcal{O}(\partial^{2}_{\pi}A^{B}_{t})=\mathcal{O}(1)$ and $\mathcal{O}(d({\hat{RP}^{B}_{t})})=\mathcal{O}(d(\sigma^{\pi}_{t})^{2})=\mathcal{O}((\sigma^{\pi}_{t})^{4})$.
-Consider first the case of trivial $\{D_{t}\}$.
-The drift of $\{dS^{(0|A)}_{t}\}$ is nil in expectation under some RNE measure $\hat{\pi}^{B}_{0}\hat{W}_{B}^{Z}|_{T}$, where $d(\hat{W}^{Z}_{B})_{t}=\hat{r}_{B,\text{ }t}dt+\sigma^{Z}_{t}dw^{Z}_{t}$, $\hat{r}_{B,\text{ }t}dt=-d(\hat{RP}^{B}_{t})$, $t<T$.
-Being the RNE measure, the $(\hat{\pi}^{B}_{0}\hat{W}_{B}^{Z}|_{T})$-expectation of (B.1) must offset that of (B.2).
-Write that of (B.1) as $\nu_{t}:=(\hat{\pi}^{+}_{t}-A^{+}_{t})r^{Z\Delta}_{t}-\Sigma_{B}\hat{\pi}^{B}_{t}\hat{r}_{B,\text{ }t}$ and note $\hat{\pi}^{+}_{t}-\pi^{+}_{t}=(\hat{\pi}^{+}_{t}-A^{+}_{t})-(\pi^{+}_{t}-A^{+}_{t})$, viability thus requires:
-$\frac{-\nu_{t}}{r^{Z\Delta}_{t}}=(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)\partial_{\pi}A^{+}_{t}\cdot\frac{(\sigma^{lZ}_{t})^{2}}{r^{Z\Delta}_{t}}\bigg{[}(\sigma^{\pi}_{t})^{2}\frac{\partial_{\pi}^{2}A^{+}_{t}}{2\partial_{\pi}A^{+}_{t}}-({\pi}^{+}_{t}-A^{+}_{t})-\frac{-\nu_{t}}{r^{Z\Delta}_{t}}\bigg{]}.$ (B.5)
-By Remark 15, we have $\mathcal{O}(RHS)=\mathcal{O}((\sigma^{\pi}_{t})^{4})$ and $\mathcal{O}(LHS)=\mathcal{O}((\sigma^{\pi}_{t})^{2})$, $\forall\sigma^{\pi}_{t}\ll\frac{1}{2}$, so nil-drift is possible only if each side of the above vanishes on its own, meaning:
-$\frac{\partial_{\pi}^{2}A^{+}_{t}}{2\partial_{\pi}A^{+}_{t}}=\frac{{\pi}^{+}_{t}-A^{+}_{t}}{(\sigma^{\pi}_{t})^{2}}.$ (B.6)
+Reference model-risk beliefs $\{{\pi}^{+}_{t}\}$ and any of its RNE $\{\hat{\pi}^{+}_{t}\}$ are adjacent (Item-4, Appendix A), implying that at any fixed $t<T$ for $\sigma^{\pi}_{t}\ll\frac{1}{2}$, by (23) when $\{{\hat{\pi}}^{+}_{t}\}=\{\Pi^{+}_{t}\}$, we have $\mathcal{O}({\pi}^{+}_{t}-{\hat{\pi}}^{+}_{t})=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and so $\mathcal{O}({k}^{\hat{\pi}}_{t})=\mathcal{O}(\sigma^{\pi}_{t})$ for any RNE price-of-model-risk (11).
+In turn, by Lemma 1, the LHS and RHS of (B.4) being equal implies $\mathcal{O}({\hat{RP}^{+}_{t}})=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and $\mathcal{O}({k}^{A}_{t})=\mathcal{O}(\sigma^{\pi}_{t})$, as well as $\mathcal{O}((A^{+}_{t})^{\prime}_{\pi})=\mathcal{O}((A^{+}_{t})^{\prime\prime}_{\pi})=\mathcal{O}(1)$ and $\mathcal{O}(d{\hat{RP}^{+}_{t}})=\mathcal{O}(d(\sigma^{\pi}_{t})^{2})=\mathcal{O}((\sigma^{\pi}_{t})^{4})$.
+Consider first the case of trivial $\{D_{t}\}$, so that the drift of $\{dS^{(0|A)}_{t}\}$ is nil in expectation under some RNE measure $\hat{\pi}^{B}_{0}\hat{W}_{B}^{Z}|_{T}$ satisfying $d\hat{W}^{Z}_{B,t}=\hat{r}_{B,\text{ }t}dt+\sigma^{Z}_{t}dw^{Z}_{t}$, $\hat{r}_{B,\text{ }t}dt=-d\hat{RP}^{B}_{t}$, $t<T$.
+[…]
+$\frac{-\nu_{t}}{r^{Z\Delta}_{t}}=(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)(A^{+}_{t})^{\prime}_{\pi}\frac{(\sigma^{lZ}_{t})^{2}}{r^{Z\Delta}_{t}}\bigg{[}(A^{+}_{t}-{\pi}^{+}_{t})+(\sigma^{\pi}_{t})^{2}\frac{(A^{+}_{t})^{\prime\prime}_{\pi}}{2(A^{+}_{t})^{\prime}_{\pi}}-\frac{-\nu_{t}}{r^{Z\Delta}_{t}}\bigg{]}.$ (B.5)
+By Remark 12, we have $\mathcal{O}(LHS)=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and $\mathcal{O}(RHS)=\mathcal{O}((\sigma^{\pi}_{t})^{4})$, $\forall\sigma^{\pi}_{t}\ll\frac{1}{2}$, so nil-drift is possible only if each side vanishes on its own, meaning:
+$\frac{(A^{+}_{t})^{\prime\prime}_{\pi}}{2(A^{+}_{t})^{\prime}_{\pi}}=\frac{{\pi}^{+}_{t}-A^{+}_{t}}{(\sigma^{\pi}_{t})^{2}}.$ (B.6)
 This has solutions of the form $\{A^{+}_{t}\}=\{\Pi^{+}_{t}\}$ only, where $\{\Pi^{+}_{t}\}$ is any model-risk inference based on any measure $\Pi^{B}_{0}W_{B}^{Z}|_{T}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.
-The same can be obtained from the LHS of (B.5).
-Adding $B$-conditionally independent data $\{D_{t}\}$ under law $W_{B}^{D}|_{T}$, the argument above applies unchanged to the dynamics of $\{dS^{(0|A)}_{t}\}$, (B.1-B.2), under reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$ and some RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}$ such that (B.4) holds and $d(\hat{\mathbf{W}}_{B}^{T})_{t}=\hat{r}_{B,\text{ }t}dt+d(\mathbf{W}_{-}^{T})_{t}$, $\hat{{r}}_{B,\text{ }t}dt=-d(\hat{{RP}}^{B}_{t})$.
+Adding $B$-conditionally independent data $\{D_{t}\}$ under law $W_{B}^{D}|_{T}$, the argument above applies unchanged to the dynamics of $\{dS^{(0|A)}_{t}\}$, (B.1-B.2), under the reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$ and some RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}$ such that (B.4) holds and $d\hat{\mathbf{W}}^{T}_{B,t}=\hat{r}_{B,\text{ }t}dt+d\mathbf{W}^{T}_{-,t}$, $\hat{{r}}_{B,\text{ }t}dt=-d\hat{{RP}}^{B}_{t}$.
 To conclude, the RNE measure $\{S^{(0|A)}_{t}\}$ must have the canonical form: $\Pi^{B}_{0}\mathbf{W}_{B}^{T}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.
-Part II. Viability of canonical normal pricing $\{S^{(RP|{\Pi})}_{t}\}=\{{S}^{(0|{\Pi})}_{t}\}-\{\Sigma_{B}{RP}^{B}(t)\Pi^{B}_{t}\}$.
-Under the martingale measure $\Pi^{B}_{0}\mathbf{W}_{B}^{T}$ of process $\{{S}^{(0|{\Pi})}_{t}\}$, process $\{S^{(RP|\Pi)}_{t}\}$ is a martingale of form (B.3) plus the extra drift $R^{B}_{t}dt=-{d}(RP^{B}(t))$ that stems from $B$-sure risk-pricing:
-$dS^{(RP|\Pi)}_{t}=\big{[}(\mathbf{1}_{+}-\Pi^{+}_{t})\check{r}^{Z\Delta}_{t}dt+\sigma^{Z}_{t}dw^{Z}_{t}\big{]}+S^{\Delta}(t)(\sigma^{\Pi}_{t})^{2}\big{[}(\mathbf{1}_{+}-\Pi^{+}_{t})(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw^{\prime}_{t}\big{]}+R^{(\cdot)}_{t}dt,$ (B.7)
-where $\check{r}^{Z\Delta}_{t}dt:=-{d}{}S^{\Delta}(t)$, with $S^{\Delta}(t)\equiv Y^{\Delta}(t)-RP^{+}(t)+RP^{-}(t)$.
-As such, it is viable so long as this extra drift $\{R^{B}_{t}dt\}$ can be ’absorbed’ into the known $(\Pi^{B}_{0}\mathbf{W}_{B}^{T})$-martingale $\{d{S}^{(0|{\Pi})}_{t}\}$, that is, whenever
+Part II. The viability of canonical price process $\{S^{(RP|{\Pi})}_{t}\}=\{{S}^{(0|{\Pi})}_{t}\}-\{\Sigma_{B}{RP}^{B}(t)\Pi^{B}_{t}\}$.
+Under the martingale measure $\Pi^{B}_{0}\mathbf{W}_{B}^{T}$ of process $\{{S}^{(0|{\Pi})}_{t}\}$, process $\{S^{(RP|\Pi)}_{t}\}$ is a martingale of form (B.3) plus the extra drift $R^{B}_{t}dt=-{d}RP^{B}(t)$ that stems from $B$-sure risk-pricing:
+$dS^{(RP|\Pi)}_{t}=\big{[}(\mathbf{1}_{+}-\Pi^{+}_{t})\cdot\check{r}^{Z\Delta}_{t}dt+\sigma^{Z}_{t}dw^{Z}_{t}\big{]}+S^{\Delta}(t)(\sigma^{\Pi}_{t})^{2}\big{[}(\mathbf{1}_{+}-\Pi^{+}_{t})\cdot(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw_{t}\big{]}+R^{(\cdot)}_{t}dt,$ (B.7)
+where $\check{r}^{Z\Delta}_{t}dt:=-{d}{}S^{\Delta}(t)$, with $S^{\Delta}(t)\equiv Y^{\Delta}(t)-RP^{+}(t)+RP^{-}(t)$ (by (12-14).
+As such, it is viable if drift $\{R^{B}_{t}dt\}$ can be ’absorbed’ into the known $(\Pi^{B}_{0}\mathbf{W}_{B}^{T})$-martingale $\{d{S}^{(0|{\Pi})}_{t}\}$, that is, whenever
 $\frac{{R^{\pm}_{t}}/{S^{\Delta}(t)}}{{\sigma^{Z}_{t}}/{S^{\Delta}(t)}+{(\sigma^{\Pi}_{t})^{2}{\sigma^{lD}_{t}}\sqrt{1+\text{F}^{Z/D}_{t}}}}$
-is square-integrable up to horizon $T$ (Novikov’s Condition), where, for convenience, we have written $(\sigma^{l}_{t})^{2}=(\sigma^{lD}_{t})^{2}(1+\text{F}^{Z/D}_{t})$, with $\text{F}^{Z/D}_{t}:=(\sigma^{lZ}_{t})^{2}/(\sigma^{lD}_{t})^{2}$.
+is square-integrable up to horizon $T$ (Novikov’s Condition), where $(\sigma^{l}_{t})^{2}\equiv(\sigma^{lD}_{t})^{2}(1+\text{F}^{Z/D}_{t})$ is written in terms of the relative variance $\text{F}^{Z/D}_{t}:=(\sigma^{lZ}_{t})^{2}/(\sigma^{lD}_{t})^{2}$ of data $\{Z_{t}\}$ to $\{D_{t}\}$.
 C THE MOMENTUM AND LOW-RISK FORMULAE
-For the assets of Section 3.2-3.4, all the distributions required stem from the underlying log-LR process $\{l^{1/{\text{{\bf{0}}}}}_{t}\}$, of parameter $\sigma^{l}$, for which log-LR levels $l^{1/{\text{{\bf{0}}}}}_{t}$ follow the Normal distribution:
+For the assets of Section 3.2-3.3, all the distributions required stem from the underlying log-LR process $\{l^{1/{\text{{\bf{0}}}}}_{t}\}$, of parameter $\sigma^{l}$, for which log-LR levels $l^{1/{\text{{\bf{0}}}}}_{t}$ follow the Normal distribution:
 $\mathcal{N}\bigg{(}\mu^{l}_{[t]}(B)=\frac{(-1)^{B+1}}{2}(\sigma^{l})^{2}t,\text{ }(\sigma^{l}_{[t]})^{2}=(\sigma^{l})^{2}t\bigg{)},\text{ }B\in\{0,1\}.$ (C.1)
-The distributions of RNE beliefs $\{\Pi^{1}_{t}\}$ and so of prices ((38)) then derive via change-of-variable:
+The distributions of RNE beliefs $\{\Pi^{1}_{t}\}$ and so of prices (37) then derive via (3) by change of variable: for $\Pi^{1}_{t}=v$ given $l^{1/{\text{{\bf{0}}}}}_{t}=l$, we have $dl=(v\underline{v})^{-1}dv$, with:
 $l=H_{\Pi\pm}+\log\circ\text{ }O_{f}[v],$ (C.2)
@@
 $H_{\mathbf{p}}:=\log\circ\text{ }O_{f}[\underline{\mathbf{p}}^{1}_{0}],$ (C.4)
-for $\Pi^{1}_{t}=v$ given $l^{1/{\text{{\bf{0}}}}}_{t}=l$, with $dl=(v\underline{v})^{-1}dv$ ((3)) and $O_{f}[\Pi^{1}_{0\pm}]=O_{f}[\mathbf{p}^{1}_{0}]/(\rho K^{\pm 1})$ ((39)) given $sign[1{\text{{\bf{0}}}}]=\pm$.
+where recall (39), $O_{f}[\Pi^{1}_{0\pm}]=O_{f}[\mathbf{p}^{1}_{0}]/(\rho K^{\pm 1})$ given $sign[1{\text{{\bf{0}}}}]=\pm$.
 The $H$-variables above are inferential milestones: $H_{\Pi\pm}$ is the log-LR hurdle for event $\{\Pi^{1}_{t}\geq\frac{1}{2}\}$, and $H_{\mathbf{p}}$, for $\{\mathbf{p}^{1}_{t}\geq\frac{1}{2}\}$.
 C.1 Tracking Inferential Progress and Bias Dominance
-The degree of $B$-certainty as data accumulate may be assessed in the usual way given Normal distributions.
-The objective degree is tracked by some $|C^{\mathbf{p}}_{t}|$, where, with $t_{\mathbf{p}}:=2H_{\mathbf{p}}/(\sigma^{l})^{2}$,
+The degree of $B$-certainty as data accumulate may be assessed in the standard way for Normal distributions.
+Its objective level is tracked by some $|C^{\mathbf{p}}_{t}|$, where, with $t_{\mathbf{p}}:=2H_{\mathbf{p}}/(\sigma^{l})^{2}$,
 $C^{\mathbf{p}}_{t}(B):=\frac{H_{\mathbf{p}}-\mu^{l}_{[t]}(B)}{\sigma^{l}_{[t]}}=\frac{1}{2}\sigma^{l}_{[t]}\cdot\big{(}{(-1)^{B}}+\frac{t_{\mathbf{p}}}{t}\big{)}.$ (C.5)
-Certainty is high for $\sigma^{l}_{[t]}\ll 1$ (little data) or $\sigma^{l}_{[t]}\gg 1$ (lots of data), with $|C^{\mathbf{p}}_{t}|$ bottoming at 00 if and when $H_{\mathbf{p}}=\mu^{l}_{[t]}(B)$, that is, if and when $t=(-1)^{B+1}t_{\mathbf{p}}$ has a solution.
+Certainty level $|C^{\mathbf{p}}_{t}|$ is high if $\sigma^{l}_{[t]}\ll 1$ (little data) or $\sigma^{l}_{[t]}\gg 1$ (lots of data); it bottoms out at 0 if and when $H_{\mathbf{p}}=\mu^{l}_{[t]}(B)$ has a solution, that is, if and when $t=(-1)^{B+1}t_{\mathbf{p}}$ and $B=1$.
 For the price-implied RNE inference process $\{\Pi^{1}_{t}\}$, the same indicator, denoted $|C^{\Pi}_{t\pm}|$, reads:
@@
 $=$ $\frac{1}{2}\sigma^{l}_{[t]}\cdot(\frac{t_{\rho}}{t}\pm\frac{t_{K}}{t})+C^{\mathbf{p}}_{t}(B),$ (C.7)
-[…]
+where $t_{\Pi\pm}=t_{\rho}\pm t_{K}+t_{\mathbf{p}}$, with $t_{\rho}:={2\log(\rho)}/{(\sigma^{l})^{2}}>0$ and $t_{K}:={2\log K}/{(\sigma^{l})^{2}}>0$ (recall $K-1\not\gg 1$ in general and $K\in(1,2)$ in competitive markets).
+The pattern of $|C^{\Pi}_{t\pm}|$ parallels that of $|C^{\mathbf{p}}_{t}|$; it has 0 as a minimum if and when $t=(-1)^{B+1}t_{\Pi\pm}>0$.
 Bias creates a ’burden of proof’ $t_{\rho}$, and $t_{\rho}/t$ tracks how it is overcome.
-We consider cases of positive objective hurdle $H_{\mathbf{p}}>0$ and bias domination $H_{\Pi\pm}-H_{\mathbf{p}}\gg 0$, so $t_{\mathbf{p}}>0$ and $t_{\rho}\gg t_{K}$, where ’objective burden’ $t_{\mathbf{p}}$ is presumed ’non-extreme’ for non-trivial inference/excess.
+We consider cases of positive objective hurdle $H_{\mathbf{p}}>0$ and bias domination $H_{\Pi\pm}-H_{\mathbf{p}}\gg 0$, so $t_{\mathbf{p}}>0$ and $t_{\rho}\gg t_{K}$, where ’objective burden’ $t_{\mathbf{p}}$ is ’non-extreme’, to exclude easy inference or profits.
 C.2 The Window of Opportunity
@@
 Remark 13.
-Consider for example profitable events $\{O_{f}[{\Pi}^{1}_{t}]=\rho^{-1}\}$ or $\{O_{f}[{\Pi}^{1}_{t\pm u}]=\rho^{-1}\}$ ((44)), with $t$ fixed and inside the above window, and $u>0$ large enough for the second event to be outside.
+Consider profitable events $\{O_{f}[{\Pi}^{1}_{t}]=\rho^{-1}\}$ and $\{O_{f}[{\Pi}^{1}_{t\pm u}]=\rho^{-1}\}$ (see (43)), with $t$ fixed inside the window above and $u>0$ large enough for the second event to be outside the window.
 Their respective probability densities are $\propto(t)^{-1/2}e^{-|C^{\mathbf{p}}_{t}|^{2}/2}$ and $\propto(t\pm u)^{-1/2}e^{-|C^{\mathbf{p}}_{(t\pm u)}|^{2}/2}$.
-Thus, in-window events dominate, with a log-LR of $\propto\big{(}\log[1\pm u/t]+|C^{\mathbf{p}}_{(t\pm u)}|^{2}-|C^{\mathbf{p}}_{t}|^{2}\big{)}\gg 1$, which diverges as $(1-u/t)\to 0$ or $(1+u/t)\to\infty$, with $t$ fixed.
+Thus, in-window events dominate, with a log-LR of $\log[1\pm u/t]+|C^{\mathbf{p}}_{(t\pm u)}|^{2}-|C^{\mathbf{p}}_{t}|^{2}\gg 1$, which diverges as $(1-u/t)\to 0$ or $(1+u/t)\to\infty$, $t$ fixed.
 C.3 Momentum and Volatility Mixtures
 The probabilities relevant to volatility-conditioned mean-excess (41) obey the rule:
-$\mathbb{P}_{t}(\Pi^{1}_{t}=v)\mathbb{P}_{t}(\pm|\text{ }\Pi^{1}_{t}=v)\equiv\mathbb{P}_{t}(\pm)\mathbb{P}_{t}(\Pi^{1}_{t}=v|\text{ }\pm),$ (C.9)
+[…]
 with $\mathbb{P}_{t}(\Pi^{1}_{t}=v|\text{ }\pm)$ from Normal distribution (C.1) via (C.2).
-The likelihood-ratio of $\{\Pi^{1}_{t}=\underline{v}\}$ vs $\{\Pi^{1}_{t}=v\}$, $v\in(0,\frac{1}{2}]$, given the $sign[1{\text{{\bf{0}}}}]=\pm$ of potential change, is:
+The likelihood-ratio of events $\{\Pi^{1}_{t}=\underline{v}\}$ vs $\{\Pi^{1}_{t}=v\}$, $v\in(0,\frac{1}{2}]$, given the $sign[1{\text{{\bf{0}}}}]=\pm$ of potential change, is:
 $R_{t|\pm}(\underline{v}/v)(B):=\frac{\mathbb{P}_{t}(\underline{v}|\pm)}{\mathbb{P}_{t}(v|\pm)}(B)=({\underline{v}}/{v})^{-(t_{\Pi\pm}/{t})-(-1)^{B}},\text{ }B\in\{0,1\}.$ (C.10)
-Note that $\{\Pi^{1}_{t}=v\}$-events dominate as long as bias does (i.e. $t_{\rho}/{t}\gg 1$) regardless; its dependence on actual $B$-outcomes is weak.
+Note that $\{\Pi^{1}_{t}=v\}$-events dominate when bias does (i.e. $t_{\rho}/{t}\gg 1$), regardless of $B$-outcomes.
 To compute (41) we also need the mix-ratio function $M_{t|v}:={\mathbb{P}_{t}(+|v)}/{\mathbb{P}_{t}(-|v)}$ of the event-set $\{\Pi^{1}_{t}=v\}$, $v\in(0,1)$, between member-events with $\{sign[1{\text{{\bf{0}}}}]=+\}$ and with $\{sign[1{\text{{\bf{0}}}}]=-\}$.
-By (C.9) we have $M_{t|v}={\mathbb{P}_{t}(v|+)}/{\mathbb{P}_{t}(v|-)}$, since $\mathbb{P}_{t}(\pm)=\frac{1}{2}$ by setup. Hence:
-$M_{t|v}(B)$ $=(\frac{\rho v}{\underline{v}})^{-t_{K}/t}\cdot K^{-(t_{\mathbf{p}}/t)-(-1)^{B}},$ (C.11)
+Given (C.9), we have $M_{t|v}={\mathbb{P}_{t}(v|+)}/{\mathbb{P}_{t}(v|-)}$, since $\mathbb{P}_{t}(\pm)=\frac{1}{2}$ by setup.
+Hence:
+$M_{t|v}(B)$ $=(\frac{\rho v}{\underline{v}})^{-t_{K}/t}K^{-(t_{\mathbf{p}}/t)-(-1)^{B}},$ (C.11)
 and likewise the volatility-conditioned ratio $M_{t|v\underline{v}}:={\mathbb{P}_{t}(+|v\underline{v})}/{\mathbb{P}_{t}(-|v\underline{v})},\text{ }v\in(0,\frac{1}{2}]$:
 $M_{t|v\underline{v}}=M_{t|v}\frac{1+R_{t|+}(\underline{v}/v)}{1+R_{t|-}(\underline{v}/v)}.$ (C.12)
-Note the source of conditionality: risk-pricing $K$; if $K=1$, conditional mixes equal the unconditional mix; if $K\in(1,2)$, they are perturbations of the unconditional one.
+Note the source of the conditionalities above, model-risk pricing $K$.
+For $K=1$, unpriced model-risk, conditional mixes equal the unconditional.
+For $K\in(1,2)$, competitively priced model-risk, conditional mixes are perturbations of the unconditional.
 Remark 14.
-Both mix-ratios (C.11-C.12) are monotone declining, and peak-volatility ($v=\frac{1}{2}$) brings $M_{t|\frac{1}{2}}=M_{t|\frac{1}{2}\frac{1}{2}}\approx\rho^{-t_{K}/t}=\exp[{-(t_{\rho}/t)\log K}]$.
+Both mix-ratios (C.11-C.12) are monotone declining in $v$, and peak-volatility ($v=\frac{1}{2}$) brings $M_{t|\frac{1}{2}}=M_{t|\frac{1}{2}\frac{1}{2}}\approx\rho^{-t_{K}/t}=\exp[{-(t_{\rho}/t)\log K}]$.
 That is, for $\rho\gg 1$, in the window (C.8) of opportunity, when trading events are most abundant, the observed mix at high volatility can be highly negative vs the unconditional level.
-Mix returns to its unconditional state as $t\to\infty$ when model-risk resolves.
+Mix reverts to its unconditional level as $t\to\infty$ when model-risk resolves.
 Any uneven unconditional background merely introduces a constant factor in (C.11-C.12); all above statements remain valid.
@@
 Further, given (C.10) and under condition (C.8), the $\{\Pi^{1}_{t}=\underline{v}\}$-contributions to (41) are no more than $\mathcal{O}(\rho^{-t_{\rho}/{t}})$.
-That is, the leading order solution (44) is not affected by these contributions while $t_{\rho}/t\gg 1$, i.e.
+That is, the leading order solution (43) is not affected by these contributions whenever $t_{\rho}/t\gg 1$, i.e.
 while bias dominates.
 Remark 15.
-For uneven unconditional-mix, $\mathbb{P}_{t}(\pm)\not=\frac{1}{2}$, the same applies.
-Compared to even solution (44), its peak moves left (right), and size, down (up), for unevenness that tilts negatively (positively); peak-location is capped by $v^{+mo}_{max}$ ((43)), and peak-size, by $rp^{+mo}_{max}$ ((43)).
-Without uncertainty or bias, $\rho=1$, both the $\{\Pi^{1}_{t}=v\}$- and $\{\Pi^{1}_{t}=\underline{v}\}$-part of (42) are invariant under $v\leftrightarrow\underline{v}$ switching ((41) and (C.10-C.12)), so their derivatives both vanish at $v=\frac{1}{2}$, thus confirming ’high risk high reward’ under classical conditions.
+For uneven unconditional-mix, $\mathbb{P}_{t}(\pm)\not=\frac{1}{2}$, the same features apply.
+For unevenness that tilts negatively (positively), the location of peak-excess moves left (right) vs the even solution (43), and its size, down (up).
+Without uncertainty or bias, $\rho=1$, both the $\{\Pi^{1}_{t}=v\}$- and $\{\Pi^{1}_{t}=\underline{v}\}$-contributions to (41) are invariant under $v\leftrightarrow\underline{v}$ switching (see (40) and (C.10-C.12)), so their derivatives both vanish at $v=\frac{1}{2}$, thus confirming peak-volatility as the location of peak-reward and ’high risk high reward’ under classical conditions.

### v7

@@
 Remark 1.
-Some model-risks of this type may be hedged or diversified away.
+Some model-risks of this type may be replicated/hedged or diversified.
 Our results apply only to the economically priced: any deemed worthy of being assigned at each moment a risk-premium through some recursive optimisation scheme taking said risk as an input.
@@
   1. 1.
-    Any incremental change (if at all) to bullet-payoff $Y$ is such that the cumulative firmed-up (’cum-div’) asset-worth $\{Z_{t}\}$, with $Z_{T}\equiv Y$, follows the dynamics below, based on a standard Wiener process $\{w^{Z}_{t}\}$, with natural filtration $\{\mathcal{F}^{Z}_{t}\}$, subject to binary outcomes $\{B=b\}$ (e.g.
-’high’ or ’low’) labelled by $b\in\mathcal{B}:=\{+,-\}$,
+    Let $Y$ be revealed continuously by a public record $\{Z_{t}\}$ of realised increments (decrements) to some known base, set to $\log 1=0$, where each $Z_{t}$ is the sum of such firmed-up changes to date; by setup $Y\equiv Z_{T}$.
+Let $\{Z_{t}\}$ be a diffusion, based on some standard Wiener process $\{w^{Z}_{t}\}$, with natural filtration $\{\mathcal{F}^{Z}_{t}\}$, subject to outcome $\{B=b\}$, $b\in\mathcal{B}:=\{+,-\}$,
     $dZ_{t}(B)=r^{Z}_{B,t}dt+\sigma^{Z}_{t}dw^{Z}_{t},\text{ }t\in(0,T],$ (1)
     where its $B$-sure drifts $\{r^{Z}_{B,t}\}$ and $B$-independent volatilities $\{\sigma^{Z}_{t}\}$ are both $\{\mathcal{F}^{Z}_{t}\}$-predictable.
-The underlying $B$-sure probability space $(\mathcal{S}^{Z}|_{T},W^{Z}_{B}|_{T})$ is the restriction to $(0,T]$ of the Wiener space/measure $(\mathcal{S}^{Z},W^{Z}_{B})$ associated with process $\{Z_{t}(B)\}$, that is, $dW^{Z}_{B,t}=dZ_{t}(B)$.
+The underlying $B$-sure space $(\mathcal{S}^{Z}|_{T},W^{Z}_{B}|_{T})$ is the restriction to $(0,T]$ of the Wiener space $(\mathcal{S}^{Z},W^{Z}_{B})$ associated with $\{Z_{t}(B)\}$, that is, $dW^{Z}_{B,t}=dZ_{t}(B)$.
 The above allows the sequential testing of Wiener processes of differing drifts [^2] Text-book sequential tests (e.g.
@@
     Remark 3.
-    Existing studies ignore potential structures within dataflow ($\{\mathbf{D}_{t}\}$), as if only data relevant to tangible, firmed-up value, $\{Z_{t}\}$, exist/matter.
-By considering data that are exclusive to model-risk (price-irrelevant once model-sure), including when tangible-value data, $\{Z_{t}\}$, are absent or trivial, we tackle a frequent challenge facing partitioners: a crisis market unsettled by a bout of model-risk whose economics dominate, while non-model, routine, risks remain, independently evolving or unchanging.
+    Existing studies ignore potential structures within dataflow ($\{\mathbf{D}_{t}\}$), as if only data relevant to tangible firmed-up value, $\{Z_{t}\}$, exist/matter.
+By considering data exclusively relevant to model-risk (irrelevant once model-sure), we tackle a frequent challenge facing partitioners: a crisis market unsettled by a bout of model-risk whose economics dominate asset valuation and hence traders’ cognitive and trading activities, while non-model, routine, risks remain, independently evolving or hardly changing.
   4. 4.
@@
     Standard Continuous-Time Inference.
-Given the setup and Item-5 of Section 1.1, the log-LR process of time-homogeneous Wiener data processes has the form below (Peskir and Shirayev (2006)), in terms of some constant $\sigma^{l}>0$ characterising the signal-to-noise ratio of the given data,
+Given setup (Section 1.1), the log-LR process for testing time-homogeneous Wiener data of drift $\mathbf{1}_{+}(B)r^{Z}$ and volatility $\sigma^{Z}$ has the form (Peskir and Shirayev (2006)): in terms of signal-to-noise $\sigma^{l}=\frac{r^{Z}}{\sigma^{Z}}$,
     $dl^{+/-}_{\tau}(B)=(-1)^{\mathbf{1}_{-}(B)}\frac{(\sigma^{l})^{2}}{2}d\tau+\sigma^{l}dw_{\tau},$ (A.1)
-    with $\{w_{\tau}\}$ denoting a standard Wiener process.
-Any time-variation may be introduced via an absolutely continuous clock-change $t\mapsto\tau(t)$.
-The image process $\{l^{+/-}_{t}\}:=\{l^{+/-}_{\tau(t)}\}$ is again Wiener, but with continuous time-varying parameters (e.g.
+    with $\{w_{\tau}\}$ being some standard Wiener process.
+Time-dependence obtains via an absolutely continuous clock-change $t\mapsto\tau(t)$: image $\{l^{+/-}_{t}\}:=\{l^{+/-}_{\tau(t)}\}$ is Wiener, with continuous time-varying parameters (e.g.
 Kallsen (2006)).
-It and its resulting inferential belief process under Bayes’ Rule (3) read: in terms of some standard Wiener process $\{w_{t}\}$,
+It and the resulting belief process by Bayes’ Rule (3) read: in terms of a standard Wiener process $\{w_{t}\}$ in $t$-time,
     $dl^{+/-}_{t}(B)=(-1)^{\mathbf{1}_{-}(B)}\frac{(\sigma^{l}_{t})^{2}}{2}+\sigma^{l}_{t}dw_{t};$ (A.2)
@@
   2. 2.
+    Non-Resolution and Novikov’s Condition.
+For inference based on equivalent Wiener-measure pair ${\mathbf{W}}_{+}\sim\mathbf{W}_{-}$, the Radon-Nikodym Theorem ensures $\lim_{t\to\infty}\log\frac{d{\mathbf{W}}_{+}|_{t}(\cdot)}{d\mathbf{W}_{-}|_{t}(\cdot)}<\infty$ and so a finite total variance for the log-LR process (A.2) almost surely.
+Risk-resolution (Item-5, Section 1.1) thus equates to the LR process, the exponentiation of (A.2), or equivalently (A.3), violating Novikov’s Condition, and vice versa.
+  3. 3.
     Regular Inference.
-Its defining feature is the predictable resolution of risky outcomes.
-In standard situations, with i.i.d data, risk-resolution occurs only at infinity.
-Predictable finite-time resolution can be modelled via a clock-change that brings resolution forward to some $T_{\mathcal{B}}\leq\infty$.
-The resulting log-LR process diverges as $t\to T_{\mathcal{B}}$, but the associated beliefs and so asset-pricing remain finite and continuous at $T_{\mathcal{B}}$ under Bayes’ Rule (3).
-  3. 3.
-    Non-Resolution and Novikov’s Condition.
-For any log-LR process based on equivalent Wiener measures on the Wiener space $\mathcal{S}$ of total data, written as $\hat{\mathbf{W}}_{B}\sim\mathbf{W}_{B}$ in the financial context, one being some RNE version of the other, the Radon-Nikodym Theorem ensures $\lim_{t\to\infty}\log\frac{d\hat{\mathbf{W}}_{B}|_{t}(\cdot)}{d\mathbf{W}_{B}|_{t}(\cdot)}<\infty$ almost surely and hence a finite total variance for the associated log-LR process of form (A.2).
-The resolution of $B$-outcomes (Item-5, Section 1.1) thus equates to the associated LR process (the exponentiation of (A.2)), or equivalently, (A.3), violating Novikov’s Condition, and vice versa.
+It resolves $B$-outcomes predictably and so continuously.
+With i.i.d data risk-resolution occurs only at $T_{\mathcal{B}}=\infty$.
+Finite-time resolution can be modelled via a clock-change that brings resolution forward.
+Note that log-LR processes (A.2) diverge as $t\to T_{\mathcal{B}}\leq\infty$ while the associated beliefs (A.3) stay finite and continuous.
+For assets with horizon or maturity $T<T_{\mathcal{B}}$ or with $T_{\mathcal{B}}\leq T<\infty$ but with model-risk $B$ not $\mathcal{F}^{Z}_{T}$-measurable, our model-risk pricing applies all the same on $0,T_{\mathcal{B}}\wedge T)$ regardless.
+In the rest of the cases, where model-risk $B$ is $\mathcal{F}^{Z}_{T}$-measurable and hence replicable/hedgeable, its ’economic pricing’ (Remark 1) becomes irrelevant anyway.
   4. 4.
     Adjacency.
-Given any pair of log-LR process $\{l^{+/-}_{t}\}$ and $\{\hat{l}^{+/-}_{t}\}$ of respective defining laws $\{\mathbf{W}_{+},\mathbf{W}_{-}\}$ and $\{\hat{\mathbf{W}}_{+},\hat{\mathbf{W}}_{-}\}$ such that equivalence $\hat{\mathbf{W}}_{B}\sim\mathbf{W}_{B}$ holds on the space $\mathcal{S}$ of total data, then they almost surely can differ at most by a constant as $t\to\infty$, since:
+Given any pair of log-LR process $\{l^{+/-}_{t}\}$ and $\{\hat{l}^{+/-}_{t}\}$ of respective defining laws $\{\mathbf{W}_{+},\mathbf{W}_{-}\}$ and $\{\hat{\mathbf{W}}_{+},\hat{\mathbf{W}}_{-}\}$ such that equivalence $\hat{\mathbf{W}}_{B}\sim\mathbf{W}_{B}$ holds on the space $\mathcal{S}$ of total data, then they almost surely can differ at most by a finite amount as $t\to\infty$, since:
     $\hat{l}^{+/-}_{t}-l^{+/-}_{t}=\log{\frac{d\hat{\mathbf{W}}_{+}|_{t}}{d\mathbf{W}_{+}|_{t}}}-\log{\frac{d\hat{\mathbf{W}}_{-}|_{t}}{d\mathbf{W}_{-|_{t}}}},\text{ }\forall t\in\mathbb{R}^{+},$ (A.4)
-    where the RHS is convergent under equivalence; this is the case even when each of $\{l^{+/-}_{t}\}$ and $\{\hat{l}^{+/-}_{t}\}$ is resolving (i.e. divergent in total variance and so total drift).
+    where the RHS is bounded (Item-2), even if $\{l^{+/-}_{t}\}$ and $\{\hat{l}^{+/-}_{t}\}$ are resolving (divergent).
     Lemma 1.
-    No two resolving log-LR processes, unless identical up to a constant difference, can be deterministically and time-homogenously related to one another.
+    No two resolving log-LR processes can be related to each other deterministically and time-homogenously without being identical up to a constant difference.
 […]
     $g^{\prime\prime}=-(-1)^{\mathbf{1}_{+}(B)}(g^{\prime}-1)g^{\prime}.$ (A.5)
-    If $B=+$, then $g^{\prime}e^{-g}=g^{\prime}(0)e^{-g(0)}e^{-\mathbf{Id}}$, so $g=g(0)-\log[g^{\prime}(0)e^{-\mathbf{Id}}+1-g^{\prime}(0)]$; we then have $g^{\prime}(0)=1$ as $\{\hat{l}^{+/-}_{t}\}$ and $\{{l}^{+/-}_{t}\}$ are resolving.
-If $B=$-, so $g=g(0)+\log[g^{\prime}(0)e^{\mathbf{Id}}+1-g^{\prime}(0)]$, then almost-sure divergence $\lim_{t\to\infty}\hat{l}^{+/-}_{t}=-\infty=\lim_{t\to\infty}{l}^{+/-}_{t}$ requires the same.
+    If $B=+$: $g=g(0)-\log[g^{\prime}(0)e^{-\mathbf{Id}}+1-g^{\prime}(0)]$; so $g^{\prime}(0)=1$ as $\lim_{t\to T_{\mathcal{B}}}\hat{l}^{+/-}_{t}=\infty=\lim_{t\to T_{\mathcal{B}}}{l}^{+/-}_{t}$.
+If $B=$-: $g=g(0)+\log[g^{\prime}(0)e^{\mathbf{Id}}+1-g^{\prime}(0)]$; so $g^{\prime}(0)=1$ as $\lim_{t\to T_{\mathcal{B}}}\hat{l}^{+/-}_{t}=-\infty=\lim_{t\to T_{\mathcal{B}}}{l}^{+/-}_{t}$.
 B PROOF OF PROPOSITION 1
@@
 Adding $B$-conditionally independent data $\{D_{t}\}$ under law $W_{B}^{D}|_{T}$, the argument above applies unchanged to the dynamics of $\{dS^{(0|A)}_{t}\}$, (B.1-B.2), under the reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$ and some RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}$ such that (B.4) holds and $d\hat{\mathbf{W}}^{T}_{B,t}=\hat{r}_{B,\text{ }t}dt+d\mathbf{W}^{T}_{-,t}$, $\hat{{r}}_{B,\text{ }t}dt=-d\hat{{RP}}^{B}_{t}$.
-To conclude, the RNE measure $\{S^{(0|A)}_{t}\}$ must have the canonical form: $\Pi^{B}_{0}\mathbf{W}_{B}^{T}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.
 Part II. The viability of canonical price process $\{S^{(RP|{\Pi})}_{t}\}=\{{S}^{(0|{\Pi})}_{t}\}-\{\Sigma_{B}{RP}^{B}(t)\Pi^{B}_{t}\}$.

### v8

@@
 We thus work with three types of probabilities: 1) the objectively true ones governing events; 2) the subjective ones representing inferential beliefs about the objective world; and 3) the RNE ones for pricing.
-Let our model-risk $B$ have two outcomes, $b$ or $\overline{b}\not=b$, and our $B$-sure models, elementary (’model-risk free’).
-Such $B$-sure processes represent ’routines’ that by classical RE principles are well-modelled; only change creates model-risk (e.g.
+Let our model-risk $B$ have two outcomes, $\{B=b\}$ or $\{B=\overline{b}\not=b\}$, and our $B$-sure models, free of model-risk.
+The $B$-sure processes represent ’routines’, which by classical RE principles must be well-modelled.
+Only change may create model-risk (e.g.
 new CEO, R&D results, macroeconomic, regulatory or geopolitical events).
@@
 Remark 1.
-Some model-risks of this type may be replicated/hedged or diversified.
-Our results apply only to the economically priced: any deemed worthy of being assigned at each moment a risk-premium through some recursive optimisation scheme taking said risk as an input.
-Our model-risk shall be considered one-off.
-How it ’flares up’ and the pricing of ’flare-risk’ remains implicit; that is, we study episodic ’model crises’, one at a time.
-This contrasts with the workhorse Markov regime-switching setting of the literature, where the pricing of the entire process is treated.
-Further, despite possible uncertainty and indefinite priors, our model-risk is resolvable via regular inference, albeit not before any finite horizon; this is also the case in Hansen-Sargent’s robust-control setting (e.g.
-Hansen et al.
-(2006)).
-Correspondingly, let horizons be sufficiently long for the meaningful development of model-risk inference, and let asset-pricing be time-homogeneous (’identical risk-levels identically priced’) during such developments (i.e.
-while the given model-risk remains active and relevant).
+Some risks of this type may be diversified away, or replicable and so hedged out.
+Our study concerns only the economically priced: any deemed worthy of being given at each moment a risk-premium via some recursive optimisation scheme with said risk as an input.
+Model-risk $B$ is to be treated as one-off.
+How it ’flares up’ and the pricing of ’flare risk’ are implicit; that is, we study episodic ’model crises’, one at a time.
+Such a backdrop puts the core issues in sharp relief and is a good approximation of both real-world settings and the favourite of the literature, Markov hidden regime-switching that is rare.
+Let horizon or maturity be long enough for meaningful model-risk inference to develop, and let asset-pricing be time-homogeneous (’identical risk-levels identically priced’) during such periods.
 The law of our model-risk, the unconditional distribution of $B$-outcomes, may be unknown.
@@
 If unknown, the same may involve non-classical cognitive and economic elements.
-Nonetheless, asset-pricing may be formulated in terms of classical inference, with non-classical aspects reflected through preference and discount mechanisms; this is the usual, and our, strategy.
-Lastly, data on model-risk (e.g.
-weekly/monthly indicators of economic regimes) tend to be more frequent and impactful than data on ’routine’ risks (e.g.
-biannual dividend).
-Such structures, absent in existing studies but relevant to practice, are treated explicitly here.
-The setup as outlined above makes model-risk a prominent feature of asset-pricing, a major source of excess value, and so a focus of market effort and activities.
-From the Perspectives of Standard Model-Risk Studies
+Nonetheless, asset-pricing may be formulated in terms of classical inference, with non-classical aspects reflected through preference and discount mechanisms; this is the usual strategy and ours.
+Our main point of departure from existing treatments is the separation of model-risk from the rest, the non-model risks, in terms of its inference and discounting, to arrive at a tractable characterisation of RNE asset-pricing under model-risk despite path-dependency.
+The separation is reflected also in our information structure, through the potential presence of data exclusive to model-risk, as opposed to the standard one, which is generated by the irreducible inputs of asset-worth alone (e.g.
+dividends).
+It mimics the common scenario where data about model-risk (e.g.
+economic indicators) are far more varied and frequent than the realisations of irreducibles (e.g.
+earnings).
+It also enables and enriches the economic analysis of uncertainties with sure pre-horizon resolution.
+Such uncertainties, although common in practice, fall outside the remit of current theories, which require model or parameter risks to stay unresolved, or as in Hansen-Sargent’s robust-control, resolved at infinity (e.g.
+Hansen et al.
+(2006)), since risks surely resolvable pre-horizon in a standard information structure are replicable/hedgeable and so priced by arbitrage.
+Our setup accommodates unreplicable model-risks that are however sure to resolve continuously before horizon or maturity.
+The Perspectives of Standard Model-Risk Studies
 Most model-risk pricing schemes, ambiguity-aversion, robust-control and some parameter-learning, achieve model-risk discounts by in effect pessimistically distorting a reference belief (about model-risk), known as ’reference measure’, ’second-order prior’, or ’benchmark’.
@@
 We show how this issue may resolve itself in our approach.
-Finally, existing theories require ’total risk’ as a whole as inputs, model and non-model risks together, aggregated across all relevant assets in a given market.
-They are unwieldy for bottom-up (individual-asset) or cross-sectional enquiries, especially when the model-risk(s) can be micro-economic or heterogeneous in origin.
-Our method, with its isolation of model-risk pricing and RNE formulation, is more flexible and convenient for such applications.
 Content and Organisation
-Section 1 sets out the usual asset and data structure in continuous time, highlighting the effect of model-risk on standard risk-pricing considerations; Section 2 derives RNE asset-pricing under model-risk; Section 3 demonstrates a potential application, followed by conclusions and discussions; technical backgrounds and proofs are collected in the Appendix.
-We find that the RNE formulations of asset-pricing under model-risk are constrained by economic, informational, necessities (Definition 1 and Lemma 1) and so have a canonical form (Proposition 1), resulting in familiar properties (Corollary 1) and dynamics (Section 2.2.2).
-One implication is a potential link between Status Quo Bias and well-known anomalies, whose characteristic risk-reward curves reveal and separate ex-ante risk-premia from bias under model-uncertainty (Section 3.4.1&3.4.2).
+Section 1 sets out the asset and data structure in continuous time, highlighting the effect of model-risk on standard risk-pricing considerations; Section 2 derives RNE asset-pricing under model-risk; Section 3 demonstrates a potential application, followed by conclusions and discussions; technical backgrounds and proofs are collected in the Appendix.
+The RNE formulations of asset-pricing under model-risk are shown to be constrained by economic, informational, necessities (Definition 1 and Lemma 1) to have a canonical form (Proposition 1), resulting in familiar properties (Corollary 1) and dynamics (Section 2.2.2).
+One implication is a potential link, given model-uncertainty, between Status Quo Bias and well-known anomalies, whose characteristic risk-reward curves can identify and separate ex-ante risk-premia from bias (Section 3.4.1&3.4.2).
 1 Setup
 1.1 The Usual-Case Asset and Information Structure
-For clarity, consider an usual-case asset: it has just one, bullet, payoff, $Y\in\mathbb{R}$, at some horizon or maturity $T$ much longer than unit-time, so $1\ll T<\infty$; let value and price be in logarithm, and time- and risk-free discount exogenous and nil (in a suitable numéraire).
-Further:
+Consider an usual-case asset, one with a sole payoff $Y\in\mathbb{R}$, in logarithmic value, at a far off horizon $T$, $1\ll T<\infty$, with time- and risk-free discount set to nil (in a suitable numéraire):
   1. 1.
-    Let $Y$ be revealed continuously by a public record $\{Z_{t}\}$ of realised increments (decrements) to some known base, set to $\log 1=0$, where each $Z_{t}$ is the sum of such firmed-up changes to date; by setup $Y\equiv Z_{T}$.
-Let $\{Z_{t}\}$ be a diffusion, based on some standard Wiener process $\{w^{Z}_{t}\}$, with natural filtration $\{\mathcal{F}^{Z}_{t}\}$, subject to outcome $\{B=b\}$, $b\in\mathcal{B}:=\{+,-\}$,
-    $dZ_{t}(B)=r^{Z}_{B,t}dt+\sigma^{Z}_{t}dw^{Z}_{t},\text{ }t\in(0,T],$ (1)
-    where its $B$-sure drifts $\{r^{Z}_{B,t}\}$ and $B$-independent volatilities $\{\sigma^{Z}_{t}\}$ are both $\{\mathcal{F}^{Z}_{t}\}$-predictable.
-The underlying $B$-sure space $(\mathcal{S}^{Z}|_{T},W^{Z}_{B}|_{T})$ is the restriction to $(0,T]$ of the Wiener space $(\mathcal{S}^{Z},W^{Z}_{B})$ associated with $\{Z_{t}(B)\}$, that is, $dW^{Z}_{B,t}=dZ_{t}(B)$.
-The above allows the sequential testing of Wiener processes of differing drifts [^2] Text-book sequential tests (e.g.
-Peskir and Shirayev (2006)) concern processes of fixed drift and volatility.
-Our setup allows time-variations achieved through an absolutely continuous clock-change, or equivalently, $\{\mathcal{F}^{Z}_{t}\}$-predictable drifts and volatilities.
-For more background see Appendix A..
+    Let $Y$ be revealed continuously by a public record $\{Z_{t}\}$ of realised incremental changes to a known base level, set to $\log 1=0$, where each $Z_{t}$ is the sum of firmed-up changes to date, with $Z_{T}\equiv Y$ by definition.
+Let $\{Z_{t}\}$ be a diffusion, based on a standard Wiener process $\{w^{Z}_{t}\}$, generating standard Wiener space $\mathcal{S}^{Z}$, measure $w^{Z}$ and natural filtration $\{\mathcal{F}^{Z}_{t}\}$, subject to outcome $\{B=b\}$, $b\in\mathcal{B}:=\{+,-\}$, where its $B$-sure drift $\{r^{Z}_{B,t}\}$ and $B$-independent volatility $\{\sigma^{Z}_{t}\}$ are $\{\mathcal{F}^{Z}_{t}\}$-predictable:
+    $dZ_{t}(B)=r^{Z}_{B,t}dt+\sigma^{Z}_{t}dw^{Z}_{t},\text{ }t\in[0,\infty).$ (1)
+    The Wiener space $(\mathcal{S}^{Z},W^{Z}_{B})$ of the $B$-sure process $\{Z_{t}(B)\}$, with $dW^{Z}_{B,t}=dZ_{t}(B)$, provides the state space and measure $(\mathcal{S}^{Z}|_{T},W^{Z}_{B}|_{T})$ of $B$-sure payoffs via restrictions to $T$.
+Model-risk inference amounts to the standard sequential testing of Wiener processes (Peskir and Shirayev (2006)), assuming ${W}^{Z}_{+}\perp{W}^{Z}_{-}$ and ${W}^{Z}_{+}|_{t}\sim{W}^{Z}_{-}|_{t}$, $\forall t<\infty$, so that data $\{Z_{t}\}$ allow $B$-resolution at $T^{Z}_{\mathcal{B}}:=\inf\big{\{}t|\mathcal{F}^{Z}_{t}\ni\{B=b\}\big{\}}=\infty$ almost surely [^2] Text-book cases concern constant drifts and volatility, from which tests with $\{\mathcal{F}^{Z}_{t}\}$-predictable parameters derive via an absolutely continuous clock-change (see e.g.
+Kallsen (2006) and also Appendix A)..
   2. 2.
-    The economic impact of model-risk $B$ thus derives from drift difference $r^{Z\Delta}_{t}:=r^{Z}_{+,t}-r^{Z}_{-,t}$; let it be consistent: $Y^{\Delta}(t):=\int_{t}^{T}r^{Z\Delta}_{s}ds>0$ almost surely, at any $t<T$.
-This is intuitive, also necessary for viable and economically valid risk-pricing, as will be seen (Remark 8).
+    Then, given $Z_{t}=z_{t}$, $t<T$, the $B$-sure payoff expectation process $\{Y^{B}_{t}\}:=\{{\mathbf{E}}_{B}[Y|\mathcal{F}^{Z}_{t}]\}$, a well-known martingale, takes the following form:
+    $Y^{B}_{t}=z_{t}+\mathbf{E}_{B}[Y-z_{t}|{\mathcal{F}^{Z}_{t}}]=z_{t}+\mathbf{E}_{B}[Y-z_{t}]=:z_{t}+y_{B}(T-t,t);$ (2)
+    note $y_{B}(0,T)\equiv 0$ and $y_{B}(T-t,t)=\mathbf{E}_{B}[\int_{t}^{T}dZ_{t}]=\mathbf{E}[\int_{t}^{T}{r}^{Z}_{B,s}ds]$, where $\mathbf{E}[\cdot|\cdot]$ denotes expectations under the standard Wiener measure $w^{Z}$.
   3. 3.
-    There may be a purely $B$-informative real-vector dataflow $\{{D}_{t}(B)\}$ that is $B$-conditionally independent but otherwise similar to $\{Z_{t}(B)\}$, based on a standard Wiener process $\{w^{D}_{t}\}$,
-    $dD_{t}(B)=r^{D}_{B,t}dt+\sigma^{D}_{t}dw^{D}_{t}.$ (2)
-    It by definition is price-irrelevant in the absence of model-risk (when $B$-sure).
-Combined, total dataflow $\{\mathbf{D}_{t}\}:=\{(Z_{t},D_{t})\}$ has state space $\mathcal{S}^{T}:=(\mathcal{S}^{Z}\times\mathcal{S}^{D})|_{T}$ and natural filtration $\{\mathbf{F}_{t}\}:=\{\mathcal{F}^{Z}_{t}\}\vee\{\mathcal{F}^{D}_{t}\}$, under the product $B$-sure law $\mathbf{W}^{T}_{B}:=\mathbf{W}_{B}|_{T}$, with $\mathbf{W}_{B}:={W}^{Z}_{B}\times{W}^{D}_{B}$.
+    Define model-risk impact $Y^{\Delta}(t):=Y^{+}_{t}-Y^{-}_{t}=\mathbf{E}[\int_{t}^{T}r^{Z\Delta}_{s}ds]$, with $r^{Z\Delta}_{t}:=r^{Z}_{+,t}-r^{Z}_{-,t}$.
+Let it be consistent: $Y^{\Delta}(t)>0$, $\forall t<T$.
+This is intuitive, and necessary, as it turns out, for viable and economically valid risk-pricing (Remark 9).
+  4. 4.
+    There is a purely $B$-informative real-vector dataflow $\{{D}_{t}\}$ that is $B$-conditionally independent to $\{Z_{t}\}$ (so price-irrelevant if given $B$) based on standard Wiener process $\{w^{D}_{t}\}$,
+    $dD_{t}(B)=r^{D}_{B,t}dt+\sigma^{D}_{t}dw^{D}_{t}=:dW^{D}_{B,t},\text{ }t\in[0,T^{D}_{\mathcal{B}}),$ (3)
+    where $T^{D}_{\mathcal{B}}:=\inf\big{\{}t|\mathcal{F}^{D}_{t}\ni\{B=b\}\big{\}}\leq\infty$, that is, ${W}^{D}_{+}\perp{W}^{D}_{-}$ and ${W}^{D}_{+}|_{t}\sim{W}^{D}_{-}|_{t}$, $\forall t<T^{D}_{\mathcal{B}}$.
+As such, data $\{{D}_{t}\}$ may resolve $B$ in finite time (see Appendix A for the construction of such processes from standard diffusion).
+Model-risk $B$ is thus active on $0,T^{D}_{\mathcal{B}}\wedge T)$.
     Remark 3.
-    Existing studies ignore potential structures within dataflow ($\{\mathbf{D}_{t}\}$), as if only data relevant to tangible firmed-up value, $\{Z_{t}\}$, exist/matter.
-By considering data exclusively relevant to model-risk (irrelevant once model-sure), we tackle a frequent challenge facing partitioners: a crisis market unsettled by a bout of model-risk whose economics dominate asset valuation and hence traders’ cognitive and trading activities, while non-model, routine, risks remain, independently evolving or hardly changing.
-  4. 4.
-[…]
+    Existing studies assume, de facto, that only data feeding directly into tangible, firmed-up, values exist/matter.
+By adding data exclusive to model-risk and irrelevant if model-sure, we may tackle a frequent challenge facing partitioners: a market unsettled by a bout of model-risk whose economics dominate asset value and so market activities, while routine risks remain, independently evolving, in the background.
   5. 5.
-    Tests based on likelihood-ratios (LR) $L^{+/-}_{t}:=\frac{d\mathbf{W}^{T}_{+}|_{t}}{d\mathbf{W}^{T}_{-}|_{t}}$, or log-LR $l^{+/-}_{t}:=\log L^{+/-}_{t}$, $\forall t<T$, are optimal in the sense of Neyman-Pearson Lemma.
+    Combined, asset data $\{\mathbf{D}_{t}\}:=\{(Z_{t},D_{t})\}$ have $B$-sure state space $\mathcal{S}^{T}:=(\mathcal{S}^{Z}\times\mathcal{S}^{D})|_{T}$ and $B$-sure law $\mathbf{W}^{T}_{B}:=\mathbf{W}_{B}|_{T}$, $\mathbf{W}_{B}:={W}^{Z}_{B}\times{W}^{D}_{B}$.
+Overall model-risk inference takes place on filtered space ($)\big{(}\mathcal{S}^{T}_{\mathcal{B}}:=\mathcal{B}\times\mathcal{S}^{T}\ni\omega_{B}:=(B,\omega),\text{ }\{\mathbf{F}_{t}\}:=\{\mathcal{F}^{Z}_{t}\}\vee\{\mathcal{F}^{D}_{t}\},\text{ }\mathbf{F}_{T};\text{ }{\pi}^{B(\cdot)}_{0}\mathbf{W}^{T}_{B(\cdot)}$($\cdot$$\big{)}$, where ${\pi}^{B(\omega_{B})}_{0}$ is the unconditional belief in outcome $B(\omega_{B}):=B$.
+Then, by Item-1&4, model-risk $B$ is active on $0,T_{\mathcal{B}}\wedge T)$, with $T_{\mathcal{B}}:=\inf\big{\{}t|\mathbf{F}_{t}\ni\{B=b\}\big{\}}={T}^{D}_{\mathcal{B}}$.
+  6. 6.
+    Tests based on likelihood-ratios (LR) $\{L^{+/-}_{t}\}:=\{\frac{d\mathbf{W}^{T}_{+}|_{t}}{d\mathbf{W}^{T}_{-}|_{t}}\}$, or log-LR $\{l^{+/-}_{t}\}:=\{\log L^{+/-}_{t}\}$, are optimal in the sense of Neyman-Pearson Lemma.
 Model-risk $B$ resolves iff.
@@
 Henceforth let $\mathbf{W}_{-}$ be this base measure, under which $r^{Z}_{-,t}=0=r^{D}_{-,t}$, $\forall t$.
-  6. 6.
+  7. 7.
     The above has a Bayesian representation: any $L^{+/-}_{t}\in(0,\infty)$ maps bijectively and smoothly to an a posteriori belief ${\pi}^{B}_{t}\in(0,1)$.
@@
 That is, inferential odds are exponentiated log-LR processes; the latter, $\{l^{+/-}_{t}\}$, are Wiener processes under broad conditions ((A.2-A.3)).
-  7. 7.
-    Given $Z_{t}=z_{t}$, $t<T$, the $B$-sure expected payoff $Y^{B}_{t}:={\mathbf{E}}_{B}[Y|\mathbf{F}_{t}]\equiv{\mathbf{E}}_{B}[Y|\mathcal{F}^{Z}_{t}]$ reads ((1-2)):
-    $Y^{B}_{t}=z_{t}+\mathbf{E}_{B}[Y-z_{t}|{\mathcal{F}^{Z}_{t}}]=z_{t}+\mathbf{E}_{B}[Y-z_{t}]=z_{t}+y_{B}(T-t,t),$ (4)
-    where $y_{B}(T-t,t)-y_{B}(T-t-t^{\prime},t+t^{\prime})=\mathbf{E}_{B}[\int_{t}^{t+t^{\prime}}dZ_{s}(B)]$, $\forall t^{\prime}\leq T-t$, and so $y_{B}(T-t,t)=\int_{t}^{T}{r}^{Z}_{B,s}ds$, as by definition $\{Y^{B}_{t}\}$ is a martingale and $y_{B}(0,T)\equiv 0$.
-Note also that this relates to model-risk impact $Y^{\Delta}(t)$ via $Y^{\Delta}(t)\equiv Y^{+}_{t}-Y^{-}_{t}\equiv{y}_{+}(T-t,t)-{y}_{-}(T-t,t)$.
 1.2 The FTAP Baseline under Model-Risk
@@
 where $\Sigma_{B}{(\cdot)^{B}}{(\cdot)^{B}}$ denotes averages weighted by $B$-outcomes, such as by RNE beliefs $\hat{\pi}^{B}_{t}$ under Bayes’ Rule (4) given $\hat{\pi}^{B}_{0}$ and RNE log-LR $\hat{l}^{+/-}_{t}:=\log\frac{d\hat{\mathbf{W}}^{T}_{+}|_{t}}{d\hat{\mathbf{W}}^{T}_{-}|_{t}}$.
-Viability by no means ensures economic validity (e.g.
-any pricing based on RNE measures of the form $\tilde{\pi}^{B}_{0}\tilde{\mathbf{W}}^{T}\sim{\pi}^{B}_{0}\mathbf{W}_{B}^{T}$ is blind to $B$-informative data); indeed, any pricing that is viable to any member of the equivalence class ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}]$ is viable to all in the class.
-In this sense, ’anything goes’ under FTAP-viability.
-What characterises the processes and the RNE measures that do correspond to plausible economic theories?
-For answers, consider risk-premium, the gap between ex-ante pricing and expectation, non-negative under economic asset-pricing: at $\forall t<T$, given reference model-risk belief $\{{\pi}^{B}_{t}\}$,
-$RP_{t}:=Y_{t}-S_{t}=\Sigma_{B}{\pi}^{B}_{t}Y^{B}_{t}-S_{t}\geq 0.$ (7)
+Remark 4.
+By setup, $\hat{S}^{B}_{t}=\hat{\mathbf{E}}_{B}\big{[}Y\big{|}\mathcal{F}^{Z}_{t}\big{]}=z_{t}+\hat{y}_{B}(T-t,t)$ given $Z_{t}=z_{t}$, as in (2).
+Thus, $B$-risk, which can be measurable (resolvable) to $\mathbf{F}_{T}$ but not $\mathcal{F}^{Z}_{T}$ (Item-1), is not measurable to $\mathcal{F}^{S}_{T}$, the price-generated $\sigma$-algebra via (5).
+It is not replicable and cannot be priced by arbitrage.
+Viable prices do not mean economically valid prices.
+Indeed, any pricing that is viable to any one member of the equivalence class ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}]$ is then viable to all in the class; ’anything goes’ in this sense.
+What characterise processes and RNE measures that stem from valid economic theories?
+For answers, consider risk-premia $\{RP_{t}\}$, the gap between pricing and expectation, which should be non-negative, at least ex-ante: given reference model-risk belief $\{{\pi}^{B}_{t}\}$,
+$RP_{t}:=Y_{t}-S_{t}=\Sigma_{B}{\pi}^{B}_{t}Y^{B}_{t}-S_{t}\geq 0\text{, }\forall t<T.$ (7)
 It has RNE decomposition (by inserting $\pm\Sigma_{B}\hat{\pi}^{B}_{t}{{Y}^{B}_{t}}$):
@@
 $B\_\widehat{RP}_{t}:=RP_{t}-\Sigma_{B}\hat{\pi}^{B}_{t}{\hat{RP}^{B}_{t}}=({\pi}^{+}_{t}-\hat{\pi}^{+}_{t}){Y}^{\Delta}(t),$ (10)
-where $\Sigma_{B}\hat{\pi}^{B}_{t}{\hat{RP}^{B}_{t}}$ accounts for $B$-sure (non-model) risks, and $B\_\widehat{RP}_{t}$, model-risk $B$.
-The latter has standard deviation $Y^{\Delta}(t)\sigma^{\pi}_{t}$, $\sigma^{\pi}_{t}:=(\pi^{+}_{t}\underline{\pi}^{+}_{t})^{\frac{1}{2}}$, which, being free of risk-pricing choices, is an intrinsic quantity.
-It well-defines a RNE price-of-$B$-risk:
+where $\Sigma_{B}\hat{\pi}^{B}_{t}{\hat{RP}^{B}_{t}}$ discounts $B$-sure (non-model) risks, and $B\_\widehat{RP}_{t}$, model-risk $B$.
+The latter has standard deviation $Y^{\Delta}(t)\sigma^{\pi}_{t}$, $\sigma^{\pi}_{t}:=(\pi^{+}_{t}{\pi}^{-}_{t})^{\frac{1}{2}}$, which, being free of risk-pricing choices, is an intrinsic quantity.
+It well-defines a RNE price-of-$B$-risk process $\{k_{t}^{\hat{\pi}}\}$: $\forall t<T$,
 ${k}^{\hat{\pi}}_{t}:=\frac{B\_\widehat{RP}_{t}}{Y^{\Delta}(t)\sigma^{\pi}_{t}}\equiv\frac{{\pi}^{+}_{t}-\hat{\pi}^{+}_{t}}{(\pi^{+}_{t}\underline{\pi}^{+}_{t})^{\frac{1}{2}}}.$ (11)
@@
 $RP^{B}_{t}:=Y^{B}_{t}-S^{B}_{t}=y_{B}(T-t,t)-\check{y}_{B}(T-t,t)=:RP^{B}(t)\geq 0,$ (13)
-where $\check{\mathbf{E}}_{{B}}[\cdot|\cdot]$ denotes expectations under its $B$-sure RNE measure $\check{\mathbf{W}}^{T}_{B}\sim{\mathbf{W}}^{T}_{B}$, leading to RNE outlook $\check{y}_{B}(T-t,t)=\int_{t}^{T}\check{r}_{B,s}ds$ as in (4), and $B$-sure risk-premium $RP^{B}(t)$ satisfying: $\forall t\leq T$,
+where $\check{\mathbf{E}}_{{B}}[\cdot|\cdot]$ denotes expectations under its $B$-sure RNE measure $\check{\mathbf{W}}^{T}_{B}\sim{\mathbf{W}}^{T}_{B}$, leading to RNE outlook $\check{y}_{B}(T-t,t)={\mathbf{E}}[\int_{t}^{T}\check{r}_{B,s}ds]$ as in (2), and $B$-sure risk-premium $RP^{B}(t)$ satisfying: $\forall t<T$,
 $\check{y}_{+}(T-t,t)-\check{y}_{-}(T-t,t)=:S^{\Delta}(t)>0,\text{ almost surely},$ (14)
 $S^{B}_{t}=\lim_{\sigma^{\pi}_{t}\to 0}\hat{S}^{B}_{t},\text{ }\text{i.e. }RP^{B}(t)=\lim_{\sigma^{\pi}_{t}\to 0}\hat{RP}^{B}(t),$ (15)
-where (14) ensures consistency with asset-property Item-2, Section 1.1, while (15), the internal consistency of the pricing scheme under model-risk $B$.
+where (14) ensures consistency with asset-property Item-3, Section 1.1, while (15), the internal consistency of the pricing scheme under model-risk $B$.
 Remark 5.
@@
 $B\_RP_{t}:=RP_{t}-\Sigma_{B}{\pi}^{B}_{t}{RP^{B}(t)}=\Sigma_{B}\pi^{B}_{t}S^{B}_{t}-S_{t}=({\pi}^{+}_{t}-A^{+}_{t}){S}^{\Delta}(t),$ (18)
-$k^{A}_{t}=\frac{{\pi}^{+}_{t}-A^{+}_{t}}{(\pi^{+}_{t}\underline{\pi}^{+}_{t})^{\frac{1}{2}}}\geq 0,$ (19)
+$k^{A}_{t}:=\frac{{\pi}^{+}_{t}-A^{+}_{t}}{(\pi^{+}_{t}\underline{\pi}^{+}_{t})^{\frac{1}{2}}}\geq 0,$ (19)
 which parallels RNE decomposition (8-11), but pricing coefficients $\{A^{B}_{t}\}\in(0,1)$, unlike $\{\hat{\pi}^{B}_{t}\}$, need not be a conditional-probability process.
@@
 Remark 6.
-In ambiguity-aversion settings, the primary utility function, enforcing risk-aversion, and the secondary utility/variational-control, enforcing uncertainty-aversion, can be dialled independently and arbitrarily close to neutrality (linearity), to achieve the above separation.
+In ambiguity-aversion settings, the primary utility function, enforcing risk-aversion, and the secondary-utility/variational-control, enforcing uncertainty-aversion, can be dialled independently and arbitrarily close to neutrality (linearity), to achieve the above separation.
 In classical and some parameter-learning settings, with a single utility, the above obtains by replacing payoffs with their model-sure expectations perturbed by suitably small noise.
@@
 Asset-pricing under model-risk $B$ is said to be usual if its price-of-$B$-risk $k^{A}_{t}$, (19), is a time-homogenous twice-differentiable function of its model-risk belief [^3] Economic risk-pricing, as the solution to some optimisation scheme, is measurable to the natural filtration of its input $(\{Z_{t}\},\{\pi^{B}_{t}(\mathbf{D}_{t})\})$; so it is at any time $t$ a function of $(Z_{t},\pi^{B}_{t}(\mathbf{D}_{t}))$ (Proposition 4.9 of Breiman (1992)).
-Any explicit dependence on $t$ or $Z_{t}$, if at all, is presumed here weak or non-existent, at least for $t\ll T$.
+Any explicit dependence on $Z_{t}$ or time, if at all, is presumed weak or non-existent, at least for $t\ll T$.
 $\pi^{B}_{t}$, $\forall t\ll T$.
@@
 The difference process $\{{\pi}^{+}_{t}\}-\{{\Pi}^{+}_{t}\}=\{k^{\Pi}_{t}\sigma^{\pi}_{t}\}$ has a fixed sign.
-Unless the model-risk is consistent (see Item-2, Section 1.1, and (14)), viable model-risk premia $\{B\_RP_{t}\}$ (see (18)) must be negative sometimes and so correspond to no known economic risk-pricing theories.
+Unless the model-risk is consistent (see Item-3, Section 1.1, and (14)), viable model-risk premia $\{B\_RP_{t}\}$ (see (18)) must be negative sometimes and so correspond to no known economic risk-pricing theories.
 2.2.2 Connection with Familiar Price-Dynamics
-For better focus, let any $B$-sure risk-premia be $B$-independent henceforth, to be written as $\check{rp}(t)=\int_{t}^{T}\check{R}_{s}ds>0$, with $\check{R}_{t}$ denoting the $B$-sure risk-premium drift at any $t<T$; note then $Y^{\Delta}(t)=S^{\Delta}(t)=\int_{t}^{T}r^{Z\Delta}_{s}ds$ as a result.
-Canonical pricing (22) becomes:
+For better focus, let any $B$-sure risk-premia be indifferent to $B$-value henceforth, to be written as $\check{rp}(t)={\mathbf{E}}[\int_{t}^{T}\check{R}_{s}ds]>0$, with $\check{R}_{t}$ denoting the $B$-sure risk-premium drift at any $t<T$; note then $Y^{\Delta}(t)=S^{\Delta}(t)$ as a result, Canonical pricing (22) becomes:
 $S_{t}=S^{(0|\Pi)}_{t}-\check{rp}(t)=Y^{-}_{t}+S^{\Delta}(t)\Pi^{+}_{t}-\check{rp}(t),$ (24)
-$dS_{t}(B)=dY^{-}_{t}(B)+S^{\Delta}(t)d\Pi^{+}_{t}(B)-r^{Z\Delta}_{t}\Pi^{+}_{j}(B)dt+\check{R}_{t}dt.$ (25)
-By setup (Section 1.1) and inference dynamics (A.2-A.3), process $\{d\Pi^{+}_{t}\}$ has two sources of noise, one from $\{Z_{t}\}$, given by $\{\sigma^{lZ}_{t}dw^{Z}_{t}\}$, with $\sigma^{lZ}_{t}:=r^{Z\Delta}_{t}/\sigma^{Z}_{t}\ll 1$ (’signal-to-noise’), one from independent $B$-informative data $\{D_{t}\}$, modelled in parallel ((2)), which can be absorbed through $(\sigma^{l}_{t})^{2}:=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$ and $\sigma^{l}_{t}dw_{t}:=\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}$; then:
+$dS_{t}(B)=dY^{-}_{t}(B)+S^{\Delta}(t)d\Pi^{+}_{t}(B)-r^{Z\Delta}_{t}\Pi^{+}_{t}(B)dt+\check{R}_{t}dt.$ (25)
+By setup (Section 1.1) and inference dynamics (A.2-A.3), process $\{d\Pi^{+}_{t}\}$ has two sources of noise, one from $\{Z_{t}\}$ ((1)), given by $\{\sigma^{lZ}_{t}dw^{Z}_{t}\}$, with $\sigma^{lZ}_{t}:=r^{Z\Delta}_{t}/\sigma^{Z}_{t}\ll 1$ (’signal-to-noise’), one from independent $B$-informative data $\{D_{t}\}$ ((3)), modelled likewise, which can be absorbed through $(\sigma^{l}_{t})^{2}:=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$ and $\sigma^{l}_{t}dw_{t}:=\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}$.
+Therefore:
 $d\Pi^{+}_{t}(B)=\mu^{\Pi,B}_{t}dt+\sigma^{\Pi,l}_{t}dw_{t},\text{ with},$ (26)
@@
 The above has a drift by design relative to reference belief $\{{\pi}^{+}_{t}\}$ (with dynamics (A.3)):
-$\mu^{\Pi,\pi}_{t}:=\Sigma_{B}{\pi}^{B}_{t}{\mu^{\Pi,B}_{t}}=\sigma^{\Pi,l}_{t}\sigma^{l}_{t}\cdot(\pi^{+}_{t}-\Pi^{+}_{t})=\sigma^{\Pi,l}_{t}\sigma^{l}_{t}\cdot k^{\Pi}_{t}\sigma^{\pi}_{t}\geq 0;$ (29)
+$\mu^{\Pi,\pi}_{t}:=\Sigma_{B}{\pi}^{B}_{t}{\mu^{\Pi,B}_{t}}=\sigma^{\Pi,l}_{t}\sigma^{l}_{t}\cdot(\pi^{+}_{t}-\Pi^{+}_{t})=(\sigma^{\Pi}_{t}\sigma^{l}_{t})^{2}\cdot k^{\Pi}_{t}\sigma^{\pi}_{t}\geq 0;$ (29)
 note ${\mu^{\pi,\pi}_{t}}\equiv 0\equiv{\mu^{\Pi,\Pi}_{t}}$, reflecting the martingale property of inference and viable pricing.
@@
 $dS_{t}(B)=$ $\big{[}\check{R}_{t}dt+\sigma^{Z}_{t}dw^{Z}_{t}\big{]}$ (30)
-$+$ $\big{[}\big{(}\mathbf{1}_{+}(B)-\Pi^{+}_{j}(B)\big{)}\cdot r^{Z\Delta}_{t}dt\big{]}$ (31)
+$+$ $\big{[}\big{(}\mathbf{1}_{+}(B)-\Pi^{+}_{t}(B)\big{)}\cdot r^{Z\Delta}_{t}dt\big{]}$ (31)
 $+$ $S^{\Delta}(t)(\sigma^{\Pi}_{t})^{2}\big{[}\big{(}\mathbf{1}_{+}(B)-\Pi^{+}_{t}(B)\big{)}\cdot(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw_{t}\big{]}.$ (32)
 Remark 10.
-Practical cases often see $B$-sure dynamics (30) and model-drift (31) dominated by model-inference (32): $\check{R}_{t}\ll S^{\Delta}(t)$ and $r^{Z\Delta}_{t}\ll S^{\Delta}(t)$; indeed firmed-up values $\{Z_{t}\}$ are often nil or fixed so that asset-pricing is just a matter of model-risk inference, with $\{dS_{t}\}$ reduced to the inferential test equation of Peskir and Shirayev (2006) (for constant $\{r^{D\Delta}_{t}\}$ and $\{\sigma^{D}_{t}\}$).
-Further, it often holds that relative variance $\text{F}^{Z/D}_{t}:=(\sigma^{lZ}_{t})^{2}/(\sigma^{lD}_{t})^{2}\ll 1$ and so $(\sigma^{l}_{t})^{2}\approx(\sigma^{lD}_{t})^{2}$: model-risk is the prime price-mover, not tangible value.
-Dispersion among such assets is thus driven mainly by their differing underlying model-risk inference processes.
+Inference dynamics (32) usually dominate $B$-sure dynamics (30) and model-drift (31), as $\check{R}_{t}\ll S^{\Delta}(t)$ and $r^{Z\Delta}_{t}\ll S^{\Delta}(t)$.
+Indeed, during gaps between payoff updates, which exist in reality, our asset-price equation of motion reduces to (32), essentially the sequential test process of Peskir and Shirayev (2006).
+Further, it often holds that $(\sigma^{l}_{t})^{2}\approx(\sigma^{lD}_{t})^{2}$ (relative variance $\text{F}^{Z/D}_{t}:=(\sigma^{lZ}_{t})^{2}/(\sigma^{lD}_{t})^{2}\ll 1$), so that model-risk inference, not tangible value, is the main driver of price moves and dispersions among the assets affected.
+Finally, under model-risk $B$, even if unpriced (so the inference-induced drift (29) is nil), asset-price volatility rises regardless, unless asset-pricing is already $B$-sure (i.e.
+$(\sigma^{\Pi}_{t})^{2}=0$).
 Averaged under model-risk belief $\{{\pi}^{+}_{t}\}$, dominant price-dynamic (32) has this familiar form:
@@
 $\frac{\mu_{t}}{\sigma_{t}}=(K^{\frac{1}{2}}-K^{-\frac{1}{2}})\sigma^{\Pi}_{t}\sigma^{\pi}_{t}\sigma^{l}_{t}=\frac{K^{\frac{1}{2}}-K^{-\frac{1}{2}}}{S^{\Delta}(t)}\bigg{(}\frac{\sigma^{\pi}_{t}}{\sigma^{\Pi}_{t}}\bigg{)}\sigma_{t}.$ (34)
-Expansion in (K1)(K-1), given Remark 7, recovers the standard and ubiquitous price-of-risk vs price-volatility relationship:
+Expansion in (K1)(K-1), given Remark 8, recovers the standard and ubiquitous relationship:
 $\frac{\mu_{t}}{\sigma_{t}}\approx\bigg{(}\frac{K-1}{S^{\Delta}(t)}\bigg{)}\sigma_{t}\text{ and }\mu_{t}\approx\bigg{(}\frac{K-1}{S^{\Delta}(t)}\bigg{)}\sigma_{t}^{2}.$ (35)
-Remark 10.
-The proportionality factor characterises ex-ante pricing through an intuitive and direct metric of asset-pricing (Remark 7).
-How its level comes about is a matter of economic theory (e.g.
+The proportionality factor characterises risk-pricing via an intuitive metric, the intended gain-to-loss (Remark 8).
+How its level comes about is up to economic theories (e.g.
 risk-aversion).
-Regardless, our derivation links it firmly to ex-ante beliefs, making the above relevant to bias studies and measurement under model-uncertainty.
+Our derivation, linking it to ex-ante beliefs, makes it relevant to bias studies.
 3 Bias and Risk-Pricing Measurement via Price Anomalies
 3.1 The Joint-Hypothesis Issue
-To focus on main issues, as in Section 2.2.2, consider usual-case assets under model-risk $B$ with $B$-independent $B$-sure risk-premia $\{\check{rp}(t)\}$.
-Let $\{\mathbf{p}^{B}_{t}\}$ be the $\{\mathbf{F}_{t}\}$-adapted true conditional probabilities of model-risk outcomes, given the true and known $B$-sure laws $\mathbf{W}^{T}_{B}$ and the true but potentially unknown unconditional law $\mathbf{p}^{B}_{0}$ of model-risk $B$.
-Its ex-post price-drift to be realised on average $\{rp_{t}\}:=\{\Sigma_{B}\mathbf{p}^{B}_{t}{Y^{B}_{t}}\}-\{S_{t}\}$ satisfies:
+Take as in Section 2.2.2 usual-case assets with model-risk $B$ and $B$-indifferent $B$-sure risk-premia $\{\check{rp}(t)\}$, for better focus.
+Let $\{\mathbf{p}^{B}_{t}\}$ be the $\{\mathbf{F}_{t}\}$-adapted true conditional probabilities of $B$-values, given the true, known, $B$-sure laws $\mathbf{W}^{T}_{B}$, and the true, potentially unknown, unconditional law $\mathbf{p}^{B}_{0}$ of $B$.
+Its ex-post average price-drift $\{rp_{t}\}:=\{\Sigma_{B}\mathbf{p}^{B}_{t}{Y^{B}_{t}}\}-\{S_{t}\}$ satisfies:
 $rp_{t}-\check{rp}(t)=B\_RP_{t}+(\mathbf{p}^{+}_{t}-{\pi}^{+}_{t}){S}^{\Delta}(t),\text{ }t<T.$ (36)
@@
 from $B=+$ to $B=$-) with true probability $\mathbf{p}^{(\cdot)}_{0}$: denote the new state by ’1’, status quo, ’0’, and status quo model-sure pricing, $\{S^{\text{{\bf{0}}}}_{t}\}$.
-The triggering events may or may not follow known laws.
-The process of how/when they arise is ignored, assuming that the average time between model crises, say some $\lambda^{-1}\gg 1$, is so long that each can be considered one-off.
-The period of interest is right after the arrival of a trigger but long before the next, so $t\ll\lambda^{-1}$, and long before horizon/maturity, so $t\ll T$.
+We ignore the law governing triggering events by assuming that the average time between them, say $\lambda^{-1}\gg 1$, is so long that each is considered one-off.
+The period of interest is just after an arrival, long before the next, and long before horizon, so $t\ll\lambda^{-1}$ and $t\ll T$.
 Canonical asset-pricing (24), with the direction $sign[1{\text{{\bf{0}}}}]:=sign[Y^{1}_{t}-Y^{\bf{0}}_{t}]$ of potential change explicit and the $t$-dependence of impact $S^{\Delta}(t)$ dropped given $t\ll T$ (and so $S^{\Delta}(t)\gg r^{Z\Delta}_{t}$), reads:
@@
 If the change-risk is diversifiable, so unpriced, then $K=1$; if it is classical, so its law $\mathbf{p}^{1}_{0}$ is known, then $\rho=1$.
-Parameter $\rho$ stems from model-uncertainty and is thus unknown ex-ante.
+Parameter $\rho$ is born of model-uncertainty and thus unknown ex-ante.
 The joint-hypothesis problem takes a concrete form: observed prices reflect product (39), so external criteria seem necessary to separate ’intention’ $K$ from ’mistake’ $\rho$.
@@
 $\mathbf{p}^{1}_{0}\geq\frac{1}{2}$ and $\rho\gg 1$) would allow easy wins for ’bet on change’ and so should not persist.
-3.3 An Ideal Model-Uncertainty Driven Market
-Consider a model-uncertainty driven market, of a large number of assets, each facing some change-risk $B$, whose nature may vary from asset to asset (e.g.
-new CEO, tort, war) but whose true $B$-sure laws are known.
-Let the $B$-risks be priced (i.e.
-correlated with the market’s stochastic-discount-factor), for their effects on aggregate value/volatility/cross-sectional dispersion [^6] Dispersion within members of an equity index, as a priced factor, has been linked to market uncertainty, the business cycle and structural shifts (e.g.
+3.3 An Ideal Market Driven by Model-Uncertainty
+Consider an ideal market, of a large number $N\gg 1$ of usual-case assets, each facing a change-risk $B$, whose nature may vary from asset to asset (e.g.
+new CEO, tort) but whose $B$-sure laws are known.
+Let them be priced-risks (i.e.
+correlated with the stochastic-discount-factor) of the market, as they can elevate total volatility and cross-sectional dispersion [^6] Dispersion among members of an equity index, as a priced factor, has been linked to market uncertainty, the business cycle and structural shifts (e.g.
 Demirer and Jategaonkar (2013), Kolari et al.
-(2021))..
-Further, let them be economically priced, so that by Proposition 1 their pricing is canonical with respect to some reference inferential belief process of the relevant $B$-risk.
-Finally, let the realisation and detection of individual change-outcomes be stochastic and independent (even for macroeconomic triggers such as war or recession).
-Note that such a market, when seen as a whole, is an asset with non-binary model-uncertainty; this is so even when there is only one binary macroeconomic model-risk trigger.
+(2021))., irrespective of their impact on total value (see (32) and Remark 10).
+Let them be economically priced, so that Section 2 applies (to individual asset processes).
+Note that the market as a whole has a non-binary model-risk: even if there is just one binary trigger (e.g.
+war, recession), there can be as many as $2^{N}$ market models.
+Let the realisation and detection of each individual change-outcome be stochastic and independent, even for a macroeconomic trigger.
 For brevity, let the unconditional all-cause probability of positive/negative economic change be symmetrical: at any time half of all potential changes, of any origin, are positive/negative (unconditional asymmetry is addressed in Remark 14&15).
 Let ’pricing intention’ $K$ ((39)) and ’apparent bias’ $\rho$ ((38)) be indifferent to the economic direction and origin of change-risk, and the size of $K$, $\rho$ and risk-impact $S^{\Delta}$, uniform across the assets.
-Such uniformity in practice amounts to asset dispersions being driven by differences in the paths of model-risk inference rather than in any other aspect (Remark 9).
+Such uniformity in practice amounts to asset dispersion being driven primarily by differences in the evolutions of model-risk inference rather than in any other aspect (Remark 10).
 Similarly, let all risk-inference be based on log-LR processes that are homogenous in time and uniform across.
@@
 Momentum applies also to $\rho\ll 1$: past excess due to RNE beliefs $\{\Pi^{1}_{t}\}$ falling from $\Pi^{1}_{0}\approx 1$ to $v<\Pi^{1}_{0}$ means an average future excess of the same sign; indeed, label-switching defines a new parameter $\rho^{-1}>1$.
-Momentum thus vanishes only when $\rho=1$ (e.g.
-when there is no model-uncertainty).
+Momentum vanishes only if $\rho=1$ (no bias or model-uncertainty).
 Momentum trading (40) earns its best realisable excess $rp^{\pm mo}_{max}$ at $v^{\pm mo}_{max}$:
@@
 $v^{\sigma}_{max}:=\frac{1}{\rho+1},\text{ }rp^{\sigma}_{max}:=\frac{1}{2}\frac{K-1}{K+1}S^{\Delta}.$ (43)
-At $\rho=1$ the above is exact, by which $v^{\sigma}_{max}=\frac{1}{2}$, where price-volatility peaks, thus confirming ’high risk high reward’ in the absence of model-uncertainty or bias.
-There is a low-risk effect otherwise: peak-reward occurs at $v^{\sigma}_{max}\approx\rho^{-1}$, while peak volatility ($v=\frac{1}{2}$) is poorly rewarded ex-post: $rp^{\sigma}(\frac{1}{2})=\frac{S^{\Delta}}{2}\mathcal{O}(\frac{K-1}{\rho}-1)$.
+At $\rho=1$ it is exact and $v^{\sigma}_{max}=\frac{1}{2}$ (peak volatility), thus confirming ’high risk high reward’ in the absence of model-uncertainty or bias.
+Otherwise a low-risk effect applies: peak-reward occurs at $v^{\sigma}_{max}\approx\rho^{-1}$, and peak volatility ($v=\frac{1}{2}$) is poorly rewarded ex-post: $rp^{\sigma}(\frac{1}{2})=\frac{S^{\Delta}}{2}\mathcal{O}(\frac{K-1}{\rho}-1)$.
 Note the separation: bias $\rho$ is revealed by peak-location $v^{\sigma}_{max}$, and risk-pricing $K$, by peak-reward $rp^{\sigma}_{max}$, which is exactly that set ex-ante for peak-risk (Remark 8).
@@
 4 Summary and Discussion
-We study the risk-neutral equivalent formulation of model-risk pricing. Explicit price dynamics under binary model-risks with potentially unknown laws are derived.
-The results show that model-risk can be a main source of excess value, with a key role in Status Quo Bias and anomalies such as Momentum and Low-Risk.
-They offer an integrated and practicable approach to these highly relevant and interconnected topics.
+We propose a risk-neutral equivalent formulation of model-risk pricing.
+Explicit price dynamics under binary model-risks with potentially unknown laws are derived.
+One implication is that model-risk can be a main source of excess value, with a key role in Status Quo Bias and anomalies such as Momentum and Low-Risk.
+Our approach offers an integrated and practicable framework to study these highly relevant and interconnected subjects.
 We arrive at the above by making use of the informational constraint imposed by ’usual’ economic asset-pricing (Remark 1 and Definition 1), to show that viable asset-pricing under model-risk may take just one canonical form (Lemma 1 and Proposition 1), which leads to a unique and intuitive model-risk pricing formula with convenient properties (Corollary 1).
-The sole risk-pricing parameter is associated with a conserved quantity of inferential dynamics and is a product of ’intention’ (preference) and ’mistake’ (difference between objective and subjective laws).
+The sole risk-pricing parameter derives from a conserved quantity of inference dynamics and is a product of ’intention’ (preference) and ’mistake’ (difference of objective and subjective laws).
 Ex-ante (’intended’) pricing, classical or with ambiguity-aversion/robust-control, can be captured by our parameter $K\in[1,\infty)$, (39), with $K=1$ corresponding to ’no model-risk discounts’.
@@
 Asset-prices under max-min utility already have this characteristic, albeit with a built-in pessimism absent in inertial pricing.
-Both are viable in our setup, as ’models’ are represented by probability laws that are equivalent before any risk-resolution (that is predictable).
-The theoretical and practical aspects of this topic will be examined in a follow-up study.
+Both are viable in our setup, as ’models’ are represented by probability laws that are equivalent before any risk-resolution.
+The theoretical and practical aspects of inertial pricing will be examined in a follow-up study.
 APPENDIX
@@
     Standard Continuous-Time Inference.
-Given setup (Section 1.1), the log-LR process for testing time-homogeneous Wiener data of drift $\mathbf{1}_{+}(B)r^{Z}$ and volatility $\sigma^{Z}$ has the form (Peskir and Shirayev (2006)): in terms of signal-to-noise $\sigma^{l}=\frac{r^{Z}}{\sigma^{Z}}$,
+Given setup (Section 1.1), the log-LR process for testing time-homogeneous Wiener data of drift $\mathbf{1}_{+}(B)r^{Z}$ and volatility $\sigma^{Z}$ has the following form (Peskir and Shirayev (2006)): in terms of signal-to-noise $\sigma^{l}=\frac{r^{Z}}{\sigma^{Z}}$,
     $dl^{+/-}_{\tau}(B)=(-1)^{\mathbf{1}_{-}(B)}\frac{(\sigma^{l})^{2}}{2}d\tau+\sigma^{l}dw_{\tau},$ (A.1)
-    with $\{w_{\tau}\}$ being some standard Wiener process.
-Time-dependence obtains via an absolutely continuous clock-change $t\mapsto\tau(t)$: image $\{l^{+/-}_{t}\}:=\{l^{+/-}_{\tau(t)}\}$ is Wiener, with continuous time-varying parameters (e.g.
+    with $\{w_{\tau}\}$ being a standard Wiener process.
+Time-dependence obtains via an absolutely continuous clock-change $t\mapsto\tau(t)$: the image $\{l^{+/-}_{t}\}:=\{l^{+/-}_{\tau(t)}\}$ is Wiener, with continuous time-varying parameters (e.g.
 Kallsen (2006)).
-It and the resulting belief process by Bayes’ Rule (3) read: in terms of a standard Wiener process $\{w_{t}\}$ in $t$-time,
+It and the resulting belief process by Bayes’ Rule (4) then read: in terms of a standard Wiener process $\{w_{t}\}$ in $t$-time,
     $dl^{+/-}_{t}(B)=(-1)^{\mathbf{1}_{-}(B)}\frac{(\sigma^{l}_{t})^{2}}{2}+\sigma^{l}_{t}dw_{t};$ (A.2)
@@
     Non-Resolution and Novikov’s Condition.
-For inference based on equivalent Wiener-measure pair ${\mathbf{W}}_{+}\sim\mathbf{W}_{-}$, the Radon-Nikodym Theorem ensures $\lim_{t\to\infty}\log\frac{d{\mathbf{W}}_{+}|_{t}(\cdot)}{d\mathbf{W}_{-}|_{t}(\cdot)}<\infty$ and so a finite total variance for the log-LR process (A.2) almost surely.
-Risk-resolution (Item-5, Section 1.1) thus equates to the LR process, the exponentiation of (A.2), or equivalently (A.3), violating Novikov’s Condition, and vice versa.
+For inference using equivalent Wiener-measure pair ${\mathbf{W}}_{+}\sim\mathbf{W}_{-}$, the Radon-Nikodym Theorem ensures $\lim_{t\to\infty}\log\frac{d{\mathbf{W}}_{+}|_{t}(\cdot)}{d\mathbf{W}_{-}|_{t}(\cdot)}<\infty$ and so a finite total variance for the log-LR process (A.2) and non-resolution.
+Risk-resolution (Item-6, Section 1.1) thus equates to the LR process, the exponentiation of (A.2), or equivalently (A.3), violating Novikov’s Condition, and vice versa.
   3. 3.
     Regular Inference.
-It resolves $B$-outcomes predictably and so continuously.
-With i.i.d data risk-resolution occurs only at $T_{\mathcal{B}}=\infty$.
-Finite-time resolution can be modelled via a clock-change that brings resolution forward.
-Note that log-LR processes (A.2) diverge as $t\to T_{\mathcal{B}}\leq\infty$ while the associated beliefs (A.3) stay finite and continuous.
-For assets with horizon or maturity $T<T_{\mathcal{B}}$ or with $T_{\mathcal{B}}\leq T<\infty$ but with model-risk $B$ not $\mathcal{F}^{Z}_{T}$-measurable, our model-risk pricing applies all the same on $0,T_{\mathcal{B}}\wedge T)$ regardless.
-In the rest of the cases, where model-risk $B$ is $\mathcal{F}^{Z}_{T}$-measurable and hence replicable/hedgeable, its ’economic pricing’ (Remark 1) becomes irrelevant anyway.
+Uncertainty resolves predictably and so continuously under regular inference.
+With i.i.d data resolution occurs at $T_{\mathcal{B}}=\infty$.
+It can be brought forward under a clock-change that transforms processes on $(0,\infty)$ to those on some $(0,T_{\mathcal{B}}\leq\infty)$.
+Note that log-LR processes (A.2) diverge as $t\to T_{\mathcal{B}}$ while the associated beliefs (A.3) stay finite and continuous.
+For assets with horizon or maturity $T<T_{\mathcal{B}}$, or with $T_{\mathcal{B}}\leq T<\infty$ but a model-risk $B$ that is not $\mathcal{F}^{Z}_{T}$-measurable, our model-risk pricing theory applies all the same on $0,T_{\mathcal{B}}\wedge T)$.
+In all other cases, where model-risk $B$ is $\mathcal{F}^{Z}_{T}$-measurable and hence replicable/hedgeable, arbitrage pricing applies.
   4. 4.
@@
 […]
-Under reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$, with $d\mathbf{W}^{T}_{-,t}=\sigma^{Z}_{t}dw^{Z}_{t}+\sigma^{D}_{t}dw^{D}_{t}$, $\{dw^{Z}_{t}\}$ and $\{dw^{Z}_{t}\}$ being independent standard Wiener processes (Item-3&5, Section 1.1), we have:
+Under reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$, with $d\mathbf{W}^{T}_{-,t}=\sigma^{Z}_{t}dw^{Z}_{t}+\sigma^{D}_{t}dw^{D}_{t}$, $\{dw^{Z}_{t}\}$ and $\{dw^{D}_{t}\}$ being independent standard Wiener processes (Item-4&6, Section 1.1), we have:
 […]
 $(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)(A^{+}_{t})^{\prime}_{\pi}\bigg{[}($ $\mathbf{1}_{+}-\pi^{+}_{t})\cdot(\sigma^{l}_{t})^{2}dt+(\sigma^{\pi}_{t})^{2}\frac{(A^{+}_{t})^{\prime\prime}_{\pi}}{2(A^{+}_{t})^{\prime}_{\pi}}\cdot(\sigma^{l}_{t})^{2}dt\text{ }+\text{ }\sigma^{l}_{t}dw_{t}\bigg{]},$ (B.2)
-with model-drift $r^{Z\Delta}_{t}dt=-{d}Y^{\Delta}(t)$ (Item-2, Section 1.1) and $(\sigma^{l}_{t})^{2}:=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$, where $\sigma^{lZ}_{t}:={r^{Z\Delta}_{t}}/{\sigma^{Z}_{t}}$ (’signal-to-noise’) stems from data $\{Z_{t}(B)\}$ (Peskir and Shirayev (2006)), with $\sigma^{lD}_{t}$ likewise from $B$-conditionally independent data $\{D_{t}(B)\}$, and so $\sigma^{l}_{t}dw_{t}:=\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}$.
+with model-drift $r^{Z\Delta}_{t}dt=-{d}Y^{\Delta}(t)$ (Item-3, Section 1.1) and $(\sigma^{l}_{t})^{2}:=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$, where $\sigma^{lZ}_{t}:={r^{Z\Delta}_{t}}/{\sigma^{Z}_{t}}$ (’signal-to-noise’) stems from data $\{Z_{t}(B)\}$ (Peskir and Shirayev (2006)), with $\sigma^{lD}_{t}$ likewise from $B$-conditionally independent data $\{D_{t}(B)\}$, so $\sigma^{l}_{t}dw_{t}:=\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}$.
 Without risk-pricing, $\{A^{+}_{t}\}=\{\pi^{+}_{t}\}$, we have trivial viability under reference belief $\pi^{B}_{0}\mathbf{W}^{T}_{B}$:
 $dS^{(0|\pi)}_{t}=\big{[}(\mathbf{1}_{+}-\pi^{+}_{t})\cdot r^{Z\Delta}_{t}dt+\sigma^{Z}_{t}dw^{Z}_{t}\big{]}+(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)\big{[}(\mathbf{1}_{+}-\pi^{+}_{t})\cdot(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw_{t}\big{]}.$ (B.3)
-With viable model-risk only pricing, under some RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}\sim\pi^{B}_{0}\mathbf{W}_{B}^{T}$, we have $\{S^{(0|A)}_{t}\}=\{\Sigma_{B}\hat{\pi}^{B}_{t}\hat{S}^{B}_{t}\}$, and the following, by (8-11) and (19),
+With viable model-risk only pricing, under a RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}\sim\pi^{B}_{0}\mathbf{W}_{B}^{T}$, we have $\{S^{(0|A)}_{t}\}=\{\Sigma_{B}\hat{\pi}^{B}_{t}\hat{S}^{B}_{t}\}$, and the following, by (8-11) and (19),
 $\pi^{+}_{t}-A^{+}_{t}=k^{{A}}_{t}\sigma^{\pi}_{t}=k^{\hat{\pi}}_{t}\sigma^{\pi}_{t}+\frac{\Sigma_{B}\hat{\pi}^{B}_{t}\hat{RP}^{B}_{t}}{{Y}^{\Delta}(t)}=(\pi^{+}_{t}-\hat{\pi}^{+}_{t})+\frac{\Sigma_{B}\hat{\pi}^{B}_{t}\hat{RP}^{B}_{t}}{{Y}^{\Delta}(t)}.$ (B.4)
@@
 In turn, by Lemma 1, the LHS and RHS of (B.4) being equal implies $\mathcal{O}({\hat{RP}^{+}_{t}})=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and $\mathcal{O}({k}^{A}_{t})=\mathcal{O}(\sigma^{\pi}_{t})$, as well as $\mathcal{O}((A^{+}_{t})^{\prime}_{\pi})=\mathcal{O}((A^{+}_{t})^{\prime\prime}_{\pi})=\mathcal{O}(1)$ and $\mathcal{O}(d{\hat{RP}^{+}_{t}})=\mathcal{O}(d(\sigma^{\pi}_{t})^{2})=\mathcal{O}((\sigma^{\pi}_{t})^{4})$.
-Consider first the case of trivial $\{D_{t}\}$, so that the drift of $\{dS^{(0|A)}_{t}\}$ is nil in expectation under some RNE measure $\hat{\pi}^{B}_{0}\hat{W}_{B}^{Z}|_{T}$ satisfying $d\hat{W}^{Z}_{B,t}=\hat{r}_{B,\text{ }t}dt+\sigma^{Z}_{t}dw^{Z}_{t}$, $\hat{r}_{B,\text{ }t}dt=-d\hat{RP}^{B}_{t}$, $t<T$.
 […]
 $\frac{-\nu_{t}}{r^{Z\Delta}_{t}}=(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)(A^{+}_{t})^{\prime}_{\pi}\frac{(\sigma^{lZ}_{t})^{2}}{r^{Z\Delta}_{t}}\bigg{[}(A^{+}_{t}-{\pi}^{+}_{t})+(\sigma^{\pi}_{t})^{2}\frac{(A^{+}_{t})^{\prime\prime}_{\pi}}{2(A^{+}_{t})^{\prime}_{\pi}}-\frac{-\nu_{t}}{r^{Z\Delta}_{t}}\bigg{]}.$ (B.5)
-By Remark 12, we have $\mathcal{O}(LHS)=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and $\mathcal{O}(RHS)=\mathcal{O}((\sigma^{\pi}_{t})^{4})$, $\forall\sigma^{\pi}_{t}\ll\frac{1}{2}$, so nil-drift is possible only if each side vanishes on its own, meaning:
+By Remark 12, $\mathcal{O}(LHS)=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and $\mathcal{O}(RHS)=\mathcal{O}((\sigma^{\pi}_{t})^{4})$, $\forall\sigma^{\pi}_{t}\ll\frac{1}{2}$, so nil-drift requires each side above to vanish on its own, meaning:
 $\frac{(A^{+}_{t})^{\prime\prime}_{\pi}}{2(A^{+}_{t})^{\prime}_{\pi}}=\frac{{\pi}^{+}_{t}-A^{+}_{t}}{(\sigma^{\pi}_{t})^{2}}.$ (B.6)
-This has solutions of the form $\{A^{+}_{t}\}=\{\Pi^{+}_{t}\}$ only, where $\{\Pi^{+}_{t}\}$ is any model-risk inference based on any measure $\Pi^{B}_{0}W_{B}^{Z}|_{T}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.
+It is solved by $\{A^{+}_{t}\}=\{\Pi^{+}_{t}\}$, where $\{\Pi^{+}_{t}\}$ is any model-risk inference based on any measure of the form $\Pi^{B}_{0}W_{B}^{Z}|_{T}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.
 Adding $B$-conditionally independent data $\{D_{t}\}$ under law $W_{B}^{D}|_{T}$, the argument above applies unchanged to the dynamics of $\{dS^{(0|A)}_{t}\}$, (B.1-B.2), under the reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$ and some RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}$ such that (B.4) holds and $d\hat{\mathbf{W}}^{T}_{B,t}=\hat{r}_{B,\text{ }t}dt+d\mathbf{W}^{T}_{-,t}$, $\hat{{r}}_{B,\text{ }t}dt=-d\hat{{RP}}^{B}_{t}$.
 Part II. The viability of canonical price process $\{S^{(RP|{\Pi})}_{t}\}=\{{S}^{(0|{\Pi})}_{t}\}-\{\Sigma_{B}{RP}^{B}(t)\Pi^{B}_{t}\}$.
-Under the martingale measure $\Pi^{B}_{0}\mathbf{W}_{B}^{T}$ of process $\{{S}^{(0|{\Pi})}_{t}\}$, process $\{S^{(RP|\Pi)}_{t}\}$ is a martingale of form (B.3) plus the extra drift $R^{B}_{t}dt=-{d}RP^{B}(t)$ that stems from $B$-sure risk-pricing:
+Under the martingale measure $\Pi^{B}_{0}\mathbf{W}_{B}^{T}$ of process $\{{S}^{(0|{\Pi})}_{t}\}$, process $\{S^{(RP|\Pi)}_{t}\}$ is a martingale of form (B.3) plus the extra drift $R^{B}_{t}dt=-{d}RP^{B}(t)$ due to $B$-sure risk-pricing:
 $dS^{(RP|\Pi)}_{t}=\big{[}(\mathbf{1}_{+}-\Pi^{+}_{t})\cdot\check{r}^{Z\Delta}_{t}dt+\sigma^{Z}_{t}dw^{Z}_{t}\big{]}+S^{\Delta}(t)(\sigma^{\Pi}_{t})^{2}\big{[}(\mathbf{1}_{+}-\Pi^{+}_{t})\cdot(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw_{t}\big{]}+R^{(\cdot)}_{t}dt,$ (B.7)
@@
 C THE MOMENTUM AND LOW-RISK FORMULAE
-For the assets of Section 3.2-3.3, all the distributions required stem from the underlying log-LR process $\{l^{1/{\text{{\bf{0}}}}}_{t}\}$, of parameter $\sigma^{l}$, for which log-LR levels $l^{1/{\text{{\bf{0}}}}}_{t}$ follow the Normal distribution:
+For the assets of Section 3.2-3.3, all the distributions required derive from the underlying log-LR process $\{l^{1/{\text{{\bf{0}}}}}_{t}\}$, of parameter $\sigma^{l}$, for which log-LR levels $l^{1/{\text{{\bf{0}}}}}_{t}$ follow the Normal distribution:
 $\mathcal{N}\bigg{(}\mu^{l}_{[t]}(B)=\frac{(-1)^{B+1}}{2}(\sigma^{l})^{2}t,\text{ }(\sigma^{l}_{[t]})^{2}=(\sigma^{l})^{2}t\bigg{)},\text{ }B\in\{0,1\}.$ (C.1)
-The distributions of RNE beliefs $\{\Pi^{1}_{t}\}$ and so of prices (37) then derive via (3) by change of variable: for $\Pi^{1}_{t}=v$ given $l^{1/{\text{{\bf{0}}}}}_{t}=l$, we have $dl=(v\underline{v})^{-1}dv$, with:
+The distributions of RNE beliefs $\{\Pi^{1}_{t}\}$ and so of prices (37) then derive via (4) by change of variable: for $\Pi^{1}_{t}=v$ given $l^{1/{\text{{\bf{0}}}}}_{t}=l$, we have $dl=(v\underline{v})^{-1}dv$, since:
 $l=H_{\Pi\pm}+\log\circ\text{ }O_{f}[v],$ (C.2)
-$H_{\Pi\pm}:=\log\circ\text{ }O_{f}[\underline{\Pi}^{1}_{0\pm}]=H_{\mathbf{p}}+\log(\rho K^{\pm 1}),$ (C.3)
-$H_{\mathbf{p}}:=\log\circ\text{ }O_{f}[\underline{\mathbf{p}}^{1}_{0}],$ (C.4)
+$H_{\Pi\pm}:=\log\circ\text{ }O_{f}[\underline{\Pi}^{1}_{0\pm}]=\log\circ\text{ }O_{f}[\underline{\mathbf{p}}^{1}_{0}]+\log(\rho K^{\pm 1})=:H_{\mathbf{p}}+\log(\rho K^{\pm 1}),$ (C.3)
 where recall (39), $O_{f}[\Pi^{1}_{0\pm}]=O_{f}[\mathbf{p}^{1}_{0}]/(\rho K^{\pm 1})$ given $sign[1{\text{{\bf{0}}}}]=\pm$.
@@
 The pattern of $|C^{\Pi}_{t\pm}|$ parallels that of $|C^{\mathbf{p}}_{t}|$; it has 0 as a minimum if and when $t=(-1)^{B+1}t_{\Pi\pm}>0$.
-Bias creates a ’burden of proof’ $t_{\rho}$, and $t_{\rho}/t$ tracks how it is overcome.
+Bias thus creates a ’burden of proof’ $t_{\rho}$, and $t_{\rho}/t$ tracks how it is overcome.
 We consider cases of positive objective hurdle $H_{\mathbf{p}}>0$ and bias domination $H_{\Pi\pm}-H_{\mathbf{p}}\gg 0$, so $t_{\mathbf{p}}>0$ and $t_{\rho}\gg t_{K}$, where ’objective burden’ $t_{\mathbf{p}}$ is ’non-extreme’, to exclude easy inference or profits.
@@
 Remark 13.
-Consider profitable events $\{O_{f}[{\Pi}^{1}_{t}]=\rho^{-1}\}$ and $\{O_{f}[{\Pi}^{1}_{t\pm u}]=\rho^{-1}\}$ (see (43)), with $t$ fixed inside the window above and $u>0$ large enough for the second event to be outside the window.
+Take profitable events $\{O_{f}[{\Pi}^{1}_{t}]=\rho^{-1}\}$ and $\{O_{f}[{\Pi}^{1}_{t\pm u}]=\rho^{-1}\}$ (see (43)), with $t$ fixed inside the window above and $u>0$ large enough for the second event to be outside the window.
 Their respective probability densities are $\propto(t)^{-1/2}e^{-|C^{\mathbf{p}}_{t}|^{2}/2}$ and $\propto(t\pm u)^{-1/2}e^{-|C^{\mathbf{p}}_{(t\pm u)}|^{2}/2}$.
@@
 Hence:
-$M_{t|v}(B)$ $=(\frac{\rho v}{\underline{v}})^{-t_{K}/t}K^{-(t_{\mathbf{p}}/t)-(-1)^{B}},$ (C.11)
+$M_{t|v}(B)$ $=(\frac{\rho v}{\underline{v}})^{-t_{K}/t}K^{-(t_{\mathbf{p}}/t)-(-1)^{B}};$ (C.10)
 and likewise the volatility-conditioned ratio $M_{t|v\underline{v}}:={\mathbb{P}_{t}(+|v\underline{v})}/{\mathbb{P}_{t}(-|v\underline{v})},\text{ }v\in(0,\frac{1}{2}]$:
 $M_{t|v\underline{v}}=M_{t|v}\frac{1+R_{t|+}(\underline{v}/v)}{1+R_{t|-}(\underline{v}/v)}.$ (C.11)
-Note the source of the conditionalities above, model-risk pricing $K$.
-For $K=1$, unpriced model-risk, conditional mixes equal the unconditional.
-For $K\in(1,2)$, competitively priced model-risk, conditional mixes are perturbations of the unconditional.
+The source of the conditionalities above is model-risk pricing $K$: if $K=1$ (unpriced model-risk), conditional mixes equal the unconditional; if $K\in(1,2)$ (competitively priced model-risk), conditional mixes are perturbations of the unconditional.
 Remark 14.
-Both mix-ratios (C.11-C.12) are monotone declining in $v$, and peak-volatility ($v=\frac{1}{2}$) brings $M_{t|\frac{1}{2}}=M_{t|\frac{1}{2}\frac{1}{2}}\approx\rho^{-t_{K}/t}=\exp[{-(t_{\rho}/t)\log K}]$.
+Both mix-ratios (C.10-C.11) are monotone declining in $v$, and peak-volatility ($v=\frac{1}{2}$) brings $M_{t|\frac{1}{2}}=M_{t|\frac{1}{2}\frac{1}{2}}\approx\rho^{-t_{K}/t}=\exp[{-(t_{\rho}/t)\log K}]$.
 That is, for $\rho\gg 1$, in the window (C.7) of opportunity, when trading events are most abundant, the observed mix at high volatility can be highly negative vs the unconditional level.
 Mix reverts to its unconditional level as $t\to\infty$ when model-risk resolves.
-Any uneven unconditional background merely introduces a constant factor in (C.11-C.12); all above statements remain valid.
+Any uneven unconditional background merely introduces a constant factor in (C.10-C.11); all above statements remain valid.
 C.4 Peak-Reward Location and Size for Volatility-Conditioned Trading
@@
 Further, given (C.9) and under condition (C.7), the $\{\Pi^{1}_{t}=\underline{v}\}$-contributions to (41) are no more than $\mathcal{O}(\rho^{-t_{\rho}/{t}})$.
-That is, the leading order solution (43) is not affected by these contributions whenever $t_{\rho}/t\gg 1$, i.e.
-while bias dominates.
+That is, the leading order solution (43) is not affected by these contributions under $t_{\rho}/t\gg 1$ (i.e.
+bias domination).
 Remark 15.
@@
 For unevenness that tilts negatively (positively), the location of peak-excess moves left (right) vs the even solution (43), and its size, down (up).
-Without uncertainty or bias, $\rho=1$, both the $\{\Pi^{1}_{t}=v\}$- and $\{\Pi^{1}_{t}=\underline{v}\}$-contributions to (41) are invariant under $v\leftrightarrow\underline{v}$ switching (see (40) and (C.10-C.12)), so their derivatives both vanish at $v=\frac{1}{2}$, thus confirming peak-volatility as the location of peak-reward and ’high risk high reward’ under classical conditions.
+Without uncertainty or bias, $\rho=1$, both the $\{\Pi^{1}_{t}=v\}$- and $\{\Pi^{1}_{t}=\underline{v}\}$-contributions to (41) are invariant under $v\leftrightarrow\underline{v}$ switching (see (40) and (C.9-C.11)), so their derivatives both vanish at $v=\frac{1}{2}$, thus confirming peak-volatility as the location of peak-reward and ’high risk high reward’ under classical conditions.